### v2

# Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface

Abstract
This paper considers downlink multi-user transmission facilitated by a reconfigurable intelligent surface (RIS). First, focusing on the multi-group multicast beamforming scenario, we develop a fast and scalable algorithm for the joint base station (BS) and RIS beamforming optimization to minimize the transmit power subject to the user quality-of-service (QoS) constraints. By exploring the structure of this QoS problem, we show that the joint beamforming optimization can be naturally decomposed into a BS multicast beamforming QoS problem and an RIS passive multicast beamforming max-min-fair (MMF) problem. We propose an alternating multicast beamforming (AMBF) algorithm to solve the two subproblems alternatingly. For the BS QoS subproblem, we utilize the optimal multicast beamforming structure to obtain the BS beamformers efficiently. Furthermore, we reformulate the challenging RIS MMF subproblem and employ a first-order projected subgradient algorithm (PSA), which yields closed-form updates. The computational complexity of the AMBF algorithm grows linearly with the number of RIS elements and BS antennas. We further show that the AMBF approach is also an efficient method for the RIS-assisted downlink multi-user unicast beamforming problem, providing semi-closed-form updates. Next, we study the MMF problem for the RIS-assisted downlink beamforming design and propose a PSA-based fast algorithm to compute the BS and RIS beamforming solutions with closed-form updates per iteration, leading to a highly computationally efficient solution. Simulation results show the efficacy of our proposed algorithms in both performance and computational cost compared to other alternative methods.
I Introduction
Reconfigurable intelligent surface (RIS) is an emerging technology that can actively control and improve the wireless propagation channel conditions and thus creates a smart reconfigurable wireless environment to enhance communications performance [2, 3]. RIS uses a planar surface consisting of passive reflective elements to control the phase shifts of the reflected wireless signal towards the desired direction with little energy usage, thus forming passive beamforming to enhance the transmission performance. It offers a range of advantages over the traditional options (e.g., base station (BS) or relay), such as no need for signal processing, low energy consumption and hardware cost, convenience for deployment, etc. Thus, RIS is a promising technology that has been explored to support a wide range of applications for next-generation wireless systems, such as coverage extension, capacity enhancement, improved localization and sensing [4, 5, 6, 7].
For downlink transmissions, besides the typical dedicated unicast data transmission to each user, wireless multicasting has become increasingly popular to support content distribution and delivery in emerging applications. RIS can be explored to enhance multicast transmission performance. In conventional wireless systems, physical-layer multicast beamforming has been widely considered an efficient transmission technology for the BS to deliver common data to multiple users simultaneously. Various optimization algorithms or signal processing techniques have been developed in the literature to provide design solutions for different transmission scenarios or systems [8, 9, 10, 11, 12, 13, 14]. For delivering common data to a group, multicast beamforming performance is typically limited by the user in the group with the worst channel condition. Such limitation is dictated by the wireless prorogation environment. RIS can be utilized to break this barrier by adding an additional controllable channel path. In particular, it can form passive beamforming by controlling the phase shifts of the reflective elements to improve the user channel conditions to further enhance the overall transmission performance. Thus, it is important to explore this potential for RIS-assisted multicast beamforming design.
In conventional systems, downlink multicast beamforming is a more challenging design problem than multi-user unicast beamforming, as the former is in general an NP-hard problem [8]. This is also the case for the design of RIS-assisted downlink beamforming. RIS-assisted multicast beamforming design has recently been investigated for the single-group [15, 16] and multi-group settings [17, 18, 19, 20]. Several approaches for joint BS and RIS beamforming optimization have been proposed for either maximizing sum group rate [17, 18], or minimizing the BS transmit power subject to the user quality-of-service (QoS) target constraints, i.e., the QoS problem [19, 20]. In these works, the alternating optimization (AO) technique is commonly adopted for joint BS and RIS beamforming optimization. Depending on the specific problem formulation, smoothing techniques or SDR-based approaches have been proposed to solve the subproblems in each AO iteration. However, some limitations exist in these existing methods. For the QoS problem, existing studies are limited, and the efficacy of the proposed AO approaches in [19, 20] is uncertain as there is no guarantee of the quality of the computed solution. Also, the above existing methods generally have high computational complexity, which cannot be effectively implemented for a large number of RIS elements and BS antennas. Since RIS is beneficial only when the number of reflective elements becomes large, developing an effective design with both high performance and low computational complexity is crucial.
I-A Contribution
In this paper, we consider the RIS-assisted downlink transmission design focusing on the multicast beamforming scenario. We develop our efficient fast algorithms for joint BS and RIS beamforming optimization for both the QoS problem and the max-min fair (MMF) problem, where the MMF problem refers to maximizing the weighted minimum SINR among users subject to the BS transmit power budget. Furthermore, we generalize our proposed approach and show that it can be extended to the unicast beamforming scenario, lending to a fast beamforming solution.
Several unique features of our proposed algorithms distinguish our work from the existing ones. First, we propose a fast alternating multicast beamforming (AMBF) approach for the QoS problem. Different from the existing methods that directly apply AO to this problem, we explore the structure of the joint BS and RIS beamforming optimization and show that this QoS problem can be naturally broken down into two multicast beamforming subproblems, which are solved alternatingly: a BS multicast beamforming QoS problem and an RIS passive multicast beamforming MMF problem. This structure-based AMBF approach can effectively improve the BS and RIS beamforming solutions over iterations, avoiding the issue faced in the existing direct AO approaches.
Next, with the two subproblems identified, we can judiciously exploit the optimal multicast beamforming structure [11] to compute a solution efficiently for the BS multicast problem, whose computational complexity does not depend on the number of BS antennas. This departs from the existing methods that are purely based on computational optimization techniques. For the challenging RIS passive multicast MMF problem, we develop our technique to reformulate the problem and adopt a first-order projected subgradient algorithm (PSA) [12] to solve it, which yields computationally cheap closed-form updates and has a convergence guarantee. Finally, we show that since our AMBF approach utilizes the structure of the problem, it is readily applicable to the downlink multi-user unicast beamforming scenario to compute the solution efficiently. Our fast algorithm is scalable, with computational complexity growing linearly with the number of RIS elements and the BS antennas.
For the RIS-assisted downlink multicast MMF problem, we first show it is an inverse problem to the earlier QoS problem. Following this, we utilize the optimal BS beamforming structure again to transform the MMF problem into an equivalent problem, which allows us to propose a PSA-based iterative algorithm to compute the BS and RIS beamforming solutions. In particular, the algorithm uses closed-form updates per iteration, leading to a highly computationally efficient solution. The proposed algorithm is also directly applicable to the general downlink RIS-assisted multi-user beamforming problem with the MMF objective.
The simulation results demonstrate the effectiveness of our proposed algorithms, in terms of the performance and computational advantage, for both the multicast and unicast scenarios compared to the alternative methods.
I-B Related Work
Multicast beamforming optimization is inherently a non-convex and NP-hard problem [8]. The earlier literature works adopted the semi-definite relaxation (SDR) approach [8, 9] to find an approximate solution. With the number of antennas increasing, the successive convex approximation (SCA) technique has become more attractive for its lower computational complexity and better performance [10]. The optimal multi-group multicast beamforming structure was recently obtained in [11]. It shows that to obtain the beamforming vectors, we only need to optimize the per-user weight, which is a much smaller optimization problem that is independent to the number of BS antennas. This structure has been exploited to develop fast and scalable algorithms for multicast beamforming in massive multiple-input multiple-output (MIMO) systems [12, 13, 14].
Several recent works have studied the RIS-assisted multicast beamforming design for the single group [15, 16] and the general multi-group settings [17, 18, 19, 20]. Joint BS and RIS beamforming optimization for maximizing sum group rate is considered in [17, 18]. In [17], the alternating optimization (AO) technique is used, and a smoothing technique is proposed along with a majorization minimization method to solve the subproblems. In [18], the smoothing technique is applied directly to the rate objective, along with a low-complexity alternating projected gradient algorithm. The QoS problem for the BS transmit power minimization is studied in [19], where AO is directly applied to the problem, resulting in a feasibility problem for the RIS passive beamforming design. The SDR-based approach is then proposed to solve the subproblems. A similar approach is also considered in [20] for an RIS-assisted symbiotic radio multicast system comprising the primary receivers and an internet-of-things receiver. However, as the feasibility problem cannot effectively search for an improved solution over iteration, there is no guarantee of the quality of this AO approach.
The unicast scenario for RIS-assisted downlink multi-user beamforming has been studied in many recent works. Among them, only a few works have studied the BS power minimization problem [5, 21]. In [5], directly applying AO to the joint beamforming optimization problem is first considered, and a low-complexity heuristic method based on the sum of channel gains is proposed. RIS-assisted non-orthogonal multiple access (NOMA) network is considered in [21], and the joint beamforming design using the direct AO approach is also adopted. Both [5] and [21] adopt semi-definite programming (SDP) or SDR to solve the corresponding alternating subproblems. The minimum rate (or SINR) maximization problem under various system configurations has been considered in [22, 23, 24]. In [22], an iterative algorithm providing an alternating descent path is proposed. The work in [23] uses asymptotic analysis and some assumptions for the channel model to provide the approximations to the optimal BS transmit beamformer. Then, a projected gradient ascent algorithm is proposed for the RIS beamformer based on the large-scale channel statistics. RIS-assisted multi-cell coordinated transmit beamforming is considered in [24], where RIS can adjust both reflecting phases and amplitudes. The AO approach is adopted, and the second-order cone programming (SOCPs) and SDR are used to solve the subproblems. An inexact AO approach with SCA is further considered to reduce the computational complexity. However, these existing approaches generally have computational complexity.
I-C Organization and Notations
The rest of this paper is organized as follows. Section II presents the system model and the problem formulation for the RIS-assisted multicast beamforming QoS problem. In Section III, we develop our AMBF approach to break the QoS problem into two subproblems to solve alternatingly. In Section IV, we present the fast algorithms to solve each beamforming subproblem. In Section V, we generalize our approach to solve the multi-user unicast beamforming. The MMF problem is considered in Section VI, where we discuss the relation between the QoS and MMF problems and propose a fast first-order algorithm to solve the MMF problem. The simulation results is provided in Section VII, followed by the conclusion in Section VIII.
[…]
II System Model
II-A System Model
We consider an RIS-assisted downlink multi-group multicast scenario: a BS equipped with $N$ antennas multicasts messages to $G$ user groups, and an RIS consisting of $M$ passive reflective elements is deployed to assist the data transmission between the BS and the users. We assume that group $i$ consists of $K_{i}$ single-antenna users, and these users receive a common message from the BS that is independent of other groups. The total number of users is denoted by $K_{\text{tot}}=\sum_{i=1}^{G}K_{i}$. The BS controls the phases of the RIS array elements via a RIS controller to adjust the signals received at the RIS to the desired directions. We denote the set of group indices by ${\cal G}\triangleq\{1,\ldots,G\}$, the set of user indices in group $i$ by ${\cal K}_{i}\triangleq\{1,\ldots,K_{i}\}$, $i\in{\cal G}$, and the set of RIS element indices by ${\cal M}\triangleq\{1,\ldots,M\}$.
Let ${\bf H}_{\text{r}}$ denote the $M\times N$ channel matrix from the BS to the RIS, ${\bf h}^{\text{d}}_{ik}$ the $N\times 1$ channel vector from the BS to user $k$ in group $i$, and ${\bf h}^{\text{r}}_{ik}$ the $M\times 1$ channel vector from the RIS to user $k$ in group $i$, for $k\in{\cal K},i\in{\cal G}$. We use ${\bf w}_{i}$ to represent the $N\times 1$ multicast beamforming vector at the BS for group $i\in{\cal G}$. Also, ${\bf e}\triangleq[e_{1},\ldots,e_{M}]^{T}$ represents the vector containing the $M$ reflection coefficients at the RIS, where $e_{m}=e^{j\theta_{m}}$ with $\theta_{m}\in(-\pi,\pi]$ being the phase shift of element $m$. We refer to ${\bf e}$ as the RIS passive beamforming vector. The signal received at user $k$ in group $i$ is given by
$\!\!y_{ik}$ $=\sum_{j=1}^{G}{\bf w}_{j}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf H}_{\text{r}}^{H}\text{diag}({\bf e}){\bf h}^{\text{r}}_{ik}\big{)}s_{j}+n_{ik}$
$={\bf w}_{i}^{H}\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}s_{i}+\!\sum_{j\neq i}\!{\bf w}_{j}^{H}\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}s_{j}+n_{ik}$ (1)
where $s_{i}$ is the symbol intended for group $i$ with $\mathbb{E}[|s_{i}|^{2}]=1$, ${\bf G}_{ik}\triangleq{\bf H}_{\text{r}}^{H}\text{diag}({\bf h}_{ik}^{\text{r}})$ is the $N\times M$ cascaded channel from the BS to user $k$ in group $i$ via the RIS, $n_{ik}\sim{\cal C}{\cal N}(0,\sigma^{2})$ is the receiver additive white Gaussian noise with variance $\sigma^{2}$.
The received SINR at user $k$ in group $i$ is given by
$\textrm{SINR}_{ik}=\frac{|{\bf w}_{i}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}|^{2}}{\sum_{j\neq i}|{\bf w}_{j}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}|^{2}+\sigma^{2}},$ (2)
and the total transmit power at the BS is $\sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$.
II-B Problem Formulation
We first focus on the QoS problem for the RIS-assisted multicast beamforming design. Our objective is to jointly optimize the multicast beamforming vectors ${\bf w}_{i}$ and RIS reflection coefficient vector ${\bf e}$ to minimize the BS transmit power while ensuring that the SINR target at each user is met. This joint optimization problem is formulated as
${\cal P}_{o}:\ \min_{{\bf w},{\bf e}}$ $\ \ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$ (3a)
s.t. $\ \ \textrm{SINR}_{ik}\geq\gamma_{ik},\ k\in{\cal K}_{i},~{}i\in{\cal G}$ (3b)
$\ \ |e_{m}|^{2}=1,\ m\in{\cal M}$ (3c)
where ${\bf w}\triangleq[{\bf w}_{1}^{H},\ldots,{\bf w}_{G}^{H}]^{H}$, and $\gamma_{ik}$ is the SINR target for user $k$ in group $i$.
Problem ${\cal P}_{o}$ is a non-convex quadratically constrained quadratic programming (QCQP) problem. Note that for the conventional downlink transmission without RIS, the multicast beamforming problem is already known to be NP-hard in its general form. The joint optimization problem ${\cal P}_{o}$ is even more challenging with the additional RIS passive beamformer ${\bf e}$ affecting $\textrm{SINR}_{ik}$ to be optimized under the non-convex constraints in (3c) imposed. Furthermore, since the number of RIS elements $M$ is expected to be large to show benefit, ${\cal P}_{o}$ is typically a large-scale problem. Thus, our goal is to develop an effective computational method to provide a solution with high performance but low computational complexity.
III RIS-assisted Multicast Beamforming Design
We consider the alternating optimization (AO) technique to solve ${\cal P}_{o}$, i.e., solving ${\cal P}_{o}$ with respect to (w.r.t.) the BS multicast beamformer ${\bf w}$ and RIS passive beamformer ${\bf e}$ alternatingly. However, directly applying AO to ${\cal P}_{o}$ is not an effective method. This is because the objective function in ${\cal P}_{o}$ is only a function of ${\bf w}$. With given ${\bf w}$, ${\cal P}_{o}$ is reduced to a feasibility problem w.r.t. ${\bf e}$, for which the value of ${\bf e}$ from the previous AO iteration is already a feasible solution. Thus, the algorithm could get stuck at the value of $({\bf w},{\bf e})$ after the first round of updates, and the quality of this solution merely depends on the quality of the initial point.
In our approach, we first analyze the structure of ${\cal P}_{o}$ w.r.t. ${\bf w}$ and ${\bf e}$ and then utilize it to devise an effective algorithm to solve this QoS problem.
III-A QoS Problem Reformulation
To explore the structural relation between ${\bf w}$ and ${\bf e}$ in ${\cal P}_{o}$, we first convert the set of SINR constraints in (3b) into the following equivalent constraint:
$\textrm{SINR}_{ik}\geq\gamma_{ik},\ \forall k,\forall i\quad$ $\Leftrightarrow\quad\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1,\ \forall k,\forall i$
$\Leftrightarrow\quad\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1.$ (4)
We can replace (3b) with (4) and equivalently rewrite ${\cal P}_{o}$ as
${\cal P}_{1}:\ \min_{{\bf w},{\bf e}}$ $\ \ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$ (5a)
s.t. $\ \ \min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1$ (5b)
$\ \ |e_{m}|^{2}=1,\ m\in{\cal M}.$ (5c)
Noticing that RIS passive beamforming vector ${\bf e}$ only affects $\textrm{SINR}_{ik}$’s in (5b), we can combine (5b) and (5c) and further transform ${\cal P}_{1}$ into the following equivalent problem:
${\cal P}_{2}:\ \min_{{\bf w}}$ $\ \ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$ (6a)
s.t. $\ \ \max_{{\bf e}:|e_{m}|^{2}=1,m\in{\cal M}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1.$ (6b)
Note that for showing the equivalence, it is straightforward to show ${\cal P}_{1}$ and ${\cal P}_{2}$ have the same feasible set.
Comparing ${\cal P}_{2}$ with ${\cal P}_{o}$, we see that the original set of SINR constraints in (3b) is replaced by the constraint in (6b). The left-hand side of (6b) is a max-min weighted SINRs optimization problem w.r.t. ${\bf e}$, and the unit-modulus constraints on ${\bf e}$ in (3c) are now the constraints for this max-min optimization problem.
III-B Alternating Multicast Beamforming Approach
Once transforming ${\cal P}_{o}$ into ${\cal P}_{2}$, we notice that ${\cal P}_{2}$ inherently contains two optimization subproblems w.r.t. ${\bf w}$ and ${\bf e}$, respectively. As we will show below, these two subproblems are essentially two types of multicast beamforming problems: 1) BS multicast beamforming QoS problem w.r.t. ${\bf w}$; 2) RIS passive multicast beamforming MMF problem w.r.t. ${\bf e}$. Following this, we naturally break down ${\cal P}_{2}$ into two subproblems to solve alternatingly. Our overall proposed alternating multicast beamforming (AMBF) approach includes two steps: the alternating optimization step and a final processing step. The final processing step is to ensure a feasible solution to ${\cal P}_{o}$. Below, we will first describe the two subproblems and the final processing step, leaving the fast algorithm design for solving the subproblems to Sections IV-A and IV-B.
III-B1 BS Multicast Beamforming QoS Problem for ${\bf w}$
For given RIS passive beamforming vector ${\bf e}$, ${\cal P}_{o}$ is reduced to the following BS multicast beamforming problem w.r.t. ${\bf w}$:
${\cal P}_{{\bf e}}:\ \min_{{\bf w}}$ $\ \ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$ (7a)
s.t. $\ \ \textrm{SINR}_{ik}\geq\gamma_{ik},\ k\in{\cal K}_{i},~{}i\in{\cal G}.$ (7b)
Problem ${\cal P}_{\bf e}$ is a typical downlink multi-group multicast beamforming QoS problem. The effective channel between the BS and user $k$ in group $i$ consists of the direct path and the RIS-path given by ${\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}$, as shown in $\textrm{SINR}_{ik}$ in (2).
III-B2 RIS Passive Multicast Beamforming MMF Problem for ${\bf e}$
Optimizing RIS passive beamformer ${\bf e}$ with given ${\bf w}$ is the main challenge in ${\cal P}_{o}$. As discussed earlier, directly applying AO to the original problem ${\cal P}_{o}$ is not effective. However, once we convert ${\cal P}_{o}$ into ${\cal P}_{2}$, for given ${\bf w}$, we naturally have the following optimization problem w.r.t. ${\bf e}$, which is to maximize the minimum ratio of the received SINR over its target, or the weighted SINR, i.e., ${\textrm{SINR}_{ik}}/{\gamma_{ik}}$:
${\cal S}_{{\bf w}}:\ \max_{{\bf e}}\min_{i,k}$ $\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ (8a)
s.t. $\ \ |e_{m}|^{2}=1,\ m\in{\cal M}.$ (8b)
Remark 1.
Note that ${\cal S}_{{\bf w}}$ can be viewed as a weighted MMF single-group multicast beamforming problem, with ${\bf e}$ being the beamformer to all the users in the system. However, this problem has three key differences from the conventional MMF multicast problem: i) This single-group multicast beamforming problem contains self-interference, as ${\bf e}$ appears in the denominator of $\textrm{SINR}_{ik}$ in (2). ii) From the point of view of optimizing ${\bf e}$, $\textrm{SINR}_{ik}$ in (2) can in fact be interpreted as the signal-to-leakage-and-noise ratio (SLNR) for user $k$ in group $i$. Thus, the MMF problem ${\cal S}_{{\bf w}}$ w.r.t. ${\bf e}$ is based on an SLNR metric. iii) Different from the conventional transmit power constraint, the constraints in (8b) can be viewed as a non-convex per-element power constraint, where each element is required to consume exactly one unit of power.
Due to the difference discussed above, ${\cal S}_{\bf w}$ is a more challenging problem to solve than the conventional MMF problem, which is a non-convex QCQP problem that is NP-hard. To tackle this problem, we note that the constraints in (8b) are equivalent to
$|e_{m}|^{2}\leq 1,m\in{\cal M};\quad\quad{\bf e}^{T}{\bf e}=M.$ (9)
Thus, we replace the constraints in (8b) with (9) and transform ${\cal S}_{\bf w}$ into the following equivalent problem:
${\cal S}_{{\bf w}}^{\text{eq}}:\ \max_{{\bf e}}\min_{i,k}$ $\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ (10a)
s.t. $\ \ |e_{m}|^{2}\leq 1,\ m\in{\cal M}$ (10b)
$\ \ {\bf e}^{T}{\bf e}=M.$ (10c)
To make the problem more tractable, we relax ${\cal S}_{\bf w}^{\text{eq}}$ by transferring the constraint in (10c) into the objective function in (10a) as a penalty term with a penalty weight $\zeta>0$. The relaxed problem is given by
$\widetilde{{\cal S}}_{{\bf w}}:\ \max_{{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\zeta\frac{{\bf e}^{H}{\bf e}}{M}$ (11a)
s.t. $\ \ |e_{m}|^{2}\leq 1,\ m\in{\cal M}.$ (11b)
Remark 2.
The penalty term in (11a) is to ensure that ${\bf e}^{T}{\bf e}$ is as close to $M$ as possible in solving $\widetilde{{\cal S}}_{{\bf w}}$. Note that we choose the penalty term in this form such that the two terms in the objective function are both normalized. In particular, for the weighted SINR term, using $({\bf w},{\bf e})$ from ${\cal P}_{\bf e}$, we have $\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}=1$ based on (7b). Thus, we expect this term to be close to $1$, especially as the AO iteration between ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{\bf w}$ converges. The RIS passive beamformer norm term satisfies $\frac{{\bf e}^{H}{\bf e}}{M}\leq 1$ based on (10b), and it is also expected to be close to $1$ as the AO iteration converges. In this way, the two terms in (11a) have comparable values that are both around $1$, which are not affected by the actual values of SINR or $\|{\bf e}\|^{2}$ for a specific setup. This simplifies the tuning of penalty parameter $\zeta$, which can remain the same for different system configurations.
Although $\widetilde{{\cal S}}_{{\bf w}}$ is still a non-convex QCQP problem, compared with ${\cal S}_{{\bf w}}$, it is more amenable to efficient algorithm design to compute a solution, which will be discussed in Section IV-B.
Thus, following the above, the first step in our AMBF approach is to solve the two multicast beamforming problems alternatingly, i.e., ${\cal P}_{\bf e}$ for the BS beamformers $\{{\bf w}_{i}\}$ and $\widetilde{{\cal S}}_{{\bf w}}$ for the RIS beamformer ${\bf e}$.
III-B3 Final Processing
Let $({\bf w}^{\star},{\bf e}^{\star})$ denote the solution after the above iterative procedure. Since $\widetilde{{\cal S}}_{{\bf w}}$ is a relaxed problem of ${\cal S}_{{\bf w}}$, $({\bf w}^{\star},{\bf e}^{\star})$ may not be feasible to ${\cal P}_{o}$. Thus, based on $({\bf w}^{\star},{\bf e}^{\star})$, we have the final processing step to obtain the feasible solution $({\bf w}^{\text{final}},{\bf e}^{\text{final}})$ as follows:
  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which represents taking the phase component of each element in ${\bf e}^{\star}$.
  • Solve ${\cal P}_{{\bf e}^{\text{final}}}$ with ${\bf e}^{\text{final}}$ and obtain ${\bf w}^{\text{final}}$.
IV Fast Algorithms for the AMBF Approach
In this section, we propose our fast algorithms to solve subproblems ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{{\bf w}}$.
IV-A Fast Algorithm for BS Multicast Beamforming ${\cal P}_{\bf e}$
As mentioned earlier, ${\cal P}_{\bf e}$ is a typical downlink multicast beamforming QoS problem. Define $\tilde{{\bf h}}_{ik}\triangleq{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}$ as the $N\times 1$ equivalent channel from BS to user $k$ in group $i$ that includes both the direct path and the RIS path. The received SINR at user $k$ in group $i$ in (2) can be rewritten as
$\textrm{SINR}_{ik}=\frac{|{\bf w}_{i}^{H}\tilde{{\bf h}}_{ik}|^{2}}{\sum_{j\neq i}^{G}|{\bf w}_{j}^{H}\tilde{{\bf h}}_{ik}|^{2}+\sigma^{2}},\ k\in{\cal K}_{i},i\in{\cal G}.$ (12)
Although the QoS problem ${\cal P}_{\bf e}$ is known to be a challenging non-convex QCQP problem, the structure of the optimal solution is obtained in [11], which can be utilized to compute the beamforming solution ${\bf w}$ with high computational efficiency, especially when the number of BS antennas is large $M\gg 1$. We directly employ this optimal structure to compute ${\bf w}$. Specifically, the optimal ${\bf w}_{i}$ is a weighted MMSE beamforming given by [11]
${\bf w}_{i}\,=\,{\bf R}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf H}}_{i}{\bf a}_{i}\,,\;i\in{\cal G}$ (13)
where $\widetilde{{\bf H}}_{i}\triangleq[\tilde{{\bf h}}_{i1},\ldots,\tilde{{\bf h}}_{iK}]$ is the equivalent channel matrix for group $i$, ${\bf a}_{i}$ is the $K\times 1$ optimal weight vector for group $i$ containing the (complex) weight for each user channel in the group, and ${\bf R}(\boldsymbol{\lambda})\triangleq{\bf{I}}+\sum_{i=1}^{G}\sum_{k=1}^{K}\lambda_{ik}\gamma_{ik}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}$ is the noise-plus-weighted-channel covariance matrix with $\lambda_{ik}$ being the optimal Lagrange multipliers associated with the SINR constraint in (7b) and $\boldsymbol{\lambda}$ being the vector containing all $\lambda_{ik}^{o}$’s.
The solution ${\bf w}_{i}$ in (13) is given in a semi-closed form with $\boldsymbol{\lambda}$ and $\{{\bf a}_{i}\}$ to be numerically determined. Although the optimal value for $\boldsymbol{\lambda}$ is difficult to obtain, an algorithm to compute (suboptimal) values of $\boldsymbol{\lambda}$ efficiently using the fixed-point method is proposed in [11]. With the optimal structure of ${\bf w}_{i}$ in (13) and $\boldsymbol{\lambda}$, ${\cal P}_{\bf e}$ can be transformed into a weight optimization problem w.r.t. weight vectors $\{{\bf a}_{i}\}$, which is a much smaller problem with $\sum_{i=1}^{G}K_{i}$ variables compared to the original problem ${\cal P}_{\bf e}$ with $GN$ variables for $N\gg K_{i}$. To solve this non-convex weight optimization problem for $\{{\bf a}_{i}\}$, we can adopt the SCA approach discussed in [11]. Further improvement in the computational efficiency can be considered by adopting the fast ADMM algorithm proposed recently in [13] based on the optimal structure, which provides each SCA update using closed-form expressions with ultra-low complexity. We omit the details and refer the readers to [11, 13] for these algorithms.
IV-B Fast Algorithm for RIS Passive Multicast Beamforming $\widetilde{{\cal S}}_{\bf w}$
Problem $\widetilde{{\cal S}}_{{\bf w}}$ is expected to be large-size, since the number of RIS elements is large ($M\gg 1$). Existing methods in the literature for solving ${\bf e}$ often adopt SDR or SCA [17, 20], but suffer from high computational complexity. It is essential to develop an effective low-complexity algorithm to find a solution to $\widetilde{{\cal S}}_{{\bf w}}$, especially when this subproblem needs to be solved in each AO iteration.
From Remark 1, we note that $\widetilde{{\cal S}}_{{\bf w}}$ is a variant of the MMF problem. PSA is a fast first-order algorithm recently proposed in[12] to solve the multi-group multicast beamforming MMF problem in the conventional system. Although the form of the objective function and the constraints in $\widetilde{{\cal S}}_{{\bf w}}$ are different from those of the conventional MMF problem considered in [12], the method can still be adapted to solve $\widetilde{{\cal S}}_{{\bf w}}$. Thus, we propose to apply PSA to $\widetilde{{\cal S}}_{{\bf w}}$ to compute a near-stationary solution of $\widetilde{{\cal S}}_{{\bf w}}$ efficiently.
IV-B1 Problem Reformulation
Based on the objective function in (11a) and using the SINR expression in (2), we define
$\phi_{ik}({\bf e})\triangleq-\bigg{(}\frac{1}{\gamma_{ik}}\frac{|{\bf w}_{i}^{H}({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})|^{2}}{\sum_{j\neq i}|{\bf w}_{j}^{H}({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})|^{2}\!+\!\sigma^{2}}+\zeta\frac{{\bf e}^{H}{\bf e}}{M}\bigg{)},$ (14)
for $k\in{\cal K}_{i}$, $i\in{\cal G}$. Furthermore, we denote the feasible set of $\widetilde{{\cal S}}_{{\bf w}}$ as ${\cal E}\triangleq\{{\bf e}:|e_{m}|^{2}\leq 1,\ m\in{\cal M}\}$. Then, $\widetilde{{\cal S}}_{{\bf w}}$ can be rewritten as a min-max problem:
$\widetilde{{\cal S}}_{{\bf w}}:\min_{{\bf e}\in{\cal E}}\max_{i,k}\phi_{ik}({\bf e}).$ (15)
The inner maximization problem is an integer program. To tackle this, let $\boldsymbol{\phi}({\bf e})\triangleq[\phi_{11}({\bf e}),\ldots,\phi_{1K_{1}}({\bf e}),\ldots,\phi_{G1}({\bf e}),\ldots,\phi_{GK_{G}}({\bf e})]^{T}$, and also let ${\bf y}$ be a $K_{\text{tot}}\times 1$ probability vector satisfying ${\bf y}\succcurlyeq 0$ and $\mathbf{1}^{T}{\bf y}=1$. Then, we can further transform problem (15) into the following equivalent problem
$\widetilde{{\cal S}}_{{\bf w}}^{\prime}:\min_{{\bf e}\in{\cal E}}\max_{{\bf y}\in{\cal Y}}$ $\ f({\bf e},{\bf y})$ (16)
where $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y}$ and ${\cal Y}\triangleq\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$. Note that ${\cal Y}$ is a probability simplex. Thus, an optimal solution to the inner maximization problem in $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is ${\bf y}=[0,\cdots,1,\cdots,0]^{T}$, with 1 at some position that corresponds to $\max_{i,k}\phi_{ik}({\bf e})$, which is equivalent to the inner maximization in (15). Note that both ${\cal E}$ and ${\cal Y}$ are compact convex sets.
IV-B2 The Projected Subgradient Algorithm
Note that $f({\bf e},{\bf y})$ is concave in ${\bf y}$ but non-convex in ${\bf e}$. Thus, problem $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is a non-convex-concave min-max problem, which is typically NP-hard. Furthermore, the objective function $\max_{{\bf y}\in{\cal Y}}f({\bf e},{\bf y})$ may not be differentiable. To solve $\widetilde{{\cal S}}^{\prime}$, we notice that $\widetilde{{\cal S}}^{\prime}$ has the same structure as the min-max problem considered in [12] (i.e., with objective function $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y}$ and a convex feasible set). Thus, we adopt PSA in [12] to find a near-stationary point of $\widetilde{\cal S}^{\prime}_{{\bf w}}$.
Specifically, PSA is an iterative algorithm, with the updating procedure at iteration $j$ given by [^1] Note that PSA in [12] is provided using all real-valued variables. We provide the complex-valued version of this procedure, which can be shown to be equivalent to the real-valued version.
${\bf y}^{(j)}$ $\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}f({\bf e}^{(j)},{\bf y});$ (17)
${\bf e}^{(j+1)}$ $=\Pi_{{\cal E}}\big{(}{\bf e}^{(j)}-\alpha\nabla_{{\bf e}}f({\bf e}^{(j)},{\bf y}^{(j)})\big{)}$ (18)
where $\alpha>0$ is the step size, and $\Pi_{{\cal E}}({\bf e})$ is the projection of ${\bf e}$ onto the feasible set ${\cal E}$. Since the constraints in ${\cal E}$ is for each $e_{m}$, $\Pi_{{\cal E}}({\bf e})$ performs per-element projection as
$\Pi_{{\cal E}}({\bf e})=\begin{cases}e_{m}&\mbox{if }|e_{m}|\leq 1\\ \frac{e_{m}}{|e_{m}|}&\text{o.w}\end{cases},\ \text{for}\ m\in{\cal M}.$ (19)
In (17), ${\bf y}^{(j)}$ can be obtained by finding index $\hat{i}\hat{k}$ such that $\phi_{\hat{i}\hat{k}}({\bf e})=\max_{i,k}\phi_{ik}({\bf e})$, and we have $f({\bf e},{\bf y}^{(j)})=\phi_{\hat{i}\hat{k}}({\bf e})$. To compute the gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (18), with a slight abuse of notation to ease our presentation, we use $\phi_{ik}({\bf e})$ to represent $\phi_{i^{\prime}k^{\prime}}({\bf e})$ obtained in (17). Then, $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ has the following expression:
$\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})=\nabla_{{\bf e}}\phi_{\hat{i}\hat{k}}({\bf e})$
$\quad=-\frac{{\bf q}_{\hat{i}\hat{i}\hat{k}}({\bf e})I_{\hat{i}\hat{k}}\!({\bf e})-\!\sum_{j\neq\hat{i}}{\bf q}_{j\hat{i}\hat{k}}({\bf e})S_{\hat{i}\hat{k}}({\bf e})}{\gamma_{\hat{i}\hat{k}}I_{\hat{i}\hat{k}}^{2}({\bf e})}-\frac{\zeta}{M}{\bf e}$ (20)
where $I_{\hat{i}\hat{k}}({\bf e})\triangleq\sum_{j\neq\hat{i}}|{\bf w}_{j}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})|^{2}+\sigma^{2}$, $S_{\hat{i}\hat{k}}({\bf e})\triangleq|{\bf w}_{\hat{i}}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})|^{2}$, and ${\bf q}_{j\hat{i}\hat{k}}({\bf e})={\bf G}_{\hat{i}\hat{k}}^{H}{\bf w}_{j}{\bf w}_{j}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})$.
The updates in each iteration are closed-form evaluations as in (17)–(IV-B2), with only matrix and vector multiplications. Thus, PSA is computationally cheap and is particularly suitable for solving $\widetilde{{\cal S}}_{{\bf w}}$ with a large value of $M$. Furthermore, the convergence analysis of PSA in [12] is applicable to our problem. It shows that the above PSA procedure is guaranteed to converge in finite time to a point at the vicinity of a stationary point of $\widetilde{{\cal S}}_{{\bf w}}$.
IV-C Discussion on the AMBF Algorithm
We summarize our proposed fast AMBF algorithm for the RIS-assisted multicast QoS problem ${\cal P}_{o}$ in Algorithm 1. It consists of the proposed AMBF approach in Section III-B and the fast algorithms in Sections IV-A and IV-B for each subproblem. A few aspects of the algorithm are discussed below:
Algorithm 1 Alternating Multicast Beamforming Algorithm (AMBF) for RIS-Assisted QoS Problem ${\cal P}_{o}$.
1:Initialization: Set feasible initial point ${\bf e}^{0}$; Set $n=0$.
2:repeat
3:     With ${\bf e}^{n}$, solve ${\cal P}_{\bf e}$ to obtain ${\bf w}^{n+1}$ by using (13).
4:     With ${\bf w}^{n}$, set ${\bf e}^{(0)}={\bf e}^{n}$; Solve $\widetilde{{\cal S}}_{{\bf w}}$ using PSA updates (17) and (18) to obtain ${\bf e}^{n+1}$.
5:     Set $n\leftarrow n+1$.
6:until convergence
7:Set ${\bf e}^{\text{final}}=\Pi_{{\cal E}}({\bf e}^{n})$.
8:Obtain ${\bf w}^{\text{final}}$ by solving ${\cal P}_{{\bf e}^{\text{final}}}$.
9:return $({\bf w}^{\text{final}},{\bf e}^{\text{final}})$.
IV-C1 Initialization
The AMBF algorithm requires an initial feasible point ${\bf e}^{0}$ that satisfies the constraints in (3c) for the BS multicast beamforming subproblem ${\cal P}_{\bf e}$. It can be easily obtained by generating a random phase for each element in ${\bf e}^{0}$. In each AO iteration $n$, solving $\tilde{{\cal S}}_{\bf w}$ by PSA also requires an initial point ${\bf e}^{(0)}$ for the update in (17). We should set ${\bf e}^{(0)}={\bf e}^{n}$, which is from the previous AO iteration and is also used in ${\cal P}_{\bf e}$ in the current AO iteration.
IV-C2 Computational Complexity
Algorithm 1 consists of solving two optimization subproblems ${\cal P}_{{\bf e}}$ and $\widetilde{{\cal S}}_{{\bf w}}$ at each AO iteration. We discuss the computational complexity for solving each subproblem below:
Problem ${\cal P}_{\bf e}$
As mentioned in Section IV-A, using the optimal beamforming structure ${\bf w}$ in (13), ${\cal P}_{{\bf e}}$ is converted into the weight optimization problem w.r.t. $\{{\bf a}_{i}\}$ with total $K_{\text{tot}}(=\sum_{i=1}^{G}K_{i})$ variables, which can be solved by SCA [11]. If each SCA subproblem is solved using a typical interior point method [25] by the standard convex solver, the computational complexity is $O(K_{\text{tot}}^{3})$ per SCA iteration, which does not grow with either $M$ or $N$. We point out that this computational complexity can be further reduced by using the fast ADMM algorithm proposed recently in [26], which only involves closed-form updates and has the computational complexity in the order of $O(K_{\text{tot}}^{2})$.
Problem ${\cal S}_{{\bf w}}$
The problem is solved using PSA updates in (17) and (18) per iteration. Obtaining ${\bf y}^{(j)}$ in (17) involves computing $\textrm{SINR}_{ik}$ in (2) for all $K_{\text{tot}}$ users and finding the maximum $\phi_{ik}({\bf e})$, which requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+4)$ flops. The required computation in (18) is mainly at computing the gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (IV-B2). We note that all the terms in (IV-B2), i.e., ${\bf q}_{j\hat{i}\hat{k}}$, $I_{\hat{i}\hat{k}}({\bf e})$, and $S_{\hat{i}\hat{k}}({\bf e})$ are part of the SINR expression that has already been computed. Thus, calculating $\nabla_{{\bf e}}\phi_{\hat{i}\hat{k}}({\bf e})$ only requires $2GM+2M$ flops. Thus, the entire updates per iteration requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+1)+2GM+2M$ flops, with the leading complexity being $K_{\text{tot}}N(G+M)$ flops. As typically we expect $M\gg G$, the leading computational complexity is $K_{\text{tot}}NM$. Thus, the complexity of PSA grows linearly over $M$, $N$, and $K_{\text{tot}}$. The complexity of the final per-element projection in (19) is $O(M)$. Thus, the overall computational complexity in each AO iteration of Algorithm 1 is as low as $O(K_{\text{tot}}^{2}+K_{\text{tot}}MN)$, which is linear in both $M$ and $N$.
IV-C3 Convergence
For solving ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{\bf w}$ alternatingly, note that the objective of the minimization problem ${\cal P}_{\bf e}$ is lower bounded and that of the maximization problem $\widetilde{{\cal S}}_{\bf w}$ is upper bounded. Thus, the alternating procedure is guaranteed to converge.
V Generalization to RIS-Assisted Multi-user Downlink Beamforming
Problem ${\cal P}_{o}$ in fact represents the general formulation of downlink RIS-assisted multi-user beamforming QoS problem that includes both multicast and unicast transmission scenarios. Unicast transmission is the most commonly considered scenario in the downlink, where the BS sends data dedicated to each user. It is a special case of multicast beamforming when $K=1$ for each group, and the multicast beamforming problem is reduced to the conventional unicast beamforming for $G$ users. Thus, our proposed AMBF algorithm in Algorithm 1 can be directly used to solve the RIS-assisted multi-user unicast beamforming problem.
More importantly, we note that the transformed problem ${\cal P}_{2}$ in (III-A) reflects the general RIS-assisted downlink beamforming problem for either unicast or multicast transmissions. As a result, the AMBF approach in Section III-B to alternatingly solve the BS beamforming problem ${\cal P}_{\bf e}$ and the RIS passive beamforming problem ${\cal S}_{\bf w}$ is applicable in both scenarios.
So far, we have focused on the multicast beamforming problem ${\cal P}_{o}$ for $K>1$, which is a more challenging problem to solve than the unicast beamforming problem. For the unicast scenario, the RIS beamforming problem ${\cal S}_{\bf w}$ is still a multicast beamforming MMF problem as discussed in Remark 1 of Section III-B2, which can be efficiently solved by our proposed algorithm in Section IV-B. On the other hand, ${\cal P}_{\bf e}$ is reduced to the conventional BS multi-user beamforming problem w.r.t. ${\bf w}$ [27], which yields a closed-form optimal solution as briefly described below.
For $K=1$, we directly denote the single user in group $i\in{\cal G}$ as user $i$, with a total of $G$ users. The optimal beamforming solution ${\bf w}_{i}$ in (13) still holds as a special case, where the equivalent channel matrix $\widetilde{{\bf H}}_{i}$ for group $i$ is reduced to the equivalent channel $\tilde{{\bf h}}_{i}$ for user $i$, and weight vector ${\bf a}_{i}$ now becomes a scalar, denoted as $a_{i}$, i.e.,
${\bf w}_{i}=a_{i}{\bf R}^{-1}(\boldsymbol{\lambda})\tilde{{\bf h}}_{i},\quad i\in{\cal G}.$ (21)
The weights $\{a_{i}\}$ for all $G$ users can be derived in closed-form. This is because when the transmit beamformer can be individually designed for each user, at the optimality, all SINR constraints in (5b) is attained with equality. Specifically, let ${\bf c}_{i}\triangleq{\bf R}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf h}}_{i1}$, $i\in{\cal G}$, and we have the following the SINR equality constraint
$\frac{a_{i}^{2}}{\gamma_{i1}}|{\bf c}_{i}^{H}{\bf h}_{i1}|^{2}=\!\!\sum_{j\in{\cal G},j\neq i}\!\!a_{j}^{2}|{\bf c}_{j}^{H}{\bf h}_{j1}|^{2}+\sigma^{2},\quad\text{for}~{}i\in{\cal G}.$ (22)
Thus, $\{a_{i}^{2}\}$ can be obtained by solving the above $G$ linear equations. Define the $G\times G$ matrix ${\bf V}$ with its $ij$-th entry as
$[{\bf V}]_{ij}=\begin{cases}\frac{1}{\gamma_{i1}}|{\bf c}_{i}^{H}{\bf h}_{i1}|^{2},&j=i\\ -|{\bf c}_{i}^{H}{\bf h}_{j1}|^{2},&j\neq i.\end{cases}$ (23)
Define $\tilde{{\bf a}}=[a_{1}^{2},...,a_{G}^{2}]^{T}$. Then, the solution to the $G$ linear equations in (22) can be compactly written as
$\tilde{{\bf a}}=\sigma^{2}{\bf V}^{-1}\mathbf{1}$ (24)
where $\mathbf{1}$ is an all-one vector.
Finally, without the loss of optimality, we obtain
$a_{i}=\sqrt{[\tilde{{\bf a}}]_{i}},\quad i\in{\cal G}.$ (25)
Thus, the joint optimization of $({\bf w},{\bf e})$ for ${\cal P}_{o}$ by the AMBF approach involves alternating between computing the closed-form solution ${\bf w}_{i}$ in (21) and solving $\widetilde{{\cal S}}_{{\bf w}}$ for ${\bf e}$ using the updates in (17) and (18).
Remark 3.
RIS-assisted downlink multi-user beamforming design for the unicast scenario has been studied in many recent works. However, only a few of them considered the QoS problem for transmit power minimization [5, 21]. Perhaps partially this is due to the difficulty in using AO effectively to solve the joint BS and RIS beamforming problem. In [5], directly applying AO to the joint BS and RIS beamforming to minimize the BS transmit power is considered. As discussed at the beginning of Section III, direct AO renders the RIS beamforming a feasibility problem, which can not provide an effective solution. Thus, some modification to the feasibility problem is proposed to find a better solution, and a low-complexity heuristic method based on the sum of channel gains is also proposed. The joint beamforming design for RIS-assisted non-orthogonal multiple access (NOMA) networks is considered in [21], where the direct AO approach is also adopted, with the RIS beamforming being treated as a feasibility problem. Both [5] and [21] adopt SDP or SDR to solve the corresponding alternating subproblems. In our work, we first transform the QoS problem into ${\cal P}_{2}$ in (III-A). It allows us to break the problem into two subproblems, for which AO is effective in obtaining a good solution. The subproblems can be solved by the closed-form solution in (21) and the fast algorithm in Section IV-B, which is a scalable solution for large-scale systems.
VI RIS-Assisted Multicast MMF Problem
We now consider the weighted max-min fair (MMF) design for the RIS-assisted multicast scenario. The objective is to maximize the minimum weighted SINR among users, subject to the BS transmit power budget and the unit-modulus constraint for the RIS elements. The weighted SINR is the ratio between SINR and a pre-specified SINR target, which sets certain fairness among users. This joint optimization problem w.r.t. ${\bf w}$ and ${\bf e}$ is formulated as
${\cal S}_{o}:\max_{{\bf w},{\bf e}}$ $\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ (26a)
s.t. $\ \ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}\leq P,$ (26b)
$\ \ |e_{m}|^{2}=1,\ m\in{\cal M}.$ (26c)
where $P$ is the transmit power budget. Similar to the QoS problem ${\cal P}_{o}$, problem ${\cal S}_{o}$ is also non-convex and NP-hard.
For the conventional downlink multicast beamforming design at the BS, it is known that the QoS and MMF problems are inverse problems [28]: the role of the objective function and the constraints are switched in the two problems, and we can solve the MMF problem by iteratively solving the QoS problem along with a bi-section search over the minimum SINR target for the QoS problem until the total transmit power meets the power budget in the MMF problem. Naturally, we want to know whether such a relation also holds for the RIS-assisted multicast beamforming design, which now also includes the RIS passive beamformer optimization, in addition to the BS transmit beamformers. We show below that this is indeed the case.
Proposition 1.
For RIS-assisted multicast beamforming, the QoS problem ${\cal P}_{o}$ and the MMF problem ${\cal S}_{o}$ are inverse problems. In particular, we explicitly parameterize ${\cal S}_{o}$ as ${\cal S}_{o}(\boldsymbol{\gamma},P)$ for given $\boldsymbol{\gamma}$ and $P$ with the maximum objective value denoted as $t^{o}={\cal S}_{o}(\boldsymbol{\gamma},P)$, and we also parameterize ${\cal P}_{o}$ as ${\cal P}_{o}(\boldsymbol{\gamma})$ with $\boldsymbol{\gamma}$ being the SINR target vector, and the minimum power obtained as $P={\cal P}_{o}(\boldsymbol{\gamma})$. Then, ${\cal P}_{o}(\boldsymbol{\gamma})$ and ${\cal S}_{o}(\boldsymbol{\gamma},P)$ have the following relations:
$t^{o}$ $={\cal S}_{o}(\boldsymbol{\gamma},{\cal P}_{o}(t^{o}\boldsymbol{\gamma})),\quad P={\cal P}_{o}({\cal S}_{o}(\boldsymbol{\gamma},P)\boldsymbol{\gamma}).$ (27)
Proof:
See Appendix A.
Based on the above relation, we can solve the MMF problem ${\cal S}_{o}$ using Algorithm 1. In particular, by the relation in (27), ${\cal S}_{o}$ can be solved by solving its inverse QoS problem ${\cal P}_{o}$ iteratively along with a bi-section search for $t^{o}$. However, since the size of the RIS reflection elements is typically large $M\gg 1$, iteratively solving ${\cal P}_{o}$ will incur relatively high computational complexity. We need a more computationally efficient algorithm to solve the large-scale joint optimization problem ${\cal S}_{o}$. Aiming at this goal, we present a fast algorithm to solve ${\cal S}_{o}$ directly.
VI-A Proposed Fast Algorithm for Solving ${\cal S}_{o}$
For the conventional downlink scenario without RIS, as mentioned in Section IV-B, PSA is proposed in [12] to obtain a near-stationary solution for the multicast beamforming MMF problem. Below, we show that we can adopt PSA to solve the MMF problem efficiently for the RIS-assisted multicast scenario directly as well.
From Proposition 1, problems ${\cal P}_{o}$ and ${\cal S}_{o}$ are the inverse problems. This means the optimal structure of the BS multicast beamformer ${\bf w}$ still has the form given in (13). Indeed, with given ${\bf e}$, if we treat the RIS-assisted channel between the BS and user $k$ in group $i$ as the equivalent channel $\tilde{{\bf h}}_{ik}$, ${\cal S}_{o}$ reduces to the conventional multicast MMF problem, for which the optimal beamforming structure is (13) [11]. However, the challenge for the MMF problem is that computing $\boldsymbol{\lambda}$ in ${\bf R}(\boldsymbol{\lambda})$ is not straightforward. Fortunately, the asymptotic expression of ${\bf R}(\boldsymbol{\lambda})$ for large $N$ is obtained in closed form in [11], which can be used as an approximate expression of ${\bf R}(\boldsymbol{\lambda})$ to further simplify the computation. Specifically, let $\tilde{{\bf h}}_{iK}=\sqrt{\tilde{\beta}_{ik}}{\bf g}_{ik}$ where $\tilde{\beta}_{ik}$ represents the large-scale channel variance and $\tilde{{\bf g}}_{ik}\sim{\cal C}{\cal N}(\mathbf{0},{\bf I})$. Then, the asymptotic expression of ${\bf R}(\boldsymbol{\lambda})$ has the following simple closed form
${\bf R}(\boldsymbol{\lambda})\approx{\bf I}+\frac{P\bar{\beta}_{{\bf h}}}{\sigma^{2}K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}{\bf g}_{ik}{\bf g}_{ik}^{H}\triangleq\widetilde{{\bf R}}$ (28)
where $\bar{\beta}_{{\bf h}}\triangleq 1/\Big{(}\frac{1}{K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}\frac{1}{\tilde{\beta}_{ik}}\Big{)}$ is the harmonic mean of the large-scale channel variance of all users. Using $\widetilde{{\bf R}}$ in (28), we can replace ${\bf w}_{i}$ in (13) with the following expression:
${\bf w}_{i}\,\approx\,\widetilde{{\bf R}}^{-1}\widetilde{{\bf H}}_{i}{\bf a}_{i}\,,\;i\in{\cal G}.$ (29)
Below, we develop our fast algorithm to solve ${\cal S}_{o}$ based on ${\bf w}_{i}$ in (29) to reduce the computational complexity.
Using ${\bf w}_{i}$ in (29), we first convert ${\cal S}_{o}$ into a joint optimization problem w.r.t. $\{{\bf a}_{i}\}$ and ${\bf e}$, given by
${\cal S}_{o}^{\text{eq}}:\max_{{\bf a},{\bf e}}$ $\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ (30a)
s.t. $\ \ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P,$ (30b)
$\ \ |e_{m}|^{2}\leq 1,\ m\in{\cal M}$ (30c)
$\ \ {\bf e}^{T}{\bf e}=M.$ (30d)
where ${\bf a}\triangleq[{\bf a}_{1}^{H},\cdots,{\bf a}_{G}^{H}]^{H}$, $\widetilde{{\bf C}}_{i}\triangleq\widetilde{{\bf R}}^{-1}\widetilde{{\bf H}}_{i}$, and the constraints in (26c) are replaced by the equivalent set of constraints in (30c) and (30d) based on (9). Similar to $\widetilde{{\cal S}}_{{\bf w}}$ in (III-B2), we relax $S_{o}^{\text{eq}}$ into the following problem by moving the equality constraint in (30d) into the objective function as a penalty term with a penalty weight $\delta>0$:
$\widetilde{{\cal S}}_{o}:\max_{{\bf a},{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\delta\frac{{\bf e}^{H}{\bf e}}{M}$ (31a)
s.t. $\ \ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P,$ (31b)
$\ \ |e_{m}|^{2}\leq 1,\ m\in{\cal M}.$ (31c)
Problem $\widetilde{{\cal S}}_{o}$ is similar to $\widetilde{{\cal S}}_{\bf w}$ in (III-B2), except that $\widetilde{{\cal S}}_{o}$ is a joint optimization problem for $({\bf a},{\bf e})$ while $\widetilde{{\cal S}}_{\bf w}$ is for ${\bf e}$ only. Nonetheless, we can still apply PSA to compute a near-stationary joint solution $({\bf a},{\bf e})$ for $\widetilde{{\cal S}}_{o}$ efficiently, similar to Section IV-B, as shown below.
VI-A1 Problem Reformulation
Let ${\bf x}\triangleq[{\bf a}^{H},{\bf e}^{H}]^{H}$. By substituting the SINR expression in (2) into the objective function in (31a), we define
$\varphi_{ik}({\bf x})\triangleq$ $-\!\frac{1}{\gamma_{ik}}\frac{|{\bf a}_{i}^{H}\widetilde{{\bf C}}_{i}^{H}\!({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})|^{2}}{\sum_{j\neq i}|{\bf a}_{j}^{H}\widetilde{{\bf C}}_{j}^{H}\!({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})|^{2}\!+\!\sigma^{2}}-\delta\frac{{\bf e}^{H}{\bf e}}{M},$ (32)
for $k\in{\cal K}_{i}$, $i\in{\cal G}$. Following the similar arguments from (14)–(16), let $\boldsymbol{\varphi}({\bf x})$ be a $K_{\text{tot}}\times 1$ vector containing all $\varphi_{ik}({\bf x})$’s, and let ${\bf y}$ denote a $K_{\text{tot}}\times 1$ vector with ${\bf y}\succcurlyeq 0$ and ${\bf 1}^{T}{\bf y}=1$. Then, $\widetilde{{\cal S}}_{o}$ can be transform into the following equivalent problem
$\widetilde{{\cal S}}_{o}^{\prime}:\min_{{\bf x}\in{\cal U}}\max_{{\bf y}\in{\cal Y}}$ $\ g({\bf x},{\bf y})$ (33)
where $g({\bf x},{\bf y})\triangleq\boldsymbol{\varphi}^{T}({\bf x}){\bf y}$, ${\cal Y}\triangleq\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$, and ${\cal U}\triangleq\{{\bf x}:|e_{m}|^{2}\leq 1,\ m\in{\cal M};\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\}$. Both ${\cal Y}$ and ${\cal U}$ are compact convex sets.
VI-A2 The Projected Subgradient Algorithm
Problem $\widetilde{{\cal S}}_{o}^{\prime}$ is a non-convex-concave min-max problem, whose structure is the same as $\widetilde{{\cal S}}_{\bf w}^{\prime}$ in (16). Thus, we can apply PSA with the updating steps similar to (17) and (18). In particular, the updating procedure at iteration $j$ is given by
${\bf y}^{(j)}\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}g({\bf x}^{(j)},{\bf y});$ (34)
${\bf x}^{(j+1)}=\Pi_{{\cal U}}\big{(}{\bf x}^{(j)}-\eta\nabla_{{\bf x}}g({\bf x}^{(j)},{\bf y}^{(j)})\big{)}$ (35)
where $\eta>0$ is the step size, and $\Pi_{{\cal U}}({\bf x})$ is the projection of ${\bf x}$ onto the feasible set ${\cal U}$. Since the constraints on ${\bf a}$ and ${\bf e}$ in ${\cal U}$ are separate, $\Pi_{{\cal U}}({\bf x})$ can be performed for ${\bf a}$ and ${\bf e}$ separately:
$\Pi_{{\cal U}}({\bf x})=\begin{cases}{\bf a}&\mbox{if }\sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\\ \sqrt{\frac{P}{P_{\text{tot}}}}{\bf a}&\text{o.w.}\\ e_{m}&\mbox{if }|e_{m}|\leq 1,\ m\in{\cal M}\\ \frac{e_{m}}{|e_{m}|}&\text{o.w.}\end{cases}.$ (36)
Since ${\cal Y}$ is a probability simplex, an optimal solution to the maximization problem in (34) is ${\bf y}=[0,\cdots,1,\cdots,0]^{T}$, with 1 at the position that corresponds to $\max_{i,k}\varphi_{ik}({\bf e})$, i.e.,
$g({\bf x},{\bf y}^{(j)})=\varphi_{\hat{i}\hat{k}}({\bf x})=\max_{i,k}\varphi_{ik}({\bf x}),$ (37)
for some $(\hat{i},\hat{k})$. Then, the gradient $\nabla_{{\bf x}}g({\bf x},{\bf y}^{(j)})$ is given by
$\nabla_{{\bf x}}g({\bf x},{\bf y}^{(j)})$ $=\nabla_{{\bf x}}\varphi_{\hat{i}\hat{k}}({\bf x})=\Bigl{[}\nabla_{{\bf a}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})\!,\nabla_{{\bf e}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})\Bigl{]}^{H}.$ (38)
For $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$, note that the expression for $\varphi_{ik}({\bf x})$ is the same as $\phi_{ik}({\bf e})$. Thus, the gradient $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ can be computed using (IV-B2), except that ${\bf w}_{i}$ is computed using (29), and $\zeta$ is replaced by $\delta$.
To compute $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$, we rewrite $\varphi_{ik}({\bf x})$ in (32) as
$\varphi_{ik}({\bf x})=$ $-\frac{1}{\gamma_{ik}}\frac{{\bf a}_{i}^{H}\widetilde{{\bf A}}_{iik}{\bf a}_{i}}{\sum_{j\neq i}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{jik}{\bf a}_{j}+\sigma^{2}}-\delta\frac{{\bf e}^{H}{\bf e}}{M},$
where $\widetilde{{\bf A}}_{jik}\triangleq\widetilde{{\bf C}}_{j}^{H}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}^{H}\widetilde{{\bf C}}_{j}$, with $\tilde{{\bf h}}_{ik}={\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e}$ being the equivalent channel. Then, we have
$\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$ $=[\nabla_{{\bf a}_{1}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x}),\cdots,\nabla_{{\bf a}_{G}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})]^{H}$ (39)
where
$\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})\!=\!\begin{cases}\displaystyle-\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}}}{\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}}&\mbox{if }i=\hat{i}\\[15.00002pt] \displaystyle\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{({\bf a}_{\hat{i}}^{H}\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}})\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}}{\bigl{(}\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}\bigl{)}^{2}}&\text{o.w.}\end{cases}$ (40)
From the above, we note that the updates in (34) and (35) can be computed using closed-form evaluations.
VI-A3 Final Processing
Since we are solving $\widetilde{{\cal S}}_{o}$, which is a relaxed problem of ${\cal S}_{o}$, the computed solution ${\bf e}^{\star}$ may not satisfy (26c) for ${\cal S}_{o}$. Thus, we require this final processing step to obtain a feasible solution to ${\cal S}_{o}$. Let ${\bf x}^{\star}=[{{\bf a}^{\star}}^{H},{{\bf e}^{\star}}^{H}]^{H}$ be the solution produced by the updating procedure in (34) and (35). If ${\bf e}^{\star}$ does not satisfy (26c), we have the following step to obtain a feasible solution based on $({\bf a}^{\star},{\bf e}^{\star})$:
  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which takes the phase component of each element in ${\bf e}^{\star}$;
  • Given ${\bf e}^{\text{final}}$, we solve ${\cal S}_{o}^{\text{eq}}$ in (VI-A) w.r.t. ${\bf a}$ and obtain the final beamforming solution ${\bf w}^{\text{final}}$. Note that in this case, ${\cal S}_{o}^{\text{eq}}$ is reduced to the following conventional MMF problem:
    ${\cal S}^{\text{eq}}({\bf e}^{\text{final}}):\max_{{\bf a}}$ $\min_{i,k}\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ s.t. $\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P.$ (41)
    where $\textrm{SINR}_{ik}$ is given in (12) with the equivalent channel $\tilde{{\bf h}}_{ik}$ computed based on ${\bf e}^{\text{final}}$. As mentioned earlier, the PSA-based fast algorithm for this MMF problem is provided in [12]. We directly apply this algorithm to compute the solution ${\bf a}^{\text{final}}$ to ${\cal S}^{\text{eq}}({\bf e}^{\text{final}})$. For initialization in this algorithm, we use ${\bf a}^{\star}$ as a good initial point to this algorithm to achieve fast convergence and good performance.
We summarize our proposed fast algorithm for the MMF problem ${\cal S}_{o}$ in Algorithm 2. Again, based on the existing convergence analysis [12], the PSA updating procedure in (34) and (35) is guaranteed to converge in finite time to a point at the vicinity of a stationary point for $\widetilde{{\cal S}}_{o}$.
Algorithm 2 The Proposed Fast Algorithm for RIS-Assisted MMF Problem ${\cal S}_{o}$.
1:Initialization: Set feasible point ${\bf x}^{(0)}$; Set $n=0$.
2:repeat
3:     Update ${\bf x}^{(n)}$ using (34) and (35) to obtain ${\bf x}^{(n+1)}$.
4:     Set $n\leftarrow n+1$.
5:until convergence
6:Set ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{(n)})$.
7:Solving ${\cal S}({\bf e}^{\text{final}})$ to obtain ${\bf w}^{\text{final}}$.
8:return (${\bf w}^{\text{final}}$, ${\bf e}^{\text{final}}$).
VI-B Discussions
VI-B1 Initialization
The proposed updating procedure in (34) and (35) requires an initial point ${\bf x}^{(0)}=[{{\bf a}^{(0)}}^{H},{{\bf e}^{(0)}}^{H}]^{H}$. This initial point can be generated randomly, as the projection (35) will ensure that the subsequent points $\{{\bf x}^{(n)}\}$ are feasible. However, a good initial point improves the rate of convergence. Thus, instead of a random initial point, we consider the following approach.
We first generate ${\bf e}^{(0)}$ with a random phase for each element. Given ${\bf e}^{(0)}$, ${\cal S}_{o}^{\text{eq}}$ in (VI-A) is reduced to the MMF problem w.r.t. ${\bf a}$, which is similar to ${\cal S}^{\text{eq}}({\bf e}^{\text{final}})$ in (41). It can be further equivalently written as
${\cal S}^{\text{eq}}({\bf e}^{(0)}):$ $\max_{{\bf a},t}\ t$ s.t. $\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq t,\ k\in{\cal K}_{i},i\in{\cal G}$ (42a)
$\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P.$ (42b)
Note that any feasible solution to the above problem is feasible to $\widetilde{{\cal S}}_{o}$. Thus, we generate ${\bf a}^{(0)}$ in an attempt to solve this MMF problem, which is to solve its inverse problem, i.e., the QoS problem similar to ${\cal P}_{\bf e}$ in (III-B1) for given $t$, along with bi-section search over $t$ [28, 11]. However, since our goal is only to generate a feasible point ${\bf a}^{(0)}$ for $\widetilde{{\cal S}}_{o}$, we only need to solve the QoS problem [^2] To solve the QoS problem, we use the classical approach of SDR with the Gaussian randomization procedure. along with one round of bi-section search over $t$ to obtain the initial point ${\bf a}^{(0)}$.
VI-B2 Computational Complexity
For the updates in (34) and (35) at each iteration, the update ${\bf y}^{(j)}$ is obtained in (37) by finding the maximum $\varphi_{ik}({\bf e})$ among $K_{\text{tot}}$ users, which requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+2N+4)$ flops. The update in (35) requires to calculate $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$ and $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$. Note that $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ is computed using (IV-B2), which requires $2GM+2M$ flops as discussed in Section IV-C2.ii). To compute $\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})$ in (40), we need ${\bf a}_{j}^{H}\widetilde{{\bf C}}_{j}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}$ for $j\in{\cal G}$, which is part of the SINR expression that has already been computed in $\varphi_{ik}({\bf e})$. We also need to calculate $\widetilde{{\bf C}}_{i}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}$ for $\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}=\widetilde{{\bf C}}_{i}^{H}\tilde{{\bf h}}_{\hat{i}\hat{k}}\tilde{{\bf h}}_{\hat{i}\hat{k}}^{H}\widetilde{{\bf C}}_{i}{\bf a}_{i}$, for $i\in{\cal G}$. Note that $\widetilde{{\bf C}}_{i}$’s only need to be computed once before the update procedure, and $\tilde{{\bf h}}_{\hat{i}\hat{k}}$ is already computed as part of the SINR expression. Thus, compute $\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}$’s for $i\in{\cal G}$ require $K_{\text{tot}}(N+1)$. Lastly, we need $K_{\text{tot}}+G$ flops to calculate the final value of $\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})$ for all ${\bf a}_{i}$’s. In summary, the complexity in obtaining the update in (35) requires $2(G+1)M+K_{\text{tot}}(N+2)+G$ flops. Thus, the overall leading complexity of the updating procedure in each PSA iteration is dominated by computing ${\bf y}^{(j)}$ in (37), which is $GK_{\text{tot}}N+K_{\text{tot}}MN$ flops. It grows linearly with $M$ and $N$.
VI-C General RIS-Assisted Multi-User Downlink Beamforming
Similar to the discussion in Section V about the generalization to downlink RIS-assisted multi-user beamforming for the QoS power minimization problem, our proposed fast algorithm (Algorithm 2) is directly applicable to the general downlink RIS-assisted multi-user beamforming problem with the max-min fair objective for both unicast and multicast scenarios. Specific to the unicast case with $K_{i}=1,\forall i$, the computational complexity per iteration in the proposed iterative algorithm is reduced to $GMN+G^{2}N$ flops, which is computationally cheap.
For the multi-user unicast scenario, the MMF problem for RIS-assisted downlink beamforming under various system settings has been considered in the recent literature, and various approaches have been proposed [22, 23, 24]. In [22], for the same setting considered in this work, an iterative algorithm that provides an alternating descent path is proposed to maximize the minimum SINR, and the computational complexity in each iteration as $O((GN)^{3}+M^{4})$. In comparison, the computational complexity of our proposed fast algorithm only grows linearly with $G$, $M$, and $N$. A different system model for RIS-assisted transmission is considered in [23], where only the RIS path is available and the channel correlation matrix between RIS and users are assumed to be the same. Using an approximation of the optimal BS transmit beamformer, the authors propose an algorithm based on projected gradient ascent to obtain the RIS beamformer based on the large-scale channel statistics. The computational complexity of the proposed algorithm is in the higher polynomial order of $N$ and $M$. RIS-assisted multi-cell coordinated transmit beamforming is considered in [24], where the RIS can adjust both reflecting phases and amplitudes. The AO approach is adopted, with two subproblems being solved by the second-order cone programming (SOCP) and SDR. An inexact AO approach with SCA is further considered to reduce the computational complexity. However, since these proposed approaches are second-order algorithms, their computational complexity is still high.
VII Simulation Results
We consider a downlink RIS-assisted transmission scenario with a BS and an RIS. The locations of the BS and the RIS using the $(x,y,z)$-coordinates in meters are $(0,0,0)$ and $(70,70,0)$, respectively. The RIS is a rectangular surface with the array elements placed on the $(y,z)$-plane. The users are located randomly within a circle centered at the RIS with a radius of $20$ m on the $(x,y)$-plane. We set the target SINR $\gamma_{ik}=\gamma$, $\forall i,k$. Unless specified otherwise, our default system setup is given as $N=16$ antennas at the BS, and $M=10\times 10$ reflective elements at the RIS, $G=2$ groups, $K_{i}=K=2$ users, $\forall i$.
We assume a line-of-sight (LOS) channel between the BS and the RIS and model ${\bf H}_{\text{r}}$ as a Rician fading channel matrix given by ${\bf H}_{\text{r}}=\beta^{\text{BR}}\big{(}\sqrt{\frac{K_{\text{r}}}{1+K_{\text{r}}}}{\bf H}_{\text{r}}^{\text{LOS}}+\sqrt{\frac{1}{1+K_{\text{r}}}}{\bf H}_{\text{r}}^{\text{NLOS}}\big{)}$, where Rician factor $K_{\text{r}}=10$, $\beta^{\text{BR}}$ is the path gain modeled as $\beta^{\text{BR}}\text{[dB]}=-30-22\log 10(d^{\text{B-R}})$ with $d^{\text{B-R}}$ being the BS-RIS distance in meters, ${\bf H}_{\text{r}}^{\text{NLOS}}$ is the non-line-of-sight (NLOS) component modeled as $[{\bf H}_{\text{r}}^{\text{NLOS}}]_{ij}\sim{\cal C}{\cal N}(0,{\bf I})$, and ${\bf H}_{\text{r}}^{\text{LOS}}$ is the LOS component. The LOS component ${\bf H}_{\text{r}}^{\text{LOS}}$ is a function of the BS and RIS locations and is modeled as
${\bf H}_{\text{r}}^{\text{LOS}}={\bf b}_{\text{RIS}}(\psi_{2},\theta_{2}){\bf b}_{\text{BS}}(\psi_{1},\theta_{1})^{H}$ (43)
where $\psi_{1}$ and $\theta_{1}$ are the azimuth and elevation angles of departure (AoD) from the BS to the IRS, and ${\bf b}_{\text{BS}}(\psi_{1},\theta_{1})$ is the BS antenna array steering vector, with its $n$th entry as
$[{\bf b}_{\text{BS}}(\psi_{1},\theta_{1})]_{n}=e^{\frac{2\pi(n-1)\Delta_{\text{BS}}}{\lambda_{c}}}\cos(\psi_{1})\cos(\theta_{1}),\ n=1,\ldots,N$
where $\Delta_{\text{BS}}$ is the distance between two adjacent BS antennas, $\lambda_{c}$ is the carrier wavelength, and $\cos(\psi_{1})\cos(\theta_{1})=\frac{x_{\text{RIS}}-x_{\text{BS}}}{d^{\text{B-R}}}$. We assume $\Delta_{\text{BS}}=\frac{\lambda_{c}}{2}$. Similarly, $\psi_{2}$ and $\theta_{2}$ are the azimuth and elevation angles of arrival (AoA) from the BS to the RIS, and ${\bf b}_{\text{RIS}}(\psi_{2},\theta_{2})$ is the RIS element streaming vector, with the $m$th entry, for $m\in{\cal M}$, given by
$[{\bf b}_{\text{RIS}}(\psi_{2},\theta_{2})]_{m}=e^{\frac{2\pi\Delta_{\text{RIS}}}{\lambda_{c}}}[y_{m}\sin(\psi_{2})\cos(\theta_{2})+z_{m}\sin(\theta_{2})]$
where $\Delta_{\text{RIS}}$ is the distance between two adjacent RIS elements, $y_{m}=\text{mod}(m-1,M_{y})$ with $M_{y}$ being the number of RIS elements along the $y$-axis, $z_{m}=\lfloor m/M_{y}\rfloor+1$, $\sin(\psi_{2})\cos(\theta_{2})=\frac{y_{\text{BS}}-y_{\text{RIS}}}{d_{\text{B-R}}}$, and $\sin(\theta_{2})=\frac{z_{\text{BS}}-z_{\text{RIS}}}{d_{\text{B-R}}}$. We assume $\Delta_{\text{RIS}}=\frac{\lambda_{c}}{2}$.
We consider a NLOS channel between the RIS and each user modeled as Rayleigh fading. The channel ${\bf h}^{\text{r}}_{ik}$ between the RIS and user $k$ in group $i$ is generated i.i.d as ${\bf h}^{\text{r}}_{ik}\sim{\cal C}{\cal N}({\bf 0},\beta^{\text{r}}_{ik}{\bf I})$, where $\beta^{\text{r}}_{ik}$ is the path gain modeled as $\beta^{\text{r}}_{ik}[\text{dB}]=-30-22\log 10(d^{\text{R-U}}_{ik})$, with $d^{\text{R-U}}_{ik}$ being the distance between the RIS and the user. We also model the channel between BS and user $k$ in group $i$ as Rayleigh fading, which is generated i.i.d as ${\bf h}^{\text{d}}_{ik}\sim{\cal C}{\cal N}({\bf 0},\beta^{\text{d}}_{ik}{\bf I})$, where $\beta^{\text{d}}_{ik}$ is the path gain modeled as $\beta^{\text{d}}_{ik}[\text{dB}]=-32.6-36.7\log 10(d^{\text{B-U}}_{ik})$, with $d^{\text{B-U}}_{ik}$ being the respective BS-user distances. The receiver noise power is set to $\sigma^{2}=-100~{}\text{dBm}$.
[…]
VII-A Convergence Behavior of the Proposed Algorithms
We first study the convergence behavior of our proposed AMBF algorithm in Algorithm 1 for the QoS problem ${\cal P}_{o}$, which alternatingly solves two multicast subproblems at the BS and the RIS using our proposed fast algorithms. Fig. 0(a) shows the convergence behavior of solving subproblem $\widetilde{{\cal S}}_{{\bf w}}$ using PSA in a specific iteration of the AMBF algorithm. The convergence behavior for three random channel realizations are shown. We see PSA converges after $1000\sim 2000$ iterations. Note that the closed-form updates for PSA in each iteration is computationally cheap, and thus, the overall computation time for $1000\sim 2000$ iterations is fast. Furthermore, denote the total transmit power at the BS as $P_{\text{tot}}=\sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$. Fig. 0(b) shows the trajectory of $P_{\text{tot}}$ over iterations by Algorithm 1 for three random channel realizations. We observe that our proposed AMBF approach converges within about $6\sim 12$ iterations, which is relatively fast, especially given that we use fast algorithms for computing the solution to each subproblem.
(a) The subproblem $\widetilde{{\cal S}}_{{\bf w}}$.
(b) The overall AMBF Algorithm.
Figure 1: Convergence behavior of the AMBF algorithm in Algorithm 1 ($N=16,M=100$, $(G,K)=(2,2)$, $\gamma=10$ dB).
Fig. 2 shows the convergence behavior of the MMF problem ${\cal S}_{o}$ solved by Algorithm 2 under three random channel realizations. We see that for two channel realizations, Algorithm 2 converges very fast within just a few iterations. Even for a worse scenario under the third channel realization, Algorithm 2 converges after $800$ iterations. Note that the PSA updates in each iteration is closed-form and computationally cheap.
Figure 2: Convergence behavior of the Algorithm 2 ($N=30$, $M=100$, $(G,K)=(2,2)$, $\gamma=10$ dB, $P=10$ dB).
VII-B Performance Comparison for the QoS Problem
We now evaluate the performance of our proposed AMBF algorithm in Algorithm 1. For performance comparison, we consider the following methods:
  1. i)
    No RIS: A conventional downlink multicast scenario without the RIS.
  2. ii)
    Random RIS: Apply random phase-shift for the RIS elements in ${\bf e}$. With given ${\bf e}$, solve the BS multicast beamforming problem ${\cal P}_{\bf e}$ for ${\bf w}$.
  3. iii)
    Direct AO [19]: Apply the AO approach directly to the QoS problem ${\cal P}_{o}$, where the subproblem w.r.t. ${\bf e}$ is a feasibility problem with the SINR constraints in (3b) and the unit-modulus constraints in (3c).
  4. iv)
    AMBF-SDR: Apply Algorithm 1, except that instead of solving subproblem $\tilde{{\cal S}}_{\bf w}$ w.r.t. ${\bf e}$, we relax ${\cal S}_{{\bf w}}^{\text{eq}}$ by dropping the constraint in (10c) and solve the relaxed problem:
    ${\cal S}_{{\bf w}}^{\text{Relaxed}}:\max_{{\bf e}}\min_{i,k}$ $\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ s.t. $|e_{m}|^{2}\leq 1,m\in{\cal M},$
    We solve ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ by using its epigraph form and applying SDR with a bi-section search for the objective value (i.e., let ${\bf X}={\bf e}{\bf e}^{H}$ and drop the rank-1 condition). Then, we recover ${\bf e}$ from ${\bf X}$ using the Gaussian randomization method.
  5. v)
    AMBF-SCA: Similar to AMBF-SDR that apply Algorithm 1, except that we solve ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ by SCA with bi-section search over $t$. In particular, we convexify ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ into a sequence of convex approximation problems and solved them using a standard convex solver.
  6. vi)
    IP Method: Apply MATLAB non-linear optimization solver Fmincon to ${\cal P}_{o}$. It uses the interior-point (IP) algorithm to solve ${\cal P}_{o}$.
Figure 3: Total transmit power $P_{\text{tot}}$ vs. $M$ ($N=16$, $(G,K)=(2,2)$, $\gamma=10$ dB).
TABLE I: Average Computation Time (sec).
$M$
25
50
100
200
300
RIS-SDR
24.73
46.95
84.68
672
1743
RIS-SCA
23.12
38.52
62.17
436
1349
IP Method
3.25
10.73
58.33
796
2062
Algorithm 1
2.19
3.12
5.22
8.83
13.62
VII-B1 Performance over the Number of RIS elements $M$
Fig. 3 shows the total BS transmit power $P_{\text{tot}}$ vs. the number of RIS elements $M$ by different methods, for the target SINR $\gamma=10$ dB. We see that the BS transmit power for random RIS, direct AO, and IP Method remain roughly flat as $M$ increases. This indicates that they do not provide an effective RIS beamforming solution. In contrast, under Algorithm 1, the resulting BS transmit power is reduced substantially by $4$ dB as $M$ increases from 25 to 300. RIS-SCA has a nearly identical performance to Algorithm 1, while RIS-SDR performs worse than Algorithm 1 with a noticeable gap as large as $2$ dB. This is expected, as SDR is an approximation method with a degraded performance when there are many constraints, and the problem size is relatively large.
We also compare the average computation time of different algorithms in Table I. The computational advantage of Algorithm 1 over RIS-SDR, RIS-SCA, and IP Method as the number of RIS elements $M$ increases is clearly observed, demonstrating the scalability of Algorithm 1 over $M$. In particular, the computational complexity of Algorithm 1 only grows mildly with $M$. Thus, Algorithm 1 provides both the best performance and the fastest computational time than the other three methods, especially for the large value of $M$.
VII-B2 Performance over Target SINR $\gamma$
Fig. 4 shows $P_{\text{tot}}$ vs. target SINR $\gamma$ by different methods. We see that direct AO, Random RIS, and IP method provide either no or only very small improvement over the case of No RIS. A substantial gain can be achieved when the RIS passive beamforming is effectively optimized to improve the channel environment between the BS and users. In particular, our proposed Algorithm 1 provides about $4$ dB power reduction over No RIS for $\gamma\geq 4$ dB. Similar to Fig. 3, RIS-SCA has nearly identical performance to Algorithm 1, while RIS-SDR performs worse than Algorithm 1 with a noticeable $2\sim 3$ dB gap.
Figure 4: Total transmit power $P_{\text{tot}}$ vs. target SINR $\gamma$ ($N=16,M=100$, $(G,K)=(2,2)$).
VII-C Performance Comparison for the MMF Problem
We now evaluate the proposed RIS-assisted multicast beamforming design for the MMF problem. For performance comparison, besides our proposed Algorithm 2, we consider the following methods:
  1. 1.
    No RIS: A conventional downlink multi-group multicast scenario without RIS.
  2. 2.
    Random RIS: Apply random phase-shift for the RIS elements in ${\bf e}$. With given ${\bf e}$, solve the BS multicast beamforming MMF problem for ${\bf w}$.
  3. 3.
    Relaxed RIS-MMF: Only solve the relaxed problem $\widetilde{{\cal S}}_{o}$ in (VI-A), i.e., Algorithm 2 without the final processing step to obtain the feasible solution ${\bf e}$. We use this method to measure the performance loss in the final processing step in Algorithm 2 and quantify how good the relaxed problem $\widetilde{{\cal S}}_{o}$ is as an approximation of the original problem ${\cal S}_{o}$.
  4. 4.
    IP Method: Use MATLAB non-linear optimization solver Fmincon to solve the MMF problem ${\cal S}_{o}$.
  5. 5.
    The alternating projected gradient (APG) method [18]: APG is a recently proposed iterative algorithm to maximize the achievable sum group-rate of all groups. The group rate is defined as the minimum rate among users in the group, which corresponds to the minimum SINR in the group.
Note that the design objective for the APG method [18] is different from the max-min SINR (or rate) objective in our MMF problem. However, for the single-group multicast scenario when $G=1$, the two objectives becomes identical. Since there are no existing RIS-assisted methods designed for the MMF problem, we compare Algorithm 2 with APG in this special single-group case. Besides [18], we note that [17] has also proposed the RIS-assisted multicast beamforming algorithms for the sum-group-rate maximization. Since it is shown in [18] that APG outperforms the method in [17] in both performance and computational complexity, we choose APG as the state-of-the-art method for comparison.
In the following simulation results, we set the BS transmit power $P=10~{}{\text{dBm}}$.
VII-C1 Multiple Groups
Fig. 6 shows the average minimum SINR vs. the number of RIS elements $M$ by different methods for $N=30$. We see that the minimum SINR achieved by Algorithm 2 increases noticeably as $M$ increases, significantly better than random RIS and IP Method, especially when $M>N$. This demonstrates the effectiveness of Algorithm 2 to design RIS beamforming to improve the user channel conditions. Furthermore, we see that Algorithm 2 and Relaxed RIS-MMF provide a nearly identical performance for all values of $M$. This indicates that the solution to $\widetilde{{\cal S}}_{o}$ in Algorithm 2 is mostly feasible to ${\cal S}_{o}$, and the final processing step causes a negligible performance loss. This also demonstrates that the relaxed problem $\widetilde{{\cal S}}_{o}$ is a good approximation of the original problem ${\cal S}_{o}$.
Figure 5: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(2,2)$).
Figure 5: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(2,2)$).
Figure 6: Average minimum SINR vs. $N$ ($M=100$, $(G,K)=(2,2)$).
Fig. 6 shows the average minimum SINR vs. the number of BS antennas $N$ by different methods for $M=100$. Again, we see that the gain of random RIS and IP Method over no RIS is very small. Our Algorithm 2 provides about $1.5\sim 2$ dB additional SINR gain over these methods for different values of $N$. In Fig. 8, we show the average minimum SINR vs. the number of users $K$ per group for $N=30$ and $M=100$. Algorithm 2 outperforms no RIS, random RIS, and IP Method for all values of $K$, especially for smaller values of $K$. This is expected, since the gain of the optimized multicast beamforming is higher for a smaller user group. The gain reduces as the group size increases, and the difference among algorithms reduces.
Figure 7: Average minimum SINR vs. $K$ ($N=30,M=100$, $(G,K)=(2,2)$).
Figure 7: Average minimum SINR vs. $K$ ($N=30,M=100$, $(G,K)=(2,2)$).
Figure 8: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(1,4)$).
VII-C2 Single Group
We now compare our proposed algorithm with other algorithms, including the APG method [18], in the single group scenario. Fig. 8 shows the average minimum SINR vs. the number of RIS elements $M$ for $G=1$, $K=4$, and $N=30$. We see that Algorithm 2 outperforms all other algorithms significantly. In particular, our proposed algorithm provides $1\sim 2.5$ dB gain over APG for $M$ ranging from $25$ $\sim 300$.
We also compare the average computation time of different methods in Table II for the setting used in Fig. 8. IP method has the highest computational complexity. Algorithm 2 and APG have a similar computation time as $M$ increases. Although APG is slightly faster than Algorithm 2, the difference is not significant.
TABLE II: MMF – Average Computation Time (sec).
$M$
25
50
100
200
300
APG [18]
0.29
0.71
1.61
2.05
2.7
IP Method
8.73
12.17
36.63
197
476
Algorithm 2
0.45
0.98
2.1
2.6
3.2
VII-C3 Unicast Scenario
We also consider the RIS-assisted downlink beamforming in a unicast scenario with $4$ users by setting $G=4$ and $K=1$. Note that in this case, Algorithm 2 and APG have different design objectives: APG aims to maximize the sum rate of all users, while Algorithm 2 aims to maximize the minimum SINR among users. For comparison, we plot the average minimum SINR and the sum rate performance over $M$ in Figs. 10 and 10, respectively, for $N=30$. We see that Algorithm 2 outperforms the APG method and other methods for both the minimum SINR and the sum rate objectives.
Figure 9: Unicast: Average minimum SINR vs. $M$ ($N=30$, $G=4$ users).
Figure 9: Unicast: Average minimum SINR vs. $M$ ($N=30$, $G=4$ users).
Figure 10: Unicast: Average sum rate vs. $M$ ($N=30$, $G=4$ users).
VIII Conclusion
In this paper, we consider joint BS and RIS beamforming design for RIS-assisted downlink multi-user beamforming, aiming to provide efficient and scalable algorithms. We have shown that the joint BS and RIS beamforming optimization for the QoS problem can be broken down into a BS multicast beamforming QoS problem and an RIS passive multicast beamforming MMF problem, and we have proposed an AMBF approach to solve the two subproblems alternatingly. We have utilized the optimal multicast beamforming structure to solve the BS multicast beamforming QoS problem efficiently. We have also developed a PSA-based fast iterative algorithm to solve the challenging RIS passive multicast MMF problem, which only contains closed-form updates with cheap computation. We have shown that our AMBF approach is also applicable to the unicast scenario, making it an effective method for general RIS-assisted downlink beamforming. For the RIS-assisted MMF problem, based on the relation of the QoS and MMF problems, we have used the optimal BS beamforming structure again to transform the MMF problem and propose a PSA-based iterative algorithm. Similar to that for the QoS problem, The algorithm uses the closed-form updates per iteration, leading to a highly computationally efficient solution. Simulation results show that our proposed algorithms outperform the alternative methods and that the computational complexity is substantially lower than most existing methods and grows linearly with the number of RIS reflective elements and BS antennas.
Appendix A Proof of Proposition 1
Proof:
We can equivalently rewrite ${\cal S}_{o}$ as
$\max_{{\bf w}}$ $\ \bigg{(}\max_{{\bf e}:|e_{m}|^{2}=1,m\in{\cal M}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}$ s.t. $\ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}\leq P.$
For the QoS problem ${\cal P}_{o}$, consider its equivalent formulation ${\cal P}_{2}$ in (III-A) derived in Section III-A. Comparing the above problem and ${\cal P}_{2}$, we see that the objective functions and the constraints in these two problems have the same forms but switched, indicating that they are the inverse problems (and equivalently ${\cal S}_{o}$ and ${\cal P}_{o}$ are the inverse problems). It is straightforward to see that, at the optimality, the respective constraints in ${\cal S}_{o}^{\prime}$ and ${\cal P}_{2}$ are attained with equality. As a result, we have the relations in (27). ∎

### v3

@@
-Efficient Design for Multi-user Downlink Beamforming with Reconfigurable Intelligent Surface
+Fast and Scalable Beamforming for RIS-Assisted Downlink Multi-group Multicasting
 Abstract
-This paper considers downlink multi-user transmission facilitated by a reconfigurable intelligent surface (RIS).
-First, focusing on the multi-group multicast beamforming scenario, we develop a fast and scalable algorithm for the joint base station (BS) and RIS beamforming optimization to minimize the transmit power subject to the user quality-of-service (QoS) constraints.
-By exploring the structure of this QoS problem, we show that the joint beamforming optimization can be naturally decomposed into a BS multicast beamforming QoS problem and an RIS passive multicast beamforming max-min-fair (MMF) problem.
-We propose an alternating multicast beamforming (AMBF) algorithm to solve the two subproblems alternatingly.
-For the BS QoS subproblem, we utilize the optimal multicast beamforming structure to obtain the BS beamformers efficiently.
-Furthermore, we reformulate the challenging RIS MMF subproblem and employ a first-order projected subgradient algorithm (PSA), which yields closed-form updates.
+This paper considers downlink multi-group multicasting via beamforming facilitated by a reconfigurable intelligent surface (RIS).
+We develop a fast and scalable algorithm for the joint base station (BS) and RIS beamforming optimization to minimize the transmit power while meeting user quality-of-service (QoS) targets.
+By analyzing the structure of the QoS constraints, we reformulate the problem and show that the joint beamforming optimization inherently consists of a multicast beamforming QoS problem for the BS and a passive multicast beamforming max-min-fair (MMF) problem for the RIS.
+We propose a fast alternating multicast beamforming (AMBF) algorithm to effectively solve the two subproblems alternatingly.
+For the BS multicast subproblem, we utilize the optimal multicast beamforming structure to efficiently determine the BS beamformers.
+For the RIS multicast subproblem, we reformulate the MMF problem and apply a first-order projected subgradient algorithm (PSA), which yields simple closed-form updates.
 The computational complexity of the AMBF algorithm grows linearly with the number of RIS elements and BS antennas.
-We further show that the AMBF approach is also an efficient method for the RIS-assisted downlink multi-user unicast beamforming problem, providing semi-closed-form updates.
-Next, we study the MMF problem for the RIS-assisted downlink beamforming design and propose a PSA-based fast algorithm to compute the BS and RIS beamforming solutions with closed-form updates per iteration, leading to a highly computationally efficient solution.
-Simulation results show the efficacy of our proposed algorithms in both performance and computational cost compared to other alternative methods.
+We further consider joint BS and RIS beamforming for the weighted MMF design objective subject to the BS transmit power budget.
+We propose an alternating PSA (APSA) fast algorithm to compute the beamforming solutions for the BS and RIS.
+APSA consists only closed-form updates per iteration, yielding linear computational complexity in the number of RIS elements and BS antennas.
+Simulation results show the efficacy of our proposed algorithms in terms of performance and computational cost compared to alternative methods.
 I Introduction
-Reconfigurable intelligent surface (RIS) is an emerging technology that can actively control and improve the wireless propagation channel conditions and thus creates a smart reconfigurable wireless environment to enhance communications performance [2, 3].
-RIS uses a planar surface consisting of passive reflective elements to control the phase shifts of the reflected wireless signal towards the desired direction with little energy usage, thus forming passive beamforming to enhance the transmission performance.
-It offers a range of advantages over the traditional options (e.g., base station (BS) or relay), such as no need for signal processing, low energy consumption and hardware cost, convenience for deployment, etc.
-Thus, RIS is a promising technology that has been explored to support a wide range of applications for next-generation wireless systems, such as coverage extension, capacity enhancement, improved localization and sensing [4, 5, 6, 7].
-For downlink transmissions, besides the typical dedicated unicast data transmission to each user, wireless multicasting has become increasingly popular to support content distribution and delivery in emerging applications.
-RIS can be explored to enhance multicast transmission performance.
-In conventional wireless systems, physical-layer multicast beamforming has been widely considered an efficient transmission technology for the BS to deliver common data to multiple users simultaneously.
-Various optimization algorithms or signal processing techniques have been developed in the literature to provide design solutions for different transmission scenarios or systems [8, 9, 10, 11, 12, 13, 14].
+Reconfigurable intelligent surface (RIS) is an emerging next-generation technology that can actively control and enhance wireless propagation channel conditions, thereby creating a smart reconfigurable wireless environment to enhance communications performance [2, 3].
+RIS uses a planar surface consisting of passive reflective elements to control the phase shifts of the reflected wireless signal towards desired directions with minimal energy consumption.
+This forms passive beamforming, which enhances the transmission performance.
+RIS offers several advantages over the traditional methods (e.g., base station (BS) or relay), such as the elimination of the need for complex signal processing, low energy consumption, reduced hardware cost, and convenience for deployment, etc.
+The potential role of RIS in next-generation wireless networks has been investigated to support a variety of applications, including coverage extension, capacity enhancement, and improved localization and sensing [4, 5, 6, 7, 8].
+Existing and emerging wireless services and technologies, such as video streaming, software dispatch distribution, edge storage and computing, and distributed machine learning, are driving significant growth of downlink content and data distribution.
+This demand increases the need of wireless multicast strategies to support efficiency delivery of shared content.
+In particular, with a multi-antenna BS, physical-layer multicasting via beamforming is an efficient transmission technique for the BS to deliver common data to multiple users simultaneously, which can effectively improve spectrum and power efficiency.
+Numerous optimization algorithms or signal processing techniques have been developed [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24].
 For delivering common data to a group, multicast beamforming performance is typically limited by the user in the group with the worst channel condition.
-Such limitation is dictated by the wireless prorogation environment.
-RIS can be utilized to break this barrier by adding an additional controllable channel path.
-In particular, it can form passive beamforming by controlling the phase shifts of the reflective elements to improve the user channel conditions to further enhance the overall transmission performance.
-Thus, it is important to explore this potential for RIS-assisted multicast beamforming design.
-In conventional systems, downlink multicast beamforming is a more challenging design problem than multi-user unicast beamforming, as the former is in general an NP-hard problem [8].
-This is also the case for the design of RIS-assisted downlink beamforming.
-RIS-assisted multicast beamforming design has recently been investigated for the single-group [15, 16] and multi-group settings [17, 18, 19, 20].
-Several approaches for joint BS and RIS beamforming optimization have been proposed for either maximizing sum group rate [17, 18], or minimizing the BS transmit power subject to the user quality-of-service (QoS) target constraints, i.e., the QoS problem [19, 20].
-In these works, the alternating optimization (AO) technique is commonly adopted for joint BS and RIS beamforming optimization.
-Depending on the specific problem formulation, smoothing techniques or SDR-based approaches have been proposed to solve the subproblems in each AO iteration.
-However, some limitations exist in these existing methods.
-For the QoS problem, existing studies are limited, and the efficacy of the proposed AO approaches in [19, 20] is uncertain as there is no guarantee of the quality of the computed solution.
-Also, the above existing methods generally have high computational complexity, which cannot be effectively implemented for a large number of RIS elements and BS antennas.
-Since RIS is beneficial only when the number of reflective elements becomes large, developing an effective design with both high performance and low computational complexity is crucial.
+This limitation arises from the wireless prorogation environment.
+RIS can be utilized to enhance multicast transmissions by adding an additional controllable channel path to mitigate this condition.
+By adjusting the phase shifts of the reflective elements, RIS enables passive beamforming that enhances the channel conditions for users, thereby improving the overall transmission performance.
+Therefore, it is crucial to explore effective and efficient designs for RIS-assisted multicast beamforming.
+Downlink multicast beamforming presents a challenging design problem, as it is generally a non-convex NP-hard problem [9].
+Earlier literature adopted the semi-definite relaxation (SDR) approach [9, 10, 11] to find approximate solutions.
+However, with the increasing number of BS antennas in evolved wireless systems, the successive convex approximation (SCA) technique has gained popularity due to its improved performance and lower computational complexity [12].
+To further reduce the solution complexity in massive multiple-input multiple-output (MIMO) systems, various computational methods or signal processing techniques have been developed for single-cell scenarios [13, 14, 15], multi-cell cooperation [16, 17], as well as robust designs that account for imperfect channel state information [18, 19].
+The optimal structure for multi-group multicast beamforming has been established recently in [20].
+It indicates that to design the BS beamformers, we only need to optimize the weights among users in each group.
+This significantly simplifies the optimization problem, as it is much smaller and independent of the number of BS antennas.
+This optimal structure has been leveraged to develop scalable ultra-low-complexity algorithms for massive MIMO downlink multicasting [21, 22], mixed traffic scenarios [23], and multi-cell coordination[24].
+The challenge of multicast beamforming design extends to RIS-assisted downlink multicasting.
+RIS-assisted multicast beamforming design has been investigated for the single-group [25, 26] and multi-group settings [27, 28, 29, 30].
+Several approaches have been proposed for designing joint BS and RIS beamforming [27, 28, 29, 30].
+The alternating optimization (AO) technique is commonly adopted for joint BS and RIS beamforming optimization.
+Joint BS and RIS beamforming optimization for maximizing the sum group rate is considered in [27, 28].
+In [27], a smoothing technique is proposed along with a majorization minimization method to solve the subproblems in each AO iteration.
+In [28], the smoothing technique is applied directly to the rate objective, along with a low-complexity alternating projected gradient algorithm.
+However, while maximizing the sum group rate can enhance overall data rate, it does not guarantee fairness among different groups.
+As a result, some groups may experience significant performance degradation, leading to an unacceptable user experience.
+To tackle the QoS issue, the QoS problem that minimizes the BS transmit power while meeting user QoS targets is studied in [29], where AO is directly applied to the optimization problem, resulting in a feasibility subproblem for the RIS passive beamforming design.
+The SDR-based approach is then proposed to solve the non-convex subproblems in each AO iteration.
+A similar approach is also considered in [30] for an RIS-assisted symbiotic radio multicast system comprising the primary receivers and an internet-of-things receiver.
+However, solving the feasibility problem does not ensure the quality of the solution for the RIS passive beamforming subproblem, raising questions about the effectiveness of this direct AO approach.
+Additionally, the methods mentioned above generally entail high computational complexity, which hinders their practical implementation.
+Most of these studies have only considered a relatively small number of RIS elements for performance evaluation.
+Since RIS systems typically require hundreds of elements to be beneficial, it is crucial to develop a scalable design solution that balances both high performance and low computational complexity.
 I-A Contribution
-In this paper, we consider the RIS-assisted downlink transmission design focusing on the multicast beamforming scenario.
-We develop our efficient fast algorithms for joint BS and RIS beamforming optimization for both the QoS problem and the max-min fair (MMF) problem, where the MMF problem refers to maximizing the weighted minimum SINR among users subject to the BS transmit power budget.
-Furthermore, we generalize our proposed approach and show that it can be extended to the unicast beamforming scenario, lending to a fast beamforming solution.
-Several unique features of our proposed algorithms distinguish our work from the existing ones.
-First, we propose a fast alternating multicast beamforming (AMBF) approach for the QoS problem.
-Different from the existing methods that directly apply AO to this problem, we explore the structure of the joint BS and RIS beamforming optimization and show that this QoS problem can be naturally broken down into two multicast beamforming subproblems, which are solved alternatingly: a BS multicast beamforming QoS problem and an RIS passive multicast beamforming MMF problem.
-This structure-based AMBF approach can effectively improve the BS and RIS beamforming solutions over iterations, avoiding the issue faced in the existing direct AO approaches.
-Next, with the two subproblems identified, we can judiciously exploit the optimal multicast beamforming structure [11] to compute a solution efficiently for the BS multicast problem, whose computational complexity does not depend on the number of BS antennas.
-This departs from the existing methods that are purely based on computational optimization techniques.
-For the challenging RIS passive multicast MMF problem, we develop our technique to reformulate the problem and adopt a first-order projected subgradient algorithm (PSA) [12] to solve it, which yields computationally cheap closed-form updates and has a convergence guarantee.
-Finally, we show that since our AMBF approach utilizes the structure of the problem, it is readily applicable to the downlink multi-user unicast beamforming scenario to compute the solution efficiently.
-Our fast algorithm is scalable, with computational complexity growing linearly with the number of RIS elements and the BS antennas.
-For the RIS-assisted downlink multicast MMF problem, we first show it is an inverse problem to the earlier QoS problem.
-Following this, we utilize the optimal BS beamforming structure again to transform the MMF problem into an equivalent problem, which allows us to propose a PSA-based iterative algorithm to compute the BS and RIS beamforming solutions.
-In particular, the algorithm uses closed-form updates per iteration, leading to a highly computationally efficient solution.
-The proposed algorithm is also directly applicable to the general downlink RIS-assisted multi-user beamforming problem with the MMF objective.
-The simulation results demonstrate the effectiveness of our proposed algorithms, in terms of the performance and computational advantage, for both the multicast and unicast scenarios compared to the alternative methods.
-I-B Related Work
-Multicast beamforming optimization is inherently a non-convex and NP-hard problem [8].
-The earlier literature works adopted the semi-definite relaxation (SDR) approach [8, 9] to find an approximate solution.
-With the number of antennas increasing, the successive convex approximation (SCA) technique has become more attractive for its lower computational complexity and better performance [10].
-The optimal multi-group multicast beamforming structure was recently obtained in [11].
-It shows that to obtain the beamforming vectors, we only need to optimize the per-user weight, which is a much smaller optimization problem that is independent to the number of BS antennas.
-This structure has been exploited to develop fast and scalable algorithms for multicast beamforming in massive multiple-input multiple-output (MIMO) systems [12, 13, 14].
-Several recent works have studied the RIS-assisted multicast beamforming design for the single group [15, 16] and the general multi-group settings [17, 18, 19, 20].
-Joint BS and RIS beamforming optimization for maximizing sum group rate is considered in [17, 18].
-In [17], the alternating optimization (AO) technique is used, and a smoothing technique is proposed along with a majorization minimization method to solve the subproblems.
-In [18], the smoothing technique is applied directly to the rate objective, along with a low-complexity alternating projected gradient algorithm.
-The QoS problem for the BS transmit power minimization is studied in [19], where AO is directly applied to the problem, resulting in a feasibility problem for the RIS passive beamforming design.
-The SDR-based approach is then proposed to solve the subproblems.
-A similar approach is also considered in [20] for an RIS-assisted symbiotic radio multicast system comprising the primary receivers and an internet-of-things receiver.
-However, as the feasibility problem cannot effectively search for an improved solution over iteration, there is no guarantee of the quality of this AO approach.
-The unicast scenario for RIS-assisted downlink multi-user beamforming has been studied in many recent works.
-Among them, only a few works have studied the BS power minimization problem [5, 21].
-In [5], directly applying AO to the joint beamforming optimization problem is first considered, and a low-complexity heuristic method based on the sum of channel gains is proposed.
-RIS-assisted non-orthogonal multiple access (NOMA) network is considered in [21], and the joint beamforming design using the direct AO approach is also adopted.
-Both [5] and [21] adopt semi-definite programming (SDP) or SDR to solve the corresponding alternating subproblems.
-The minimum rate (or SINR) maximization problem under various system configurations has been considered in [22, 23, 24].
-In [22], an iterative algorithm providing an alternating descent path is proposed.
-The work in [23] uses asymptotic analysis and some assumptions for the channel model to provide the approximations to the optimal BS transmit beamformer.
-Then, a projected gradient ascent algorithm is proposed for the RIS beamformer based on the large-scale channel statistics.
-RIS-assisted multi-cell coordinated transmit beamforming is considered in [24], where RIS can adjust both reflecting phases and amplitudes.
-The AO approach is adopted, and the second-order cone programming (SOCPs) and SDR are used to solve the subproblems.
-An inexact AO approach with SCA is further considered to reduce the computational complexity.
-However, these existing approaches generally have computational complexity.
-I-C Organization and Notations
+This paper aims to provide an effective joint BS and RIS beamforming design for RIS-assisted downlink data multicasting.
+We develop fast and scalable algorithmic solutions for two multicast beamforming formulations: the QoS problem, which minimizes the BS transmit power and the max-min fair (MMF) problem, which maximizes the weighted minimum signal-to-interference-and-noise ratio (SINR) among users, subject to the BS transmit power limit.
+We propose a fast alternating multicast beamforming (AMBF) approach for solving the QoS problem.
+This approach differs from existing methods that directly apply the AO technique to the joint optimization problem.
+Instead, we explore the structure of the constraints and reformulate the problem to reveal that the QoS problem inherently consists of two separate multicast beamforming optimization subproblems: a conventional multicast beamforming QoS problem for the BS and a passive multicast beamforming MMF problem for the RIS.
+The proposed AMBF approach solves these two subproblems alternatingly, effectively improving the BS and RIS beamforming solutions over iterations.
+We design fast algorithms to effectively solve the two subproblems.
+For the BS multicast problem, we utilize the optimal multicast beamforming structure [20] to compute a solution efficiently, where the computational complexity is independent of the number of BS antennas.
+This approach contrasts with existing methods for RIS-assisted beamforming, which rely on computational optimization techniques.
+For the more challenging RIS passive multicast MMF problem, we reformulate the problem to employ a first-order projected subgradient algorithm (PSA), which yields computationally inexpensive closed-form updates and has a guaranteed convergence.
+In the special case of the multi-user unicast beamforming scenario, the optimal BS beamformer can be computed exactly, and our AMBF approach requires only closed-form solutions or updates for both subproblems.
+Our AMBF algorithm is scalable, with computational complexity increasing linearly with the number of RIS elements and BS antennas.
+For the MMF problem, we show that it is an inverse problem to the earlier QoS problem.
+We leverage the optimal BS beamforming structure to transform the MMF problem into an equivalent lower-dimensional problem and propose an alternating PSA (APSA) algorithm that efficiently computes the BS and RIS beamforming solutions.
+Our APSA algorithm employs closed-form updates for both BS and RIS beamformers, maintaining linear computational complexity with respect to (w.r.t.) the number of RIS elements and BS antennas.
+Simulation results show the effectiveness of our proposed algorithms, demonstrating superior performance compared to existing AO-based algorithms and comparative to SCA-based joint optimization methods, while achieving significant reduction in computational time.
+I-B Organization and Notations
 The rest of this paper is organized as follows.
@@
 In Section IV, we present the fast algorithms to solve each beamforming subproblem.
-In Section V, we generalize our approach to solve the multi-user unicast beamforming.
-The MMF problem is considered in Section VI, where we discuss the relation between the QoS and MMF problems and propose a fast first-order algorithm to solve the MMF problem.
-The simulation results is provided in Section VII, followed by the conclusion in Section VIII.
+The MMF problem is considered in Section V, where we discuss the relation between the QoS and MMF problems and propose the APSA algorithm to solve the MMF problem efficiently.
+The simulation results is provided in Section VI, followed by the conclusion in Section VII.
 […]
@@
 II-A System Model
-We consider an RIS-assisted downlink multi-group multicast scenario: a BS equipped with $N$ antennas multicasts messages to $G$ user groups, and an RIS consisting of $M$ passive reflective elements is deployed to assist the data transmission between the BS and the users.
-We assume that group $i$ consists of $K_{i}$ single-antenna users, and these users receive a common message from the BS that is independent of other groups.
+We consider an RIS-assisted downlink multicast scenario: a BS equipped with $N$ antennas multicasts messages to $G$ user groups, and an RIS consisting of $M$ passive reflective elements is deployed to assist the data transmission between the BS and the users, as shown in Fig.
+1.
+We assume that group $i$ consists of $K_{i}$ single-antenna users, and the BS sends a common message to this group of users that is independent of other groups.
 The total number of users is denoted by $K_{\text{tot}}=\sum_{i=1}^{G}K_{i}$.
@@
 We denote the set of group indices by ${\cal G}\triangleq\{1,\ldots,G\}$, the set of user indices in group $i$ by ${\cal K}_{i}\triangleq\{1,\ldots,K_{i}\}$, $i\in{\cal G}$, and the set of RIS element indices by ${\cal M}\triangleq\{1,\ldots,M\}$.
+[…]
+Figure 1: An example of RIS-assisted downlink multi-group multicasting.
+Our design considers a general scenario where both the direct path and the RIS path are available between the BS and the users, [^1] In a rich scattering environment, although there may not be a line-of-sight path available, the non-line-of-sight paths typically still exist between the BS and a user.
+which includes the special case that only the RIS path is available.
 Let ${\bf H}_{\text{r}}$ denote the $M\times N$ channel matrix from the BS to the RIS, ${\bf h}^{\text{d}}_{ik}$ the $N\times 1$ channel vector from the BS to user $k$ in group $i$, and ${\bf h}^{\text{r}}_{ik}$ the $M\times 1$ channel vector from the RIS to user $k$ in group $i$, for $k\in{\cal K},i\in{\cal G}$.
@@
 The signal received at user $k$ in group $i$ is given by
-$\!\!y_{ik}$ $=\sum_{j=1}^{G}{\bf w}_{j}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf H}_{\text{r}}^{H}\text{diag}({\bf e}){\bf h}^{\text{r}}_{ik}\big{)}s_{j}+n_{ik}$
-$={\bf w}_{i}^{H}\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}s_{i}+\!\sum_{j\neq i}\!{\bf w}_{j}^{H}\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}s_{j}+n_{ik}$ (1)
+$y_{ik}$ $=\!\big{(}{\bf h}^{\text{d}}_{ik}+{\bf H}_{\text{r}}^{H}\text{diag}({\bf e}){\bf h}^{\text{r}}_{ik}\big{)}^{\!H}\!\sum_{j=1}^{G}{\bf w}_{j}s_{j}\!+n_{ik}$
+$=\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}^{\!H}{\bf w}_{i}s_{i}+\!\sum_{j\neq i}\!\big{(}{\bf h}^{\text{d}}_{ik}\!+\!{\bf G}_{ik}{\bf e}\big{)}^{\!H}{\bf w}_{j}s_{j}\!+\!n_{ik}$
 where $s_{i}$ is the symbol intended for group $i$ with $\mathbb{E}[|s_{i}|^{2}]=1$, ${\bf G}_{ik}\triangleq{\bf H}_{\text{r}}^{H}\text{diag}({\bf h}_{ik}^{\text{r}})$ is the $N\times M$ cascaded channel from the BS to user $k$ in group $i$ via the RIS, $n_{ik}\sim{\cal C}{\cal N}(0,\sigma^{2})$ is the receiver additive white Gaussian noise with variance $\sigma^{2}$.
 The received SINR at user $k$ in group $i$ is given by
-$\textrm{SINR}_{ik}=\frac{|{\bf w}_{i}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}|^{2}}{\sum_{j\neq i}|{\bf w}_{j}^{H}\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}|^{2}+\sigma^{2}},$ (2)
+$\textrm{SINR}_{ik}=\frac{|\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}|\big{(}{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}\big{)}^{H}{\bf w}_{j}|^{2}+\sigma^{2}},$ (1)
 and the total transmit power at the BS is $\sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$.
@@
 where ${\bf w}\triangleq[{\bf w}_{1}^{H},\ldots,{\bf w}_{G}^{H}]^{H}$, and $\gamma_{ik}$ is the SINR target for user $k$ in group $i$.
+Note that the SINR constraint in (2b) is equivalent to the per-user minimum rate constraint; it essentially defines a QoS target for each user.
+When $\gamma_{ik}=\tilde{\gamma}_{i}$, $\forall k$, this target becomes the group rate target, and constraint (2b) reduce to ensure that the minimum group rate target $\tilde{\gamma}_{i}$ is met for group $i$.
 Problem ${\cal P}_{o}$ is a non-convex quadratically constrained quadratic programming (QCQP) problem.
-Note that for the conventional downlink transmission without RIS, the multicast beamforming problem is already known to be NP-hard in its general form.
-The joint optimization problem ${\cal P}_{o}$ is even more challenging with the additional RIS passive beamformer ${\bf e}$ affecting $\textrm{SINR}_{ik}$ to be optimized under the non-convex constraints in (3c) imposed.
-Furthermore, since the number of RIS elements $M$ is expected to be large to show benefit, ${\cal P}_{o}$ is typically a large-scale problem.
-Thus, our goal is to develop an effective computational method to provide a solution with high performance but low computational complexity.
+For conventional downlink transmission without RIS, the multicast beamforming problem w.r.t.
+${\bf w}$ is known to be NP-hard.
+Therefore, the joint optimization problem ${\cal P}_{o}$ is even more challenging to solve, since it includes the additional RIS passive beamformer ${\bf e}$ affecting $\textrm{SINR}_{ik}$ to be optimized under the non-convex unit-modulus constraint (2c).
+Furthermore, the number of RIS elements $M$ is expected to be large to realize RIS benefits.
+Thus, ${\cal P}_{o}$ is typically a large-scale problem to solve.
+Our goal is to develop a fast and scalable algorithmic solution that is of high performance for RIS-assisted multicast beamforming.
 III RIS-assisted Multicast Beamforming Design
-We consider the alternating optimization (AO) technique to solve ${\cal P}_{o}$, i.e., solving ${\cal P}_{o}$ with respect to (w.r.t.) the BS multicast beamformer ${\bf w}$ and RIS passive beamformer ${\bf e}$ alternatingly.
+A common approach to solve ${\cal P}_{o}$ is to consider using alternating optimization (AO) to compute the BS multicast beamformer ${\bf w}$ and RIS passive beamformer ${\bf e}$ alternatingly.
 However, directly applying AO to ${\cal P}_{o}$ is not an effective method.
-This is because the objective function in ${\cal P}_{o}$ is only a function of ${\bf w}$.
-With given ${\bf w}$, ${\cal P}_{o}$ is reduced to a feasibility problem w.r.t.
-${\bf e}$, for which the value of ${\bf e}$ from the previous AO iteration is already a feasible solution.
-Thus, the algorithm could get stuck at the value of $({\bf w},{\bf e})$ after the first round of updates, and the quality of this solution merely depends on the quality of the initial point.
-In our approach, we first analyze the structure of ${\cal P}_{o}$ w.r.t. ${\bf w}$ and ${\bf e}$ and then utilize it to devise an effective algorithm to solve this QoS problem.
-III-A QoS Problem Reformulation
+This is because the objective is only a function of ${\bf w}$.
+Given ${\bf w}$, ${\cal P}_{o}$ is reduced to a feasibility problem w.r.t.
+${\bf e}$, which cannot ensure the quality of solution.
+For example, the value of ${\bf e}$ from the previous AO iteration is already a feasible solution.
+The algorithm could stuck at the value of $({\bf w},{\bf e})$ after the first round of updates, and the quality of this solution merely depends on the quality of the initial point.
+Thus, instead of directly applying AO, we first analyze the structure of ${\cal P}_{o}$ w.r.t.
+${\bf w}$ and ${\bf e}$ and then utilize it to devise an effective algorithm to solve ${\cal P}_{o}$.
+III-A Reformulation
 To explore the structural relation between ${\bf w}$ and ${\bf e}$ in ${\cal P}_{o}$, we first convert the set of SINR constraints in (2b) into the following equivalent constraint:
@@
 $\Leftrightarrow\quad\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1.$ (3)
-We can replace (3b) with (4) and equivalently rewrite ${\cal P}_{o}$ as
+Replacing (2b) with (3), we can equivalently rewrite ${\cal P}_{o}$ as
 ${\cal P}_{1}:\ \min_{{\bf w},{\bf e}}$ $\ \ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$ (4a)
@@
 ${\cal P}_{2}:$ $\ \min_{{\bf w}}\ \ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$ (5a)
-s.t. $\ \ \max_{{\bf e}:|e_{m}|^{2}=1,m\in{\cal M}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1.$ (6b)
-Note that for showing the equivalence, it is straightforward to show ${\cal P}_{1}$ and ${\cal P}_{2}$ have the same feasible set.
-Comparing ${\cal P}_{2}$ with ${\cal P}_{o}$, we see that the original set of SINR constraints in (3b) is replaced by the constraint in (6b).
-The left-hand side of (6b) is a max-min weighted SINRs optimization problem w.r.t.
+s.t. $\ \ \max_{\begin{subarray}{c}{\bf e}:|e_{m}|^{2}=1\\ m\in{\cal M}\end{subarray}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq 1.$ (5b)
+To see the equivalence, it is straightforward to show ${\cal P}_{1}$ and ${\cal P}_{2}$ have the same feasible set.
+From ${\cal P}_{o}$ to ${\cal P}_{2}$, the original set of SINR constraints in (2b) is now replaced by the constraint (5b).
+Note that the left hand side of this new constraint is a max-min weighted SINRs optimization problem w.r.t.
 ${\bf e}$, and the unit-modulus constraints on ${\bf e}$ in (2c) are now the constraints for this max-min optimization problem.
-III-B Alternating Multicast Beamforming Approach
-Once transforming ${\cal P}_{o}$ into ${\cal P}_{2}$, we notice that ${\cal P}_{2}$ inherently contains two optimization subproblems w.r.t.
+Thus, the joint optimization problem ${\cal P}_{o}$ is transformed into ${\cal P}_{2}$, which inherently contains two separate optimization subproblems w.r.t.
 ${\bf w}$ and ${\bf e}$, respectively.
-As we will show below, these two subproblems are essentially two types of multicast beamforming problems: 1) BS multicast beamforming QoS problem w.r.t.
+Following this, we can naturally break down ${\cal P}_{2}$ into two subproblems to solve alternatingly.
+III-B Proposed Alternating Multicast Beamforming Approach
+Examining the two subproblems in ${\cal P}_{2}$, we will show below that they are essentially two types of multicast beamforming problems: 1) BS multicast beamforming QoS problem w.r.t.
 ${\bf w}$; 2) RIS passive multicast beamforming MMF problem w.r.t.
 ${\bf e}$.
-Following this, we naturally break down ${\cal P}_{2}$ into two subproblems to solve alternatingly.
-Our overall proposed alternating multicast beamforming (AMBF) approach includes two steps: the alternating optimization step and a final processing step.
+We solve these two subproblems alternatingly.
+Our proposed alternating multicast beamforming (AMBF) approach includes two steps: 1) the alternating optimization step, and 2) a final processing step.
 The final processing step is to ensure a feasible solution to ${\cal P}_{o}$.
-Below, we will first describe the two subproblems and the final processing step, leaving the fast algorithm design for solving the subproblems to Sections IV-A and IV-B.
+Below, we will focus on describing the two subproblems and the final processing step, leaving the algorithm design for solving the subproblems to Sections IV-A and IV-B.
 III-B1 BS Multicast Beamforming QoS Problem for ${\bf w}$
-For given RIS passive beamforming vector ${\bf e}$, ${\cal P}_{o}$ is reduced to the following BS multicast beamforming problem w.r.t. ${\bf w}$:
+Given RIS passive beamforming vector ${\bf e}$, ${\cal P}_{o}$ is reduced to the following BS multicast beamforming problem w.r.t. ${\bf w}$:
 ${\cal P}_{{\bf e}}:\ \min_{{\bf w}}$ $\ \ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$ (6a)
-s.t. $\ \ \textrm{SINR}_{ik}\geq\gamma_{ik},\ k\in{\cal K}_{i},~{}i\in{\cal G}.$ (7b)
-Problem ${\cal P}_{\bf e}$ is a typical downlink multi-group multicast beamforming QoS problem.
+s.t. $\ \ \textrm{SINR}_{ik}\geq\gamma_{ik},\ k\in{\cal K}_{i},~{}i\in{\cal G},$ (6b)
+which is a conventional downlink multi-group multicast beamforming QoS problem.
 The effective channel between the BS and user $k$ in group $i$ consists of the direct path and the RIS-path given by ${\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}$, as shown in $\textrm{SINR}_{ik}$ in (1).
 III-B2 RIS Passive Multicast Beamforming MMF Problem for ${\bf e}$
-Optimizing RIS passive beamformer ${\bf e}$ with given ${\bf w}$ is the main challenge in ${\cal P}_{o}$.
-As discussed earlier, directly applying AO to the original problem ${\cal P}_{o}$ is not effective.
-However, once we convert ${\cal P}_{o}$ into ${\cal P}_{2}$, for given ${\bf w}$, we naturally have the following optimization problem w.r.t.
-${\bf e}$, which is to maximize the minimum ratio of the received SINR over its target, or the weighted SINR, i.e., ${\textrm{SINR}_{ik}}/{\gamma_{ik}}$:
+Given ${\bf w}$, RIS passive beamformer ${\bf e}$ optimization is the main challenge in ${\cal P}_{o}$.
+However, once we convert ${\cal P}_{o}$ into ${\cal P}_{2}$, we see that optimizing ${\bf e}$ is essentially to maximize the minimum ratio of the received SINR at each user over its target, i.e., the weighted SINR, as the following max-min problem:
 ${\cal S}_{{\bf w}}:$ $\ \ \max_{{\bf e}}\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ (7a)
-s.t. $\ \ |e_{m}|^{2}=1,\ m\in{\cal M}.$ (8b)
+s.t. $\ \ |e_{m}|^{2}=1,\ m\in{\cal M}$ (7b)
+where $\textrm{SINR}_{ik}$ in (1) can be re-expressed as
+$\textrm{SINR}_{ik}=\frac{|{\bf e}^{H}{\bf G}_{ik}^{H}{\bf w}_{i}+{\bf h}^{\text{d}}_{ik}\!{}^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}|{\bf e}^{H}{\bf G}_{ik}^{H}{\bf w}_{j}+{\bf h}^{\text{d}}_{ik}\!{}^{H}{\bf w}_{j}|^{2}+\sigma^{2}}.$ (8)
 Remark 1.
-Note that ${\cal S}_{{\bf w}}$ can be viewed as a weighted MMF single-group multicast beamforming problem, with ${\bf e}$ being the beamformer to all the users in the system.
-However, this problem has three key differences from the conventional MMF multicast problem: i) This single-group multicast beamforming problem contains self-interference, as ${\bf e}$ appears in the denominator of $\textrm{SINR}_{ik}$ in (2).
-ii) From the point of view of optimizing ${\bf e}$, $\textrm{SINR}_{ik}$ in (2) can in fact be interpreted as the signal-to-leakage-and-noise ratio (SLNR) for user $k$ in group $i$.
+Note that ${\cal S}_{{\bf w}}$ can be viewed as a variant of single-group multicast beamforming weighted MMF problem, with ${\bf e}$ being the multicast beamformer to all the users.
+This can be more clearly seen in the RIS-path only case, given by
+$\textrm{SINR}_{ik}=\frac{|{\bf e}^{H}{\bf G}_{ik}^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}|{\bf e}^{H}{\bf G}_{ik}^{H}{\bf w}_{j}|^{2}+\sigma^{2}}.$ (9)
+But it has a few differences from the conventional MMF multicast beamforming: i) It contains self-interference, since ${\bf e}$ appears in the denominator of $\textrm{SINR}_{ik}$ in (9).
+ii) From the perspective of RIS beamformer ${\bf e}$, $\textrm{SINR}_{ik}$ in (9) can be interpreted as the signal-to-leakage-and-noise ratio (SLNR) for user $k$ in group $i$.
+The power leakage is through the effective channels formed by other BS beamformers ${\bf w}_{j}$’s to the same user, ${\bf G}_{ik}^{H}{\bf w}_{j}$, $\forall j\neq i$.
 Thus, the MMF problem ${\cal S}_{{\bf w}}$ w.r.t.
 ${\bf e}$ is based on an SLNR metric.
-iii) Different from the conventional transmit power constraint, the constraints in (8b) can be viewed as a non-convex per-element power constraint, where each element is required to consume exactly one unit of power.
-Due to the difference discussed above, ${\cal S}_{\bf w}$ is a more challenging problem to solve than the conventional MMF problem, which is a non-convex QCQP problem that is NP-hard.
+For the general case with a direct path in (8), the combined beamed signal strength from the RIS and the direct path need to be considered.
+iii) Different from the conventional total transmit power constraint, the constraints in (7b) can be viewed as a non-convex per-element power constraint, where each element is required to consume exactly one unit of power.
+Due to the difference discussed above, ${\cal S}_{\bf w}$ is a more challenging problem to solve than the conventional MMF problem, which is a non-convex QCQP problem.
 To tackle this problem, we note that the constraints in (7b) are equivalent to
-$|e_{m}|^{2}\leq 1,m\in{\cal M};\quad\quad{\bf e}^{T}{\bf e}=M.$ (9)
+$|e_{m}|^{2}\leq 1,m\in{\cal M}\quad\text{and}\quad{\bf e}^{T}{\bf e}=M.$ (10)
 Thus, we replace the constraints in (7b) with (10) and transform ${\cal S}_{\bf w}$ into the following equivalent problem:
@@
 $\ \ {\bf e}^{T}{\bf e}=M.$ (11c)
-To make the problem more tractable, we relax ${\cal S}_{\bf w}^{\text{eq}}$ by transferring the constraint in (10c) into the objective function in (10a) as a penalty term with a penalty weight $\zeta>0$.
+To make the problem more tractable, we relax ${\cal S}_{\bf w}^{\text{eq}}$ by transferring the constraint in (11c) into the objective function in (11a) as a penalty term with a penalty weight $\rho>0$.
 The relaxed problem is given by
-$\widetilde{{\cal S}}_{{\bf w}}:\ \max_{{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\zeta\frac{{\bf e}^{H}{\bf e}}{M}$ (11a)
+$\widetilde{{\cal S}}_{{\bf w}}:\ \max_{{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\rho\frac{{\bf e}^{H}{\bf e}}{M}$ (12a)
 s.t. $\ \ |e_{m}|^{2}\leq 1,\ m\in{\cal M}.$ (12b)
@@
 The penalty term in (12a) is to ensure that ${\bf e}^{T}{\bf e}$ is as close to $M$ as possible in solving $\widetilde{{\cal S}}_{{\bf w}}$.
-Note that we choose the penalty term in this form such that the two terms in the objective function are both normalized.
-In particular, for the weighted SINR term, using $({\bf w},{\bf e})$ from ${\cal P}_{\bf e}$, we have $\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}=1$ based on (7b).
-Thus, we expect this term to be close to $1$, especially as the AO iteration between ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{\bf w}$ converges.
-The RIS passive beamformer norm term satisfies $\frac{{\bf e}^{H}{\bf e}}{M}\leq 1$ based on (10b), and it is also expected to be close to $1$ as the AO iteration converges.
-In this way, the two terms in (11a) have comparable values that are both around $1$, which are not affected by the actual values of SINR or $\|{\bf e}\|^{2}$ for a specific setup.
-This simplifies the tuning of penalty parameter $\zeta$, which can remain the same for different system configurations.
-Although $\widetilde{{\cal S}}_{{\bf w}}$ is still a non-convex QCQP problem, compared with ${\cal S}_{{\bf w}}$, it is more amenable to efficient algorithm design to compute a solution, which will be discussed in Section IV-B.
-Thus, following the above, the first step in our AMBF approach is to solve the two multicast beamforming problems alternatingly, i.e., ${\cal P}_{\bf e}$ for the BS beamformers $\{{\bf w}_{i}\}$ and $\widetilde{{\cal S}}_{{\bf w}}$ for the RIS beamformer ${\bf e}$.
+Note that we choose the penalty term in this particular form such that the two terms in the objective function are both normalized.
+In particular, for first term, we note that ${\bf w}$ is the solution to ${\cal P}_{\bf e}$, which yields $\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}=1$ in (6b).
+Thus, we expect this term to be close to $1$, especially when the iterative procedure between ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{\bf w}$ converges.
+The second term for the normalized RIS passive beamformer satisfies $\frac{{\bf e}^{H}{\bf e}}{M}\leq 1$ based on (12b), and it is also expected to be close to $1$ when the AO procedure converges.
+Therefore, the two terms in (12a) have comparable values that are both close to $1$.
+Being normalized, they do not depend on the actual values of SINR or $\|{\bf e}\|^{2}$ for a specific system configuration.
+This simplifies the tuning of penalty parameter $\rho$, which can remain the same for different system configurations.
+Compared with ${\cal S}_{{\bf w}}$, $\widetilde{{\cal S}}_{{\bf w}}$ is more amenable to efficient algorithm design to compute a solution, which will be discussed in Section IV-B.
+Combining Sections III-B1 and III-B2, the first step in our AMBF approach is to solve the two multicast beamforming problems alternatingly, i.e., ${\cal P}_{\bf e}$ for the BS beamformers $\{{\bf w}_{i}\}$ and $\widetilde{{\cal S}}_{{\bf w}}$ for the RIS beamformer ${\bf e}$.
 III-B3 Final Processing
-Let $({\bf w}^{\star},{\bf e}^{\star})$ denote the solution after the above iterative procedure.
+Let $({\bf w}^{\star},{\bf e}^{\star})$ denote the solution from the above AO procedure.
 Since $\widetilde{{\cal S}}_{{\bf w}}$ is a relaxed problem of ${\cal S}_{{\bf w}}$, $({\bf w}^{\star},{\bf e}^{\star})$ may not be feasible to ${\cal P}_{o}$.
 Thus, based on $({\bf w}^{\star},{\bf e}^{\star})$, we have the final processing step to obtain the feasible solution $({\bf w}^{\text{final}},{\bf e}^{\text{final}})$ as follows:
-  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which represents taking the phase component of each element in ${\bf e}^{\star}$.
+  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which is taking the phase component of each element in ${\bf e}^{\star}$.
   • Solve ${\cal P}_{{\bf e}^{\text{final}}}$ with ${\bf e}^{\text{final}}$ and obtain ${\bf w}^{\text{final}}$.
-IV Fast Algorithms for the AMBF Approach
-In this section, we propose our fast algorithms to solve subproblems ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{{\bf w}}$.
-IV-A Fast Algorithm for BS Multicast Beamforming ${\cal P}_{\bf e}$
-As mentioned earlier, ${\cal P}_{\bf e}$ is a typical downlink multicast beamforming QoS problem.
-Define $\tilde{{\bf h}}_{ik}\triangleq{\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e}$ as the $N\times 1$ equivalent channel from BS to user $k$ in group $i$ that includes both the direct path and the RIS path.
-The received SINR at user $k$ in group $i$ in (2) can be rewritten as
-$\textrm{SINR}_{ik}=\frac{|{\bf w}_{i}^{H}\tilde{{\bf h}}_{ik}|^{2}}{\sum_{j\neq i}^{G}|{\bf w}_{j}^{H}\tilde{{\bf h}}_{ik}|^{2}+\sigma^{2}},\ k\in{\cal K}_{i},i\in{\cal G}.$ (12)
-Although the QoS problem ${\cal P}_{\bf e}$ is known to be a challenging non-convex QCQP problem, the structure of the optimal solution is obtained in [11], which can be utilized to compute the beamforming solution ${\bf w}$ with high computational efficiency, especially when the number of BS antennas is large $M\gg 1$.
+IV Fast Algorithms for alternating multicast beamforming
+We now propose our fast algorithm for each of the multicast subproblems ${\cal P}_{\bf e}$ and $\widetilde{{\cal S}}_{{\bf w}}$ under the proposed AMBF approach.
+IV-A BS Multicast Beamforming QoS Problem ${\cal P}_{\bf e}$
+As mentioned earlier, ${\cal P}_{\bf e}$ is a conventional downlink multicast beamforming QoS problem.
+Define $\tilde{{\bf h}}_{ik}\triangleq({\bf h}^{\text{d}}_{ik}+{\bf G}_{ik}{\bf e})$ as the $N\times 1$ equivalent channel from BS to user $k$ in group $i$ that includes both the direct path and the RIS path.
+The received SINR in (1) at user $k$ in group $i$ can be rewritten as
+$\textrm{SINR}_{ik}=\frac{|\tilde{{\bf h}}_{ik}^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}^{G}|\tilde{{\bf h}}_{ik}^{H}{\bf w}_{j}|^{2}+\sigma^{2}},\ k\in{\cal K}_{i},i\in{\cal G}.$ (13)
+Despite that ${\cal P}_{\bf e}$ is an NP-hard non-convex QCQP problem, the structure of the optimal solution is obtained in [20], which can be utilized to compute the beamforming solution ${\bf w}$ with high computational efficiency, especially when the number of BS antennas is large $N\gg 1$.
 We directly employ this optimal structure to compute ${\bf w}$.
@@
 ${\bf w}_{i}\,=\,{\bf R}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf H}}_{i}{\bf a}_{i}\,,\;i\in{\cal G}$ (14)
-where $\widetilde{{\bf H}}_{i}\triangleq[\tilde{{\bf h}}_{i1},\ldots,\tilde{{\bf h}}_{iK}]$ is the equivalent channel matrix for group $i$, ${\bf a}_{i}$ is the $K\times 1$ optimal weight vector for group $i$ containing the (complex) weight for each user channel in the group, and ${\bf R}(\boldsymbol{\lambda})\triangleq{\bf{I}}+\sum_{i=1}^{G}\sum_{k=1}^{K}\lambda_{ik}\gamma_{ik}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}$ is the noise-plus-weighted-channel covariance matrix with $\lambda_{ik}$ being the optimal Lagrange multipliers associated with the SINR constraint in (7b) and $\boldsymbol{\lambda}$ being the vector containing all $\lambda_{ik}^{o}$’s.
+where $\widetilde{{\bf H}}_{i}\triangleq[\tilde{{\bf h}}_{i1},\ldots,\tilde{{\bf h}}_{iK}]$ is the equivalent channel matrix for users in group $i$, ${\bf a}_{i}$ is the $K\times 1$ optimal weight vector for group $i$ containing the (complex) weight for each user channel in the group, and ${\bf R}(\boldsymbol{\lambda})\triangleq{\bf{I}}+\sum_{i=1}^{G}\sum_{k=1}^{K}\lambda_{ik}\gamma_{ik}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}$ is the noise-plus-channel covariance matrix with $\lambda_{ik}$ being the optimal Lagrange multipliers associated with the SINR constraint in (6b) and $\boldsymbol{\lambda}$ being the vector containing all $\lambda_{ik}^{o}$’s.
 The solution ${\bf w}_{i}$ in (14) is given in a semi-closed form with $\boldsymbol{\lambda}$ and $\{{\bf a}_{i}\}$ to be numerically determined.
-Although the optimal value for $\boldsymbol{\lambda}$ is difficult to obtain, an algorithm to compute (suboptimal) values of $\boldsymbol{\lambda}$ efficiently using the fixed-point method is proposed in [11].
-With the optimal structure of ${\bf w}_{i}$ in (13) and $\boldsymbol{\lambda}$, ${\cal P}_{\bf e}$ can be transformed into a weight optimization problem w.r.t.
+We can use the algorithm proposed in [20] to compute (suboptimal) values of $\boldsymbol{\lambda}$ efficiently using the fixed-point method.
+Given $\boldsymbol{\lambda}$ and ${\bf w}_{i}$ in (14), ${\cal P}_{\bf e}$ can be transformed into a weight optimization problem w.r.t.
 weight vectors $\{{\bf a}_{i}\}$, which is a much smaller problem with $\sum_{i=1}^{G}K_{i}$ variables compared to the original problem ${\cal P}_{\bf e}$ with $GN$ variables for $N\gg K_{i}$.
-To solve this non-convex weight optimization problem for $\{{\bf a}_{i}\}$, we can adopt the SCA approach discussed in [11].
-Further improvement in the computational efficiency can be considered by adopting the fast ADMM algorithm proposed recently in [13] based on the optimal structure, which provides each SCA update using closed-form expressions with ultra-low complexity.
+To solve the non-convex weight optimization problem for $\{{\bf a}_{i}\}$, we can adopt the fast algorithm proposed in [22].
+It is based on the SCA approach and uses an ADMM-based algorithm to provide each SCA update using closed-form expressions with ultra-low complexity.
 We omit the details and refer the readers to [20, 22] for these algorithms.
-IV-B Fast Algorithm for RIS Passive Multicast Beamforming $\widetilde{{\cal S}}_{\bf w}$
-Problem $\widetilde{{\cal S}}_{{\bf w}}$ is expected to be large-size, since the number of RIS elements is large ($M\gg 1$).
-Existing methods in the literature for solving ${\bf e}$ often adopt SDR or SCA [17, 20], but suffer from high computational complexity.
-It is essential to develop an effective low-complexity algorithm to find a solution to $\widetilde{{\cal S}}_{{\bf w}}$, especially when this subproblem needs to be solved in each AO iteration.
+IV-B RIS Passive Multicast Beamforming MMF Problem $\widetilde{{\cal S}}_{\bf w}$
+The max-min problem $\widetilde{{\cal S}}_{{\bf w}}$ is expected to be large-size, since the number of RIS elements $M$ is large.
+Existing methods for solving ${\bf e}$ often adopt certain convexification methods, such as SDR or SCA [27, 30].
+However, they suffer from high computational complexity for large $M$.
+Given this subproblem needs to be solved in each AO iteration., it is essential to develop an effective low-complexity algorithm to find a solution to $\widetilde{{\cal S}}_{{\bf w}}$.
 From Remark 1, we note that $\widetilde{{\cal S}}_{{\bf w}}$ is a variant of the MMF problem.
-PSA is a fast first-order algorithm recently proposed in[12] to solve the multi-group multicast beamforming MMF problem in the conventional system.
-Although the form of the objective function and the constraints in $\widetilde{{\cal S}}_{{\bf w}}$ are different from those of the conventional MMF problem considered in [12], the method can still be adapted to solve $\widetilde{{\cal S}}_{{\bf w}}$.
-Thus, we propose to apply PSA to $\widetilde{{\cal S}}_{{\bf w}}$ to compute a near-stationary solution of $\widetilde{{\cal S}}_{{\bf w}}$ efficiently.
+We propose to apply the projected subgradient algorithm (PSA) [31] to $\widetilde{{\cal S}}_{{\bf w}}$.
+PSA is a first-order iterative algorithm.
+It has been adopted in[21] to solve the conventional multi-group multicast beamforming MMF problem efficiently.
+The structure of the objective function and constraints in $\widetilde{{\cal S}}_{\bf w}$ are different from the conventional MMF problem.
+Nonetheless, we can still employ the essential technique of this method to compute a near-stationary solution of $\widetilde{{\cal S}}_{{\bf w}}$ efficiently.
+Given our problem structure, applying PSA proves to be computationally inexpensive, which is the main advantage of this approach.
 IV-B1 Problem Reformulation
-Based on the objective function in (11a) and using the SINR expression in (2), we define
-$\phi_{ik}({\bf e})\triangleq-\bigg{(}\frac{1}{\gamma_{ik}}\frac{|{\bf w}_{i}^{H}({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})|^{2}}{\sum_{j\neq i}|{\bf w}_{j}^{H}({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})|^{2}\!+\!\sigma^{2}}+\zeta\frac{{\bf e}^{H}{\bf e}}{M}\bigg{)},$ (14)
+We first transform $\widetilde{{\cal S}}_{\bf w}$ into a min-max problem formulation. Based on the objective function in (12a) and using the SINR expression in (1), we define
+$\phi_{ik}({\bf e})\triangleq-\bigg{(}\frac{1}{\gamma_{ik}}\frac{|({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})^{H}{\bf w}_{i}|^{2}}{\sum_{j\neq i}|({\bf h}_{ik}^{\text{d}}\!+\!{\bf G}_{ik}{\bf e})^{H}{\bf w}_{j}|^{2}\!+\!\sigma^{2}}+\rho\frac{{\bf e}^{H}{\bf e}}{M}\bigg{)},$ (15)
 for $k\in{\cal K}_{i}$, $i\in{\cal G}$.
-Furthermore, we denote the feasible set of $\widetilde{{\cal S}}_{{\bf w}}$ as ${\cal E}\triangleq\{{\bf e}:|e_{m}|^{2}\leq 1,\ m\in{\cal M}\}$.
+We denote the feasible set of $\widetilde{{\cal S}}_{{\bf w}}$ as ${\cal E}\triangleq\{{\bf e}:|e_{m}|^{2}\leq 1,\ m\in{\cal M}\}$.
 Then, $\widetilde{{\cal S}}_{{\bf w}}$ can be rewritten as a min-max problem:
 $\widetilde{{\cal S}}_{{\bf w}}:\min_{{\bf e}\in{\cal E}}\max_{i,k}\phi_{ik}({\bf e}).$ (16)
-The inner maximization problem is an integer program.
-To tackle this, let $\boldsymbol{\phi}({\bf e})\triangleq[\phi_{11}({\bf e}),\ldots,\phi_{1K_{1}}({\bf e}),\ldots,\phi_{G1}({\bf e}),\ldots,\phi_{GK_{G}}({\bf e})]^{T}$, and also let ${\bf y}$ be a $K_{\text{tot}}\times 1$ probability vector satisfying ${\bf y}\succcurlyeq 0$ and $\mathbf{1}^{T}{\bf y}=1$.
+To tackle the inner maximization problem, which is an integer program, let $\boldsymbol{\phi}({\bf e})$ be the $K_{\text{tot}}\times 1$ vector containing all $\phi_{ik}({\bf e})$’s, for $k\in{\cal K}_{i}$, $i\in{\cal G}$.
+Also let ${\bf y}$ be a $K_{\text{tot}}\times 1$ probability vector satisfying ${\bf y}\succcurlyeq 0$ and $\mathbf{1}^{T}{\bf y}=1$.
+Let $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y}$, and and ${\cal Y}\triangleq\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$.
 Then, we can further transform problem (16) into the following equivalent problem
-$\widetilde{{\cal S}}_{{\bf w}}^{\prime}:\min_{{\bf e}\in{\cal E}}\max_{{\bf y}\in{\cal Y}}$ $\ f({\bf e},{\bf y})$ (16)
-where $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y}$ and ${\cal Y}\triangleq\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$.
-Note that ${\cal Y}$ is a probability simplex.
-Thus, an optimal solution to the inner maximization problem in $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is ${\bf y}=[0,\cdots,1,\cdots,0]^{T}$, with 1 at some position that corresponds to $\max_{i,k}\phi_{ik}({\bf e})$, which is equivalent to the inner maximization in (15).
-Note that both ${\cal E}$ and ${\cal Y}$ are compact convex sets.
-IV-B2 The Projected Subgradient Algorithm
-Note that $f({\bf e},{\bf y})$ is concave in ${\bf y}$ but non-convex in ${\bf e}$.
-Thus, problem $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is a non-convex-concave min-max problem, which is typically NP-hard.
-Furthermore, the objective function $\max_{{\bf y}\in{\cal Y}}f({\bf e},{\bf y})$ may not be differentiable.
-To solve $\widetilde{{\cal S}}^{\prime}$, we notice that $\widetilde{{\cal S}}^{\prime}$ has the same structure as the min-max problem considered in [12] (i.e., with objective function $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y}$ and a convex feasible set).
-Thus, we adopt PSA in [12] to find a near-stationary point of $\widetilde{\cal S}^{\prime}_{{\bf w}}$.
-Specifically, PSA is an iterative algorithm, with the updating procedure at iteration $j$ given by [^1] Note that PSA in [12] is provided using all real-valued variables.
-We provide the complex-valued version of this procedure, which can be shown to be equivalent to the real-valued version.
+$\widetilde{{\cal S}}_{{\bf w}}^{\prime}:\min_{{\bf e}\in{\cal E}}\max_{{\bf y}\in{\cal Y}}$ $\ f({\bf e},{\bf y}).$ (17)
+Both ${\cal E}$ and ${\cal Y}$ are compact convex sets, and ${\cal Y}$ is a probability simplex.
+Thus, an optimal solution to the inner maximization problem in $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is ${\bf y}=[0,\cdots,1,\cdots,0]^{T}$, with 1 at a position that corresponds to $\max_{i,k}\phi_{ik}({\bf e})$, which is equivalent to the inner maximization in (16).
+Problem $\widetilde{\cal S}^{\prime}_{{\bf w}}$ is a nonconvex-concave min-max problem, since the objective function $f({\bf e},{\bf y})$ is concave in ${\bf y}$ but nonconvex in ${\bf e}$.
+Furthermore, function $g({\bf e})\triangleq\max_{{\bf y}\in{\cal Y}}f({\bf e},{\bf y})$ may not be differentiable w.r.t.
+${\bf e}$, and thus, gradient-based methods are not applicable.
+We note that the structure of $\widetilde{{\cal S}}_{{\bf w}}^{\prime}$ is the same as the min-max problem structure considered in [21] (i.e., the objective function is in the form of $f({\bf e},{\bf y})\triangleq\boldsymbol{\phi}^{T}({\bf e}){\bf y},$ and the feasible set is convex).
+It is shown in [21] that in this formulation, $f({\bf e},{\bf y})$ is an $L$-smooth function of ${\bf e}$, and $\nabla_{\bf e}f({\bf e},{\bf y})$ is a subgradient of $g({\bf e})$.
+In addition, $g({\bf e})$ is $L$-weakly convex over ${\cal E}$.
+Thus, we can adopt PSA [31] to find a near-stationary point of $\widetilde{\cal S}^{\prime}_{{\bf w}}$.
+IV-B2 Projected Subgradient Algorithm
+PSA is an iterative algorithm.
+Its updating procedure at iteration $j$ given by [^2] Note that PSA in [21] is provided using all real-valued variables.
+We provide the complex-valued version of this procedure, which can be shown to be equivalent.
 ${\bf y}^{(j)}$ $\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}f({\bf e}^{(j)},{\bf y});$ (18)
@@
 $\Pi_{{\cal E}}({\bf e})=\begin{cases}e_{m}&\mbox{if }|e_{m}|\leq 1\\ \frac{e_{m}}{|e_{m}|}&\text{o.w}\end{cases},\ \text{for}\ m\in{\cal M}.$ (20)
-In (17), ${\bf y}^{(j)}$ can be obtained by finding index $\hat{i}\hat{k}$ such that $\phi_{\hat{i}\hat{k}}({\bf e})=\max_{i,k}\phi_{ik}({\bf e})$, and we have $f({\bf e},{\bf y}^{(j)})=\phi_{\hat{i}\hat{k}}({\bf e})$.
-To compute the gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (18), with a slight abuse of notation to ease our presentation, we use $\phi_{ik}({\bf e})$ to represent $\phi_{i^{\prime}k^{\prime}}({\bf e})$ obtained in (17).
-Then, $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ has the following expression:
+In (18), ${\bf y}^{(j)}$ can be obtained by finding an index pair $(\hat{i},\hat{k})$ such that $\phi_{\hat{i}\hat{k}}({\bf e})=\max_{i,k}\phi_{ik}({\bf e})$, and we have $f({\bf e},{\bf y}^{(j)})=\phi_{\hat{i}\hat{k}}({\bf e})$.
+The gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (19) has the following expression:
 $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})=\nabla_{{\bf e}}\phi_{\hat{i}\hat{k}}({\bf e})$
-$\quad=-\frac{{\bf q}_{\hat{i}\hat{i}\hat{k}}({\bf e})I_{\hat{i}\hat{k}}\!({\bf e})-\!\sum_{j\neq\hat{i}}{\bf q}_{j\hat{i}\hat{k}}({\bf e})S_{\hat{i}\hat{k}}({\bf e})}{\gamma_{\hat{i}\hat{k}}I_{\hat{i}\hat{k}}^{2}({\bf e})}-\frac{\zeta}{M}{\bf e}$ (20)
-where $I_{\hat{i}\hat{k}}({\bf e})\triangleq\sum_{j\neq\hat{i}}|{\bf w}_{j}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})|^{2}+\sigma^{2}$, $S_{\hat{i}\hat{k}}({\bf e})\triangleq|{\bf w}_{\hat{i}}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})|^{2}$, and ${\bf q}_{j\hat{i}\hat{k}}({\bf e})={\bf G}_{\hat{i}\hat{k}}^{H}{\bf w}_{j}{\bf w}_{j}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})$.
-The updates in each iteration are closed-form evaluations as in (17)–(IV-B2), with only matrix and vector multiplications.
+$\quad=-\frac{{\bf q}_{\hat{i}\hat{i}\hat{k}}({\bf e})I_{\hat{i}\hat{k}}\!({\bf e})-\!\sum_{j\neq\hat{i}}{\bf q}_{j\hat{i}\hat{k}}({\bf e})S_{\hat{i}\hat{k}}({\bf e})}{\gamma_{\hat{i}\hat{k}}I_{\hat{i}\hat{k}}^{2}({\bf e})}-\frac{\rho}{M}{\bf e}$ (21)
+where $I_{\hat{i}\hat{k}}({\bf e})\triangleq\sum_{j\neq\hat{i}}|({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})^{H}{\bf w}_{j}|^{2}+\sigma^{2}$, $S_{\hat{i}\hat{k}}({\bf e})\triangleq|({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})^{H}{\bf w}_{\hat{i}}|^{2}$, and ${\bf q}_{j\hat{i}\hat{k}}({\bf e})={\bf G}_{\hat{i}\hat{k}}^{H}{\bf w}_{j}{\bf w}_{j}^{H}({\bf h}_{\hat{i}\hat{k}}^{\text{d}}+{\bf G}_{\hat{i}\hat{k}}{\bf e})$.
+The updates in each PSA iteration involve closed-form evaluations as in (18)–(IV-B2), with only matrix and vector multiplications.
 Thus, PSA is computationally cheap and is particularly suitable for solving $\widetilde{{\cal S}}_{{\bf w}}$ with a large value of $M$.
-Furthermore, the convergence analysis of PSA in [12] is applicable to our problem.
+IV-B3 Convergence and Stopping Criterion
+The convergence analysis of PSA in [21] is applicable to our problem.
 It shows that the above PSA procedure is guaranteed to converge in finite time to a point at the vicinity of a stationary point of $\widetilde{{\cal S}}_{{\bf w}}$.
+Since PSA is a subgradient method, it is not guaranteed to move toward a descending direction.
+Thus, as commonly considered, we track the best solution obtained so far in $j$ iterations, i.e., the one with the minimum objective value of $\widetilde{{\cal S}}_{{\bf w}}^{\prime}$ in $j$ iterations: $g_{\min}^{(j)}=\min\{g^{(1)},...,g^{(j)}\}$, where $g^{(j)}=\max_{{\bf y}\in{\cal Y}}f({\bf e}^{(j)},{\bf y})$.
+Note that this is equivalent to track the maximum objective value in (12a) obtained so far.
+Based on its definition, $g_{\min}^{(j)}$ is monotonically decreasing and is bounded below.
+Thus, we can terminate the algorithm when $g_{\min}^{(j)}$ converges.
 IV-C Discussion on the AMBF Algorithm
 We summarize our proposed fast AMBF algorithm for the RIS-assisted multicast QoS problem ${\cal P}_{o}$ in Algorithm 1.
-It consists of the proposed AMBF approach in Section III-B and the fast algorithms in Sections IV-A and IV-B for each subproblem.
+It consists of the proposed AMBF approach in Section III-B and the fast algorithms in Sections IV-A and IV-B for the two subproblems, respectively.
 A few aspects of the algorithm are discussed below:
 Algorithm 1 Alternating Multicast Beamforming Algorithm (AMBF) for RIS-Assisted QoS Problem ${\cal P}_{o}$.
-1:Initialization: Set feasible initial point ${\bf e}^{0}$; Set $n=0$.
+1:Initialization: Set feasible initial point ${\bf e}^{0}$; Set ${\rho}$ and $\alpha$; Set $n=0$.
 2:repeat
-3:     With ${\bf e}^{n}$, solve ${\cal P}_{\bf e}$ to obtain ${\bf w}^{n+1}$ by using (13).
+3:     With ${\bf e}^{n}$, solve ${\cal P}_{\bf e}$ to obtain ${\bf w}^{n}$ by using (14).
 4:     With ${\bf w}^{n}$, set ${\bf e}^{(0)}={\bf e}^{n}$; Solve $\widetilde{{\cal S}}_{{\bf w}}$ using PSA updates (18) and (19) to obtain ${\bf e}^{n+1}$.
@@
 In each AO iteration $n$, solving $\tilde{{\cal S}}_{\bf w}$ by PSA also requires an initial point ${\bf e}^{(0)}$ for the update in (18).
-We should set ${\bf e}^{(0)}={\bf e}^{n}$, which is from the previous AO iteration and is also used in ${\cal P}_{\bf e}$ in the current AO iteration.
+We should set ${\bf e}^{(0)}={\bf e}^{n}$, i.e., the solution from the previous AO iteration.
+This is also used in ${\cal P}_{\bf e}$ to obtain ${\bf w}^{n+1}$ in the current AO iteration.
 IV-C2 Computational Complexity
-Algorithm 1 consists of solving two optimization subproblems ${\cal P}_{{\bf e}}$ and $\widetilde{{\cal S}}_{{\bf w}}$ at each AO iteration.
-We discuss the computational complexity for solving each subproblem below:
-Problem ${\cal P}_{\bf e}$
-As mentioned in Section IV-A, using the optimal beamforming structure ${\bf w}$ in (13), ${\cal P}_{{\bf e}}$ is converted into the weight optimization problem w.r.t.
-$\{{\bf a}_{i}\}$ with total $K_{\text{tot}}(=\sum_{i=1}^{G}K_{i})$ variables, which can be solved by SCA [11].
+Algorithm 1 consists of solving two optimization subproblems ${\cal P}_{{\bf e}}$ and $\widetilde{{\cal S}}_{{\bf w}}$ at each iteration.
+The computational complexity for solving each subproblem is discussed below:
+Problem ${\cal P}_{\bf e}$: As mentioned in Section IV-A, using the optimal beamforming structure ${\bf w}$ in (14), ${\cal P}_{{\bf e}}$ is converted into the weight optimization problem w.r.t.
+$\{{\bf a}_{i}\}$ with $K_{\text{tot}}(=\sum_{i=1}^{G}K_{i})$ variables, which can be solved by SCA [20].
 If each SCA subproblem is solved using a typical interior point method [32] by the standard convex solver, the computational complexity is $O(K_{\text{tot}}^{3})$ per SCA iteration, which does not grow with either $M$ or $N$.
-We point out that this computational complexity can be further reduced by using the fast ADMM algorithm proposed recently in [26], which only involves closed-form updates and has the computational complexity in the order of $O(K_{\text{tot}}^{2})$.
-Problem ${\cal S}_{{\bf w}}$
-The problem is solved using PSA updates in (17) and (18) per iteration.
-Obtaining ${\bf y}^{(j)}$ in (17) involves computing $\textrm{SINR}_{ik}$ in (2) for all $K_{\text{tot}}$ users and finding the maximum $\phi_{ik}({\bf e})$, which requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+4)$ flops.
+We can adopt low-complexity fast ADMM-based algorithm [22], which only involves closed-form updates and has the computational complexity in the order of $O(K_{\text{tot}}^{2})$.
+Problem ${\cal S}_{{\bf w}}$: The problem is solved using PSA updates in (18) and (19) per iteration.
+Obtaining ${\bf y}^{(j)}$ in (18) requires computing $\textrm{SINR}_{ik}$ in (1) for all $K_{\text{tot}}$ users and finding the maximum $\phi_{ik}({\bf e})$ in (15), which need $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+4)$ flops.
 The required computation in (19) is mainly at computing the gradient $\nabla_{{\bf e}}f({\bf e},{\bf y}^{(j)})$ in (IV-B2).
@@
 Thus, the entire updates per iteration requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+N+1)+2GM+2M$ flops, with the leading complexity being $K_{\text{tot}}N(G+M)$ flops.
-As typically we expect $M\gg G$, the leading computational complexity is $K_{\text{tot}}NM$.
-Thus, the complexity of PSA grows linearly over $M$, $N$, and $K_{\text{tot}}$.
+Since we expect $M\gg G$ in practical scenarios, the leading computational complexity is $K_{\text{tot}}NM$.
+Thus, the complexity of PSA for solving ${\cal S}_{\bf w}$ grows linearly over $M$, $N$, and $K_{\text{tot}}$.
 The complexity of the final per-element projection in (20) is $O(M)$.
-Thus, the overall computational complexity in each AO iteration of Algorithm 1 is as low as $O(K_{\text{tot}}^{2}+K_{\text{tot}}MN)$, which is linear in both $M$ and $N$.
+The overall computational complexity in each AO iteration of Algorithm 1 is $O(K_{\text{tot}}^{2})+\text{const}\cdot K_{\text{tot}}MN$, which is linear in both $M$ and $N$ and quadratic in $K_{\text{tot}}$.
 IV-C3 Convergence
@@
 Thus, the alternating procedure is guaranteed to converge.
-V Generalization to RIS-Assisted Multi-user Downlink Beamforming
-Problem ${\cal P}_{o}$ in fact represents the general formulation of downlink RIS-assisted multi-user beamforming QoS problem that includes both multicast and unicast transmission scenarios.
-Unicast transmission is the most commonly considered scenario in the downlink, where the BS sends data dedicated to each user.
-It is a special case of multicast beamforming when $K=1$ for each group, and the multicast beamforming problem is reduced to the conventional unicast beamforming for $G$ users.
-Thus, our proposed AMBF algorithm in Algorithm 1 can be directly used to solve the RIS-assisted multi-user unicast beamforming problem.
-More importantly, we note that the transformed problem ${\cal P}_{2}$ in (III-A) reflects the general RIS-assisted downlink beamforming problem for either unicast or multicast transmissions.
-As a result, the AMBF approach in Section III-B to alternatingly solve the BS beamforming problem ${\cal P}_{\bf e}$ and the RIS passive beamforming problem ${\cal S}_{\bf w}$ is applicable in both scenarios.
-So far, we have focused on the multicast beamforming problem ${\cal P}_{o}$ for $K>1$, which is a more challenging problem to solve than the unicast beamforming problem.
-For the unicast scenario, the RIS beamforming problem ${\cal S}_{\bf w}$ is still a multicast beamforming MMF problem as discussed in Remark 1 of Section III-B2, which can be efficiently solved by our proposed algorithm in Section IV-B.
-On the other hand, ${\cal P}_{\bf e}$ is reduced to the conventional BS multi-user beamforming problem w.r.t.
-${\bf w}$ [27], which yields a closed-form optimal solution as briefly described below.
-For $K=1$, we directly denote the single user in group $i\in{\cal G}$ as user $i$, with a total of $G$ users.
-The optimal beamforming solution ${\bf w}_{i}$ in (13) still holds as a special case, where the equivalent channel matrix $\widetilde{{\bf H}}_{i}$ for group $i$ is reduced to the equivalent channel $\tilde{{\bf h}}_{i}$ for user $i$, and weight vector ${\bf a}_{i}$ now becomes a scalar, denoted as $a_{i}$, i.e.,
+IV-D Special Case $K_{i}=1$
+When $K_{i}=1$, $\forall i$, there is only one user in each group. Problem ${\cal P}_{o}$ is reduced to the RIS-assisted unicast scenario, where the BS transmits dedicated data to each of $G$ users
+For this unicast scenario, the RIS passive beamforming problem ${\cal S}_{\bf w}$ is still a multicast beamforming MMF problem as in Section III-B2, which can be efficiently solved by the proposed PSA-based algorithm in Section IV-B2.
+The beamforming problem ${\cal P}_{\bf e}$ is reduced to the conventional downlink multi-user beamforming problem w.r.t.
+${\bf w}$, for which a closed-form optimal solution can be obtained [33, 20].
+In particular, SINR target for user $i$ is $\gamma_{i}$; for the optimal beamforming solution ${\bf w}_{i}$ in (14), the equivalent channel matrix $\widetilde{{\bf H}}_{i}$ for group $i$ reduces to an equivalent channel $\tilde{{\bf h}}_{i}$ for user $i$, and the weight vector ${\bf a}_{i}$ becomes a scalar, denoted as $a_{i}$.
+Then, we have
 ${\bf w}_{i}=a_{i}{\bf R}^{-1}(\boldsymbol{\lambda})\tilde{{\bf h}}_{i},\quad i\in{\cal G}.$ (22)
-The weights $\{a_{i}\}$ for all $G$ users can be derived in closed-form.
-This is because when the transmit beamformer can be individually designed for each user, at the optimality, all SINR constraints in (5b) is attained with equality.
-Specifically, let ${\bf c}_{i}\triangleq{\bf R}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf h}}_{i1}$, $i\in{\cal G}$, and we have the following the SINR equality constraint
-$\frac{a_{i}^{2}}{\gamma_{i1}}|{\bf c}_{i}^{H}{\bf h}_{i1}|^{2}=\!\!\sum_{j\in{\cal G},j\neq i}\!\!a_{j}^{2}|{\bf c}_{j}^{H}{\bf h}_{j1}|^{2}+\sigma^{2},\quad\text{for}~{}i\in{\cal G}.$ (22)
-Thus, $\{a_{i}^{2}\}$ can be obtained by solving the above $G$ linear equations. Define the $G\times G$ matrix ${\bf V}$ with its $ij$-th entry as
-$[{\bf V}]_{ij}=\begin{cases}\frac{1}{\gamma_{i1}}|{\bf c}_{i}^{H}{\bf h}_{i1}|^{2},&j=i\\ -|{\bf c}_{i}^{H}{\bf h}_{j1}|^{2},&j\neq i.\end{cases}$ (23)
-Define $\tilde{{\bf a}}=[a_{1}^{2},...,a_{G}^{2}]^{T}$. Then, the solution to the $G$ linear equations in (22) can be compactly written as
-$\tilde{{\bf a}}=\sigma^{2}{\bf V}^{-1}\mathbf{1}$ (24)
-where $\mathbf{1}$ is an all-one vector.
-Finally, without the loss of optimality, we obtain
-$a_{i}=\sqrt{[\tilde{{\bf a}}]_{i}},\quad i\in{\cal G}.$ (25)
-Thus, the joint optimization of $({\bf w},{\bf e})$ for ${\cal P}_{o}$ by the AMBF approach involves alternating between computing the closed-form solution ${\bf w}_{i}$ in (21) and solving $\widetilde{{\cal S}}_{{\bf w}}$ for ${\bf e}$ using the updates in (17) and (18).
-Remark 3.
-RIS-assisted downlink multi-user beamforming design for the unicast scenario has been studied in many recent works.
-However, only a few of them considered the QoS problem for transmit power minimization [5, 21].
-Perhaps partially this is due to the difficulty in using AO effectively to solve the joint BS and RIS beamforming problem.
-In [5], directly applying AO to the joint BS and RIS beamforming to minimize the BS transmit power is considered.
-As discussed at the beginning of Section III, direct AO renders the RIS beamforming a feasibility problem, which can not provide an effective solution.
-Thus, some modification to the feasibility problem is proposed to find a better solution, and a low-complexity heuristic method based on the sum of channel gains is also proposed.
-The joint beamforming design for RIS-assisted non-orthogonal multiple access (NOMA) networks is considered in [21], where the direct AO approach is also adopted, with the RIS beamforming being treated as a feasibility problem.
-Both [5] and [21] adopt SDP or SDR to solve the corresponding alternating subproblems.
-In our work, we first transform the QoS problem into ${\cal P}_{2}$ in (III-A).
-It allows us to break the problem into two subproblems, for which AO is effective in obtaining a good solution.
-The subproblems can be solved by the closed-form solution in (21) and the fast algorithm in Section IV-B, which is a scalable solution for large-scale systems.
-VI RIS-Assisted Multicast MMF Problem
-We now consider the weighted max-min fair (MMF) design for the RIS-assisted multicast scenario.
+The weights $a_{i}$’s for all $G$ users can be derived in closed-form in this unicast case.
+This is because when the transmit beamformer can be individually designed for each user, at the optimality, all SINR constraints in (4b) can be attained with equality.
+Denote ${\bf c}_{i}\triangleq{\bf R}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf h}}_{i}$, $i\in{\cal G}$.
+Then, we have the following SINR equality constraints
+$\frac{a_{i}^{2}}{\gamma_{i}}|{\bf c}_{i}^{H}{\bf h}_{i}|^{2}=\!\!\sum_{j\in{\cal G},j\neq i}\!\!a_{j}^{2}|{\bf c}_{j}^{H}{\bf h}_{i}|^{2}+\sigma^{2},\quad\text{for}~{}i\in{\cal G}.$ (23)
+Solving the above $G$ linear equations, we obtain $a_{i}^{2}$, for $i\in{\cal G}$. In a matrix representation, define a $G\times G$ matrix ${\bf V}$ with its $ij$-th entry as
+$[{\bf V}]_{ij}=\begin{cases}\frac{1}{\gamma_{i}}|{\bf c}_{i}^{H}{\bf h}_{i}|^{2},&j=i\\ -|{\bf c}_{j}^{H}{\bf h}_{i}|^{2},&j\neq i.\end{cases}$ (24)
+Then, the solution to the $G$ linear equations in (23) can be compactly written as
+$\tilde{{\bf a}}^{\text{\tiny sq}}=\sigma^{2}{\bf V}^{-1}\mathbf{1}$ (25)
+where $\tilde{{\bf a}}^{\text{\tiny sq}}=[a_{1}^{2},...,a_{G}^{2}]^{T}$, and $\mathbf{1}$ is an all-one vector.
+Finally, without the loss of optimality, we have $a_{i}=\sqrt{[\tilde{{\bf a}}]_{i}}$, $i\in{\cal G}$.
+In summary, in this special unicast case, the AMBF approach for solving the joint optimization of $({\bf w},{\bf e})$ for ${\cal P}_{o}$ involves alternating between computing the closed-form solution ${\bf w}_{i}$ in (22) and solving $\widetilde{{\cal S}}_{{\bf w}}$ for ${\bf e}$ using the updates in (18) and (19), which is a computationally efficient solution.
+V RIS-Assisted Multicast MMF Problem
+Another commonly considered multicast beamforming formulation is the weighted MMF problem.
 The objective is to maximize the minimum weighted SINR among users, subject to the BS transmit power budget and the unit-modulus constraint for the RIS elements.
-The weighted SINR is the ratio between SINR and a pre-specified SINR target, which sets certain fairness among users.
 This joint optimization problem w.r.t.
@@
 $\ \ |e_{m}|^{2}=1,\ m\in{\cal M}.$ (26c)
-where $P$ is the transmit power budget. Similar to the QoS problem ${\cal P}_{o}$, problem ${\cal S}_{o}$ is also non-convex and NP-hard.
-For the conventional downlink multicast beamforming design at the BS, it is known that the QoS and MMF problems are inverse problems [28]: the role of the objective function and the constraints are switched in the two problems, and we can solve the MMF problem by iteratively solving the QoS problem along with a bi-section search over the minimum SINR target for the QoS problem until the total transmit power meets the power budget in the MMF problem.
-Naturally, we want to know whether such a relation also holds for the RIS-assisted multicast beamforming design, which now also includes the RIS passive beamformer optimization, in addition to the BS transmit beamformers.
-We show below that this is indeed the case.
+where $P$ is the maximum transmit power limit, and the weight $\gamma_{ik}$ is a pre-specified SINR target, which can be used to set certain fairness among users.
+Similar to the QoS problem ${\cal P}_{o}$, the MMF problem ${\cal S}_{o}$ is non-convex NP-hard.
+We are interested in the relationship between ${\cal P}_{o}$ and ${\cal S}_{o}$.
+Note that the RIS-assisted multicast beamforming design includes optimization of the RIS passive beamformer in addition to the BS transmit beamformers.
+We show below that these two problems are inverse problems.
 Proposition 1.
 For RIS-assisted multicast beamforming, the QoS problem ${\cal P}_{o}$ and the MMF problem ${\cal S}_{o}$ are inverse problems.
-In particular, we explicitly parameterize ${\cal S}_{o}$ as ${\cal S}_{o}(\boldsymbol{\gamma},P)$ for given $\boldsymbol{\gamma}$ and $P$ with the maximum objective value denoted as $t^{o}={\cal S}_{o}(\boldsymbol{\gamma},P)$, and we also parameterize ${\cal P}_{o}$ as ${\cal P}_{o}(\boldsymbol{\gamma})$ with $\boldsymbol{\gamma}$ being the SINR target vector, and the minimum power obtained as $P={\cal P}_{o}(\boldsymbol{\gamma})$.
+Specifically, let $\boldsymbol{\gamma}$ be the vector containing all $\gamma_{ik}$’s.
+Parameterize ${\cal S}_{o}$ as ${\cal S}_{o}(\boldsymbol{\gamma},P)$ for given $\boldsymbol{\gamma}$ and $P$ with the maximum objective value denoted as $t^{o}={\cal S}_{o}(\boldsymbol{\gamma},P)$, and parameterize ${\cal P}_{o}$ as ${\cal P}_{o}(\boldsymbol{\gamma})$ with $\boldsymbol{\gamma}$ being the SINR target vector, with the minimum power objective value as $P={\cal P}_{o}(\boldsymbol{\gamma})$.
 Then, ${\cal P}_{o}(\boldsymbol{\gamma})$ and ${\cal S}_{o}(\boldsymbol{\gamma},P)$ have the following relations:
 $t^{o}$ $={\cal S}_{o}(\boldsymbol{\gamma},{\cal P}_{o}(t^{o}\boldsymbol{\gamma})),\quad P={\cal P}_{o}({\cal S}_{o}(\boldsymbol{\gamma},P)\boldsymbol{\gamma}).$ (27)
-Proof:
-See Appendix A.
-Based on the above relation, we can solve the MMF problem ${\cal S}_{o}$ using Algorithm 1.
+Proof: See Appendix A.
+Based on Proposition 1, we can solve the MMF problem ${\cal S}_{o}$ using Algorithm 1.
 In particular, by the relation in (27), ${\cal S}_{o}$ can be solved by solving its inverse QoS problem ${\cal P}_{o}$ iteratively along with a bi-section search for $t^{o}$.
-However, since the size of the RIS reflection elements is typically large $M\gg 1$, iteratively solving ${\cal P}_{o}$ will incur relatively high computational complexity.
-We need a more computationally efficient algorithm to solve the large-scale joint optimization problem ${\cal S}_{o}$.
-Aiming at this goal, we present a fast algorithm to solve ${\cal S}_{o}$ directly.
-VI-A Proposed Fast Algorithm for Solving ${\cal S}_{o}$
-For the conventional downlink scenario without RIS, as mentioned in Section IV-B, PSA is proposed in [12] to obtain a near-stationary solution for the multicast beamforming MMF problem.
-Below, we show that we can adopt PSA to solve the MMF problem efficiently for the RIS-assisted multicast scenario directly as well.
-From Proposition 1, problems ${\cal P}_{o}$ and ${\cal S}_{o}$ are the inverse problems.
-This means the optimal structure of the BS multicast beamformer ${\bf w}$ still has the form given in (13).
-Indeed, with given ${\bf e}$, if we treat the RIS-assisted channel between the BS and user $k$ in group $i$ as the equivalent channel $\tilde{{\bf h}}_{ik}$, ${\cal S}_{o}$ reduces to the conventional multicast MMF problem, for which the optimal beamforming structure is (13) [11].
-However, the challenge for the MMF problem is that computing $\boldsymbol{\lambda}$ in ${\bf R}(\boldsymbol{\lambda})$ is not straightforward.
-Fortunately, the asymptotic expression of ${\bf R}(\boldsymbol{\lambda})$ for large $N$ is obtained in closed form in [11], which can be used as an approximate expression of ${\bf R}(\boldsymbol{\lambda})$ to further simplify the computation.
-Specifically, let $\tilde{{\bf h}}_{iK}=\sqrt{\tilde{\beta}_{ik}}{\bf g}_{ik}$ where $\tilde{\beta}_{ik}$ represents the large-scale channel variance and $\tilde{{\bf g}}_{ik}\sim{\cal C}{\cal N}(\mathbf{0},{\bf I})$.
-Then, the asymptotic expression of ${\bf R}(\boldsymbol{\lambda})$ has the following simple closed form
-${\bf R}(\boldsymbol{\lambda})\approx{\bf I}+\frac{P\bar{\beta}_{{\bf h}}}{\sigma^{2}K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}{\bf g}_{ik}{\bf g}_{ik}^{H}\triangleq\widetilde{{\bf R}}$ (28)
+However, since the size of the RIS reflection elements $M$ is large, this indirect approach by iteratively solving ${\cal P}_{o}$ can be computationally inefficient.
+In the following, we propose a more efficient algorithm to solve ${\cal S}_{o}$ directly.
+In Section IV-B, PSA is proposed to obtain ${\bf e}$ for the RIS multicast beamforming MMF problem.
+Below, we propose an alternating PSA (APSA) approach to directly compute a solution for ${\cal S}_{o}$.
+V-A Proposed Fast Algorithm
+Since problems ${\cal P}_{o}$ and ${\cal S}_{o}$ are the inverse problems, it implies that the optimal structure of the BS multicast beamformer ${\bf w}$ still has the form given in (14).
+Indeed, using the equivalent channel $\tilde{{\bf h}}_{ik}$ between the BS and the user, the BS mutlicast beamforming is the same as the conventional downlink multicast MMF problem, for which the optimal beamforming structure is [20]
+${\bf w}_{i}=\widetilde{{\bf R}}^{-1}(\boldsymbol{\lambda})\widetilde{{\bf H}}_{i}{\bf a}_{i}$ (28)
+where $\widetilde{{\bf R}}={\bf{I}}+\frac{P}{\sigma^{2}}\sum_{i=1}^{G}\sum_{k=1}^{K}\frac{\lambda_{ik}\gamma_{ik}}{\boldsymbol{\lambda}^{T}\boldsymbol{\gamma}}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}$.
+However, computing $\boldsymbol{\lambda}$ in $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ is not straightforward for the MMF problem.
+Fortunately, the asymptotic expression of $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ for large $N$ has been obtained in closed-form [20].
+We can use it as an approximate expression of $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ to further simplify the computation.
+Specifically, we rewrite $\tilde{{\bf h}}_{iK}=\sqrt{\tilde{\beta}_{ik}}{\bf g}_{ik}$ where $\tilde{\beta}_{ik}$ represents the large-scale channel variance and $\tilde{{\bf g}}_{ik}\sim{\cal C}{\cal N}(\mathbf{0},{\bf I})$.
+Then, the asymptotic expression of $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ has the following simple closed form
+$\widetilde{{\bf R}}(\boldsymbol{\lambda})\approx{\bf I}+\frac{P\bar{\beta}_{{\bf h}}}{\sigma^{2}K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}{\bf g}_{ik}{\bf g}_{ik}^{H}\triangleq\widetilde{{\bf R}}^{\text{\tiny asym}}$ (29)
 where $\bar{\beta}_{{\bf h}}\triangleq 1/\Big{(}\frac{1}{K_{\text{tot}}}\sum_{i=1}^{G}\sum_{k=1}^{K_{i}}\frac{1}{\tilde{\beta}_{ik}}\Big{)}$ is the harmonic mean of the large-scale channel variance of all users.
-Using $\widetilde{{\bf R}}$ in (28), we can replace ${\bf w}_{i}$ in (13) with the following expression:
-${\bf w}_{i}\,\approx\,\widetilde{{\bf R}}^{-1}\widetilde{{\bf H}}_{i}{\bf a}_{i}\,,\;i\in{\cal G}.$ (29)
-Below, we develop our fast algorithm to solve ${\cal S}_{o}$ based on ${\bf w}_{i}$ in (29) to reduce the computational complexity.
-Using ${\bf w}_{i}$ in (29), we first convert ${\cal S}_{o}$ into a joint optimization problem w.r.t. $\{{\bf a}_{i}\}$ and ${\bf e}$, given by
+Using $\widetilde{{\bf R}}(\boldsymbol{\lambda})$ in (29), we can replace ${\bf w}_{i}$ in (14) with the following expression:
+${\bf w}_{i}\approx(\widetilde{{\bf R}}^{\text{\tiny asym}})^{-1}\widetilde{{\bf H}}_{i}{\bf a}_{i}\,,\;i\in{\cal G}.$ (30)
+We use ${\bf w}_{i}$ in (30) to develop a low-complexity fast algorithm for ${\cal S}_{o}$.
+We first convert ${\cal S}_{o}$ into a joint optimization problem w.r.t.
+$\{{\bf a}_{i}\}$ and ${\bf e}$, given by
 ${\cal S}_{o}^{\text{eq}}:\max_{{\bf a},{\bf e}}$ $\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ (31a)
@@
 $\ \ {\bf e}^{T}{\bf e}=M.$ (31d)
-where ${\bf a}\triangleq[{\bf a}_{1}^{H},\cdots,{\bf a}_{G}^{H}]^{H}$, $\widetilde{{\bf C}}_{i}\triangleq\widetilde{{\bf R}}^{-1}\widetilde{{\bf H}}_{i}$, and the constraints in (26c) are replaced by the equivalent set of constraints in (30c) and (30d) based on (9).
-Similar to $\widetilde{{\cal S}}_{{\bf w}}$ in (III-B2), we relax $S_{o}^{\text{eq}}$ into the following problem by moving the equality constraint in (30d) into the objective function as a penalty term with a penalty weight $\delta>0$:
-$\widetilde{{\cal S}}_{o}:\max_{{\bf a},{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\delta\frac{{\bf e}^{H}{\bf e}}{M}$ (31a)
+where $\widetilde{{\bf C}}_{i}\triangleq(\widetilde{{\bf R}}^{\text{\tiny asym}})^{-1}\widetilde{{\bf H}}_{i}$, ${\bf a}\triangleq[{\bf a}_{1}^{H},\cdots,{\bf a}_{G}^{H}]^{H}$, and the constraints in (26c) are replaced with the equivalent set of constraints in (31c) and (31d) based on (10).
+Using the similar relaxation technique from ${\cal S}_{{\bf w}}^{\text{eq}}$ to $\widetilde{{\cal S}}_{{\bf w}}$ in (12), we relax $S_{o}^{\text{eq}}$ into the following problem by moving the equality constraint in (31d) into the objective function as a penalty term with a penalty weight $\tilde{\rho}>0$:
+$\widetilde{{\cal S}}_{o}:\max_{{\bf a},{\bf e}}$ $\ \bigg{(}\!\min_{i,k}\ \ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}+\tilde{\rho}\frac{{\bf e}^{H}{\bf e}}{M}$ (32a)
 s.t. $\ \ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P,$ (32b)
 $\ \ |e_{m}|^{2}\leq 1,\ m\in{\cal M}.$ (32c)
-Problem $\widetilde{{\cal S}}_{o}$ is similar to $\widetilde{{\cal S}}_{\bf w}$ in (III-B2), except that $\widetilde{{\cal S}}_{o}$ is a joint optimization problem for $({\bf a},{\bf e})$ while $\widetilde{{\cal S}}_{\bf w}$ is for ${\bf e}$ only.
-Nonetheless, we can still apply PSA to compute a near-stationary joint solution $({\bf a},{\bf e})$ for $\widetilde{{\cal S}}_{o}$ efficiently, similar to Section IV-B, as shown below.
-VI-A1 Problem Reformulation
-Let ${\bf x}\triangleq[{\bf a}^{H},{\bf e}^{H}]^{H}$. By substituting the SINR expression in (2) into the objective function in (31a), we define
-$\varphi_{ik}({\bf x})\triangleq$ $-\!\frac{1}{\gamma_{ik}}\frac{|{\bf a}_{i}^{H}\widetilde{{\bf C}}_{i}^{H}\!({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})|^{2}}{\sum_{j\neq i}|{\bf a}_{j}^{H}\widetilde{{\bf C}}_{j}^{H}\!({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})|^{2}\!+\!\sigma^{2}}-\delta\frac{{\bf e}^{H}{\bf e}}{M},$ (32)
-for $k\in{\cal K}_{i}$, $i\in{\cal G}$.
-Following the similar arguments from (14)–(16), let $\boldsymbol{\varphi}({\bf x})$ be a $K_{\text{tot}}\times 1$ vector containing all $\varphi_{ik}({\bf x})$’s, and let ${\bf y}$ denote a $K_{\text{tot}}\times 1$ vector with ${\bf y}\succcurlyeq 0$ and ${\bf 1}^{T}{\bf y}=1$.
+We note that problem $\widetilde{{\cal S}}_{o}$ is structurally similar to $\widetilde{{\cal S}}_{\bf w}$ in (12), except that $\widetilde{{\cal S}}_{o}$ is a joint optimization problem for $({\bf a},{\bf e})$ while $\widetilde{{\cal S}}_{\bf w}$ is for ${\bf e}$ only.
+Thus, we propose an APSA approach to alternatingly apply PSA to compute a solution $({\bf a},{\bf e})$ for $\widetilde{{\cal S}}_{o}$.
+V-A1 Reformulation
+We first convert $\widetilde{{\cal S}}_{o}$ into a min-max problem. Substituting the SINR expression in (1) into the objective function in (32a), for each $k\in{\cal K}_{i}$, $i\in{\cal G}$, we define
+$\varphi_{ik}({\bf a},{\bf e})\triangleq$ $-\!\frac{1}{\gamma_{ik}}\frac{|({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})^{H}\widetilde{{\bf C}}_{i}{\bf a}_{i}|^{2}}{\sum_{j\neq i}|({\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e})^{H}\widetilde{{\bf C}}_{j}{\bf a}_{j}|^{2}\!+\!\sigma^{2}}-\tilde{\rho}\frac{{\bf e}^{H}{\bf e}}{M},$ (33)
+Following the similar arguments from (15)–(17), let $\boldsymbol{\varphi}({\bf a},{\bf e})$ be a $K_{\text{tot}}\times 1$ vector containing all $\varphi_{ik}({\bf a},{\bf e})$’s, and let ${\bf y}$ denote a $K_{\text{tot}}\times 1$ vector with ${\bf y}\succcurlyeq 0$ and ${\bf 1}^{T}{\bf y}=1$.
 Then, $\widetilde{{\cal S}}_{o}$ can be transform into the following equivalent problem
-$\widetilde{{\cal S}}_{o}^{\prime}:\min_{{\bf x}\in{\cal U}}\max_{{\bf y}\in{\cal Y}}$ $\ g({\bf x},{\bf y})$ (33)
-where $g({\bf x},{\bf y})\triangleq\boldsymbol{\varphi}^{T}({\bf x}){\bf y}$, ${\cal Y}\triangleq\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$, and ${\cal U}\triangleq\{{\bf x}:|e_{m}|^{2}\leq 1,\ m\in{\cal M};\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\}$.
-Both ${\cal Y}$ and ${\cal U}$ are compact convex sets.
-VI-A2 The Projected Subgradient Algorithm
-Problem $\widetilde{{\cal S}}_{o}^{\prime}$ is a non-convex-concave min-max problem, whose structure is the same as $\widetilde{{\cal S}}_{\bf w}^{\prime}$ in (16).
-Thus, we can apply PSA with the updating steps similar to (17) and (18).
-In particular, the updating procedure at iteration $j$ is given by
-${\bf y}^{(j)}\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}g({\bf x}^{(j)},{\bf y});$ (34)
-${\bf x}^{(j+1)}=\Pi_{{\cal U}}\big{(}{\bf x}^{(j)}-\eta\nabla_{{\bf x}}g({\bf x}^{(j)},{\bf y}^{(j)})\big{)}$ (35)
-where $\eta>0$ is the step size, and $\Pi_{{\cal U}}({\bf x})$ is the projection of ${\bf x}$ onto the feasible set ${\cal U}$.
-Since the constraints on ${\bf a}$ and ${\bf e}$ in ${\cal U}$ are separate, $\Pi_{{\cal U}}({\bf x})$ can be performed for ${\bf a}$ and ${\bf e}$ separately:
-$\Pi_{{\cal U}}({\bf x})=\begin{cases}{\bf a}&\mbox{if }\sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\\ \sqrt{\frac{P}{P_{\text{tot}}}}{\bf a}&\text{o.w.}\\ e_{m}&\mbox{if }|e_{m}|\leq 1,\ m\in{\cal M}\\ \frac{e_{m}}{|e_{m}|}&\text{o.w.}\end{cases}.$ (36)
+$\widetilde{{\cal S}}_{o}^{\prime}:\min_{{\bf a}\in{\cal A},{\bf e}\in{\cal E}}\max_{{\bf y}\in{\cal Y}}$ $\ \tilde{f}({\bf a},{\bf e},{\bf y})$ (34)
+where $\tilde{f}({\bf a},{\bf e},{\bf y})\triangleq\boldsymbol{\varphi}^{T}({\bf a},{\bf e}){\bf y}$, ${\cal A}\triangleq\{{\bf a}:\sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\}$, ${\cal Y}=\{{\bf y}:{\bf y}\succcurlyeq 0,\mathbf{1}^{T}{\bf y}=1\}$, and =${\cal E}=\{{\bf e}:|e_{m}|^{2}\leq 1,\ m\in{\cal M}\}$.
+All ${\cal A}$, ${\cal E}$, and ${\cal Y}$ are compact convex sets.
+V-A2 Alternating Projected Subgradient Algorithm
+Similar to the updating steps in (18) and (19) to solve $\widetilde{{\cal S}}_{\bf w}^{\prime}$ in (17), we can apply PSA to update ${\bf a}$ and ${\bf e}$ alternatingly.
+In particular, the updating procedure at iteration $n$ is given by
+${\bf y}^{(n)}\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}\tilde{f}({\bf a}^{(n)},{\bf e}^{(n)},{\bf y});$ (35)
+${\bf e}^{(n+1)}=\Pi_{{\cal E}}\big{(}{\bf e}^{(n)}-\tilde{\alpha}\nabla_{{\bf e}}\tilde{f}({\bf a}^{(n)},{\bf e}^{(n)},{\bf y}^{(n)})\big{)}$ (36)
+$\tilde{{\bf y}}^{(n)}\in\displaystyle\mathop{\arg\max}_{{\bf y}\in{\cal Y}}\tilde{f}({\bf a}^{(n)},{\bf e}^{(n+1)},{\bf y});$ (37)
+${\bf a}^{(n+1)}=\Pi_{{\cal A}}\big{(}{\bf a}^{(n)}-\tilde{\alpha}\nabla_{{\bf a}}\tilde{f}({\bf a}^{(n)},{\bf e}^{(n+1)},\tilde{{\bf y}}^{(n)})\big{)}$ (38)
+where $\tilde{\alpha}>0$ is the step size, $\Pi_{{\cal E}}({\bf e})$ is the projection of ${\bf e}$ onto ${\cal E}$ as in (20), and $\Pi_{{\cal A}}({\bf x})$ is the projection of ${\bf a}$ onto the feasible set ${\cal A}$, given by
+$\Pi_{{\cal A}}({\bf x})=\begin{cases}{\bf a}&\mbox{if }\sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P\\ \sqrt{\frac{P}{P_{\text{tot}}}}{\bf a}&\text{o.w.}\end{cases}.$ (39)
 Since ${\cal Y}$ is a probability simplex, an optimal solution to the maximization problem in (35) is ${\bf y}=[0,\cdots,1,\cdots,0]^{T}$, with 1 at the position that corresponds to $\max_{i,k}\varphi_{ik}({\bf e})$, i.e.,
-$g({\bf x},{\bf y}^{(j)})=\varphi_{\hat{i}\hat{k}}({\bf x})=\max_{i,k}\varphi_{ik}({\bf x}),$ (37)
-for some $(\hat{i},\hat{k})$. Then, the gradient $\nabla_{{\bf x}}g({\bf x},{\bf y}^{(j)})$ is given by
-$\nabla_{{\bf x}}g({\bf x},{\bf y}^{(j)})$ $=\nabla_{{\bf x}}\varphi_{\hat{i}\hat{k}}({\bf x})=\Bigl{[}\nabla_{{\bf a}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})\!,\nabla_{{\bf e}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})\Bigl{]}^{H}.$ (38)
+$\tilde{f}({\bf a},{\bf e},{\bf y}^{(n)})=\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})=\max_{i,k}\varphi_{ik}({\bf a},{\bf e}),$ (40)
+for some pair $(\hat{i},\hat{k})$. Thus, the gradient $\nabla_{{\bf x}}\tilde{f}({\bf a},{\bf e},{\bf y}^{(n)})$, for ${\bf x}={\bf a}$ or ${\bf e}$, is given by
+$\nabla_{{\bf x}}\tilde{f}({\bf a},{\bf e},{\bf y}^{(n)})$ $=\!\nabla_{{\bf x}}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e}),\quad\text{for~{}}{\bf x}={\bf a}\text{~{}or~{}}{\bf e}.$ (41)
 For $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$, note that the expression for $\varphi_{ik}({\bf x})$ is the same as $\phi_{ik}({\bf e})$.
-Thus, the gradient $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ can be computed using (IV-B2), except that ${\bf w}_{i}$ is computed using (29), and $\zeta$ is replaced by $\delta$.
-To compute $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$, we rewrite $\varphi_{ik}({\bf x})$ in (32) as
-$\varphi_{ik}({\bf x})=$ $-\frac{1}{\gamma_{ik}}\frac{{\bf a}_{i}^{H}\widetilde{{\bf A}}_{iik}{\bf a}_{i}}{\sum_{j\neq i}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{jik}{\bf a}_{j}+\sigma^{2}}-\delta\frac{{\bf e}^{H}{\bf e}}{M},$
-where $\widetilde{{\bf A}}_{jik}\triangleq\widetilde{{\bf C}}_{j}^{H}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}^{H}\widetilde{{\bf C}}_{j}$, with $\tilde{{\bf h}}_{ik}={\bf h}_{ik}^{\text{d}}\!\!+\!{\bf G}_{ik}{\bf e}$ being the equivalent channel.
-Then, we have
-$\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$ $=[\nabla_{{\bf a}_{1}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x}),\cdots,\nabla_{{\bf a}_{G}}^{H}\varphi_{\hat{i}\hat{k}}({\bf x})]^{H}$ (39)
+Thus, the gradient $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ can be computed using (IV-B2), where ${\bf w}_{i}$ is computed using (30), and $\rho$ is replaced by $\tilde{\rho}$.
+To compute $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})$, we rewrite $\varphi_{ik}({\bf a},{\bf e})$ in (33) as
+$\varphi_{ik}({\bf a},{\bf e})=$ $-\frac{1}{\gamma_{ik}}\frac{{\bf a}_{i}^{H}\widetilde{{\bf A}}_{iik}{\bf a}_{i}}{\sum_{j\neq i}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{jik}{\bf a}_{j}+\sigma^{2}}-\tilde{\rho}\frac{{\bf e}^{H}{\bf e}}{M},$
+where $\widetilde{{\bf A}}_{jik}\triangleq\widetilde{{\bf C}}_{j}^{H}\tilde{{\bf h}}_{ik}\tilde{{\bf h}}_{ik}^{H}\widetilde{{\bf C}}_{j}$. Then, we have
+$\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})$ $=[\nabla_{{\bf a}_{1}}^{H}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e}),\cdots,\nabla_{{\bf a}_{G}}^{H}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})]^{H}$ (42)
 where
-$\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})\!=\!\begin{cases}\displaystyle-\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}}}{\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}}&\mbox{if }i=\hat{i}\\[15.00002pt] \displaystyle\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{({\bf a}_{\hat{i}}^{H}\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}})\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}}{\bigl{(}\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}\bigl{)}^{2}}&\text{o.w.}\end{cases}$ (40)
-From the above, we note that the updates in (34) and (35) can be computed using closed-form evaluations.
-VI-A3 Final Processing
-Since we are solving $\widetilde{{\cal S}}_{o}$, which is a relaxed problem of ${\cal S}_{o}$, the computed solution ${\bf e}^{\star}$ may not satisfy (26c) for ${\cal S}_{o}$.
-Thus, we require this final processing step to obtain a feasible solution to ${\cal S}_{o}$.
-Let ${\bf x}^{\star}=[{{\bf a}^{\star}}^{H},{{\bf e}^{\star}}^{H}]^{H}$ be the solution produced by the updating procedure in (34) and (35).
+$\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf a},{\bf e})\!=\!\begin{cases}\displaystyle-\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}}}{\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}}&\mbox{if }i=\hat{i}\\[15.00002pt] \displaystyle\frac{1}{\gamma_{\hat{i}\hat{k}}}\frac{({\bf a}_{\hat{i}}^{H}\widetilde{{\bf A}}_{\hat{i}\hat{i}\hat{k}}{\bf a}_{\hat{i}})\widetilde{{\bf A}}_{i\hat{i}\hat{k}}{\bf a}_{i}}{\bigl{(}\sum_{j\neq\hat{i}}{\bf a}_{j}^{H}\widetilde{{\bf A}}_{j\hat{i}\hat{k}}{\bf a}_{j}\!+\!\sigma^{2}\bigl{)}^{2}}&\text{o.w.}\end{cases}$ (43)
+Thus, the updates in (35)–(38) can be computed using closed-form evaluations from (39)–(43).
+V-A3 Final Processing
+Recall that $\widetilde{{\cal S}}_{o}$ is a relaxed problem of ${\cal S}_{o}$.
+Thus, the computed solution ${\bf e}^{\star}$ may not satisfy (26c) for ${\cal S}_{o}$.
+We use this final processing step to obtain a feasible solution to ${\cal S}_{o}$.
+Let $({{\bf a}^{\star}},{{\bf e}^{\star}})$ be the solution produced by the PSA procedure in (35)–(38).
 If ${\bf e}^{\star}$ does not satisfy (26c), we have the following step to obtain a feasible solution based on $({\bf a}^{\star},{\bf e}^{\star})$:
-  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which takes the phase component of each element in ${\bf e}^{\star}$;
-  • Given ${\bf e}^{\text{final}}$, we solve ${\cal S}_{o}^{\text{eq}}$ in (VI-A) w.r.t.
-${\bf a}$ and obtain the final beamforming solution ${\bf w}^{\text{final}}$.
-Note that in this case, ${\cal S}_{o}^{\text{eq}}$ is reduced to the following conventional MMF problem:
-    ${\cal S}^{\text{eq}}({\bf e}^{\text{final}}):\max_{{\bf a}}$ $\min_{i,k}\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ s.t. $\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P.$ (41)
-    where $\textrm{SINR}_{ik}$ is given in (12) with the equivalent channel $\tilde{{\bf h}}_{ik}$ computed based on ${\bf e}^{\text{final}}$.
-As mentioned earlier, the PSA-based fast algorithm for this MMF problem is provided in [12].
-We directly apply this algorithm to compute the solution ${\bf a}^{\text{final}}$ to ${\cal S}^{\text{eq}}({\bf e}^{\text{final}})$.
-For initialization in this algorithm, we use ${\bf a}^{\star}$ as a good initial point to this algorithm to achieve fast convergence and good performance.
+  • Project ${\bf e}^{\star}$ onto its feasible set: ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{\star})$, which obtains the phase component of each element in ${\bf e}^{\star}$;
+  • Given ${\bf e}^{\text{final}}$, we solve ${\cal S}_{o}^{\text{eq}}$ in (31) w.r.t. ${\bf a}$ to obtain the final weight vector ${\bf a}^{\text{final}}$:
+    ${\cal S}^{\text{eq}}({\bf e}^{\text{final}}):$ $\max_{{\bf a}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$
+[…]
+    where $\textrm{SINR}_{ik}$ in (1) is computed based on ${\bf e}^{\text{final}}$.
+This is the conventional MMF problem, which can by solved by the PSA-based fast algorithm [21].
+Note that ${\bf a}^{\star}$ satisfies the power constraint in (V-A3) and thus serves as a good initial point to this algorithm to achieve fast convergence and good performance.
+The final beamforming solution ${\bf w}^{\text{final}}$ is computed using (30) based on ${\bf a}^{\text{final}}$.
 We summarize our proposed fast algorithm for the MMF problem ${\cal S}_{o}$ in Algorithm 2.
-Again, based on the existing convergence analysis [12], the PSA updating procedure in (34) and (35) is guaranteed to converge in finite time to a point at the vicinity of a stationary point for $\widetilde{{\cal S}}_{o}$.
-Algorithm 2 The Proposed Fast Algorithm for RIS-Assisted MMF Problem ${\cal S}_{o}$.
-1:Initialization: Set feasible point ${\bf x}^{(0)}$; Set $n=0$.
+Again, similar to that discussed in Section IV-B3, we track the best solution for $({\bf a},{\bf e})$ with the minimum objective value of $\widetilde{{\cal S}}_{o}^{\prime}$ obtained so far in $n$ iterations: $\tilde{g}_{\min}^{(n)}=\min\{\tilde{g}^{(1)},...,\tilde{g}^{(n)}\}$, where $\tilde{g}^{(n)}=\max_{{\bf y}\in{\cal Y}}f({\bf a}^{(n)},{\bf e}^{(n)},{\bf y})$.
+Since, $\tilde{g}_{\min}^{(j)}$ is monotonically decreasing and is bounded below, we terminate the algorithm when $\tilde{g}_{\min}^{(j)}$ converges.
+Algorithm 2 The APSA Algorithm for RIS-Assisted MMF Problem ${\cal S}_{o}$.
+1:Initialization: Set feasible point $({\bf a}^{(0)},{\bf e}^{(0)})$; Set $\tilde{\rho}$ and $\tilde{\alpha}$; Set $n=0$.
 2:repeat
-3:     Update ${\bf x}^{(n)}$ using (34) and (35) to obtain ${\bf x}^{(n+1)}$.
-4:     Set $n\leftarrow n+1$.
-5:until convergence
-6:Set ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{(n)})$.
-7:Solving ${\cal S}({\bf e}^{\text{final}})$ to obtain ${\bf w}^{\text{final}}$.
-8:return (${\bf w}^{\text{final}}$, ${\bf e}^{\text{final}}$).
-VI-B Discussions
-VI-B1 Initialization
-The proposed updating procedure in (34) and (35) requires an initial point ${\bf x}^{(0)}=[{{\bf a}^{(0)}}^{H},{{\bf e}^{(0)}}^{H}]^{H}$.
-This initial point can be generated randomly, as the projection (35) will ensure that the subsequent points $\{{\bf x}^{(n)}\}$ are feasible.
-However, a good initial point improves the rate of convergence.
+3:     Compute $\varphi_{ik}({\bf a}^{(n)},{\bf e}^{(n)})$, $k\in{\cal K}_{i}$, $i\in{\cal G}$. Obtain pair $(\hat{i},\hat{k})$ as in (40).
+4:     Update ${\bf e}^{(n+1)}$ using (IV-B2) and (20).
+5:     Compute $\varphi_{ik}({\bf a}^{(n)},{\bf e}^{(n+1)})$, $k\in{\cal K}_{i}$, $i\in{\cal G}$. Obtain pair $(\hat{i},\hat{k})$ as in (40).
+6:     Update ${\bf a}^{(n+1)}$ using (43) and (39).
+7:     Set $n\leftarrow n+1$.
+8:until convergence
+9:Set ${\bf e}^{\text{final}}=\exp(j\angle{\bf e}^{(n)})$.
+10:Solve ${\cal S}^{\text{eq}}({\bf e}^{\text{final}})$ to obtain ${\bf a}^{\text{final}}$.
+11:Compute ${\bf w}^{\text{final}}$ using (30).
+12:return (${\bf w}^{\text{final}}$, ${\bf e}^{\text{final}}$).
+V-B Algorithm Discussion
+V-B1 Initialization
+The proposed updating procedure requires an initial point $({{\bf a}^{(0)}},{{\bf e}^{(0)}})$.
+This initial point can be generated randomly, as the projections $\Pi_{{\cal E}}$ and $\Pi_{Ac}$ will ensure that the subsequent points $\{({{\bf a}^{(n)}},{{\bf e}^{(n)}})\}$ are feasible.
+However, a good initial point help accelerate the convergence.
 Thus, instead of a random initial point, we consider the following approach.
 We first generate ${\bf e}^{(0)}$ with a random phase for each element.
-Given ${\bf e}^{(0)}$, ${\cal S}_{o}^{\text{eq}}$ in (VI-A) is reduced to the MMF problem w.r.t.
-${\bf a}$, which is similar to ${\cal S}^{\text{eq}}({\bf e}^{\text{final}})$ in (41).
-It can be further equivalently written as
-${\cal S}^{\text{eq}}({\bf e}^{(0)}):$ $\max_{{\bf a},t}\ t$ s.t. $\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\geq t,\ k\in{\cal K}_{i},i\in{\cal G}$ (42a)
-$\ \sum_{i=1}^{G}\|\widetilde{{\bf C}}_{i}{\bf a}_{i}\|^{2}\leq P.$ (42b)
-Note that any feasible solution to the above problem is feasible to $\widetilde{{\cal S}}_{o}$.
-Thus, we generate ${\bf a}^{(0)}$ in an attempt to solve this MMF problem, which is to solve its inverse problem, i.e., the QoS problem similar to ${\cal P}_{\bf e}$ in (III-B1) for given $t$, along with bi-section search over $t$ [28, 11].
-However, since our goal is only to generate a feasible point ${\bf a}^{(0)}$ for $\widetilde{{\cal S}}_{o}$, we only need to solve the QoS problem [^2] To solve the QoS problem, we use the classical approach of SDR with the Gaussian randomization procedure.
-along with one round of bi-section search over $t$ to obtain the initial point ${\bf a}^{(0)}$.
-VI-B2 Computational Complexity
-For the updates in (34) and (35) at each iteration, the update ${\bf y}^{(j)}$ is obtained in (37) by finding the maximum $\varphi_{ik}({\bf e})$ among $K_{\text{tot}}$ users, which requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+2N+4)$ flops.
-The update in (35) requires to calculate $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$ and $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$.
+Given ${\bf e}^{(0)}$, ${\cal S}_{o}^{\text{eq}}$ in (31) is reduced to the MMF problem w.r.t.
+${\bf a}$, i.e., ${\cal S}^{\text{eq}}({\bf e}^{(0)})$ as in (V-A3) by replacing ${\bf e}^{\text{final}}$ with ${\bf e}^{0}$.
+Note that any feasible solution ${\bf a}$ to ${\cal S}^{\text{eq}}({\bf e}^{(0)})$ is feasible to $\widetilde{{\cal S}}_{o}$.
+Thus, we only need to find a feasible solution.
+To do so, we solve the inverse problem of ${\cal S}^{\text{eq}}({\bf e}^{(0)})$, i.e., the QoS problem similar to ${\cal P}_{\bf e}$ in (6) for a given minimum SINR target $t$, along with one round bi-section search over $t$, to find the transmit power objective of the QoS problem [^3] To solve the QoS problem, we can use the classical approach of SDR with the Gaussian randomization procedure.
+that is less than $P$ and obtain the initial point ${\bf a}^{(0)}$.
+V-B2 Computational Complexity
+For the updates in (35)–(38) at each iteration, the update ${\bf y}^{(j)}$ is obtained in (40) by finding the maximum $\varphi_{ik}({\bf e})$ among $K_{\text{tot}}$ users, which requires $GK_{\text{tot}}(N+1)+K_{\text{tot}}(MN+2N+4)$ flops.
+The updates in (36) and (38) require to calculate $\nabla_{{\bf a}}\varphi_{\hat{i}\hat{k}}({\bf x})$ and $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$.
 Note that $\nabla_{{\bf e}}\varphi_{\hat{i}\hat{k}}({\bf x})$ is computed using (IV-B2), which requires $2GM+2M$ flops as discussed in Section IV-C2.ii).
@@
 Lastly, we need $K_{\text{tot}}+G$ flops to calculate the final value of $\nabla_{{\bf a}_{i}}\varphi_{\hat{i}\hat{k}}({\bf x})$ for all ${\bf a}_{i}$’s.
-In summary, the complexity in obtaining the update in (35) requires $2(G+1)M+K_{\text{tot}}(N+2)+G$ flops.
+In summary, the complexity in obtaining the updates in (36) and (38) require $2(G+1)M+K_{\text{tot}}(N+2)+G$ flops.
 Thus, the overall leading complexity of the updating procedure in each PSA iteration is dominated by computing ${\bf y}^{(j)}$ in (40), which is $GK_{\text{tot}}N+K_{\text{tot}}MN$ flops.
 It grows linearly with $M$ and $N$.
-VI-C General RIS-Assisted Multi-User Downlink Beamforming
-Similar to the discussion in Section V about the generalization to downlink RIS-assisted multi-user beamforming for the QoS power minimization problem, our proposed fast algorithm (Algorithm 2) is directly applicable to the general downlink RIS-assisted multi-user beamforming problem with the max-min fair objective for both unicast and multicast scenarios.
-Specific to the unicast case with $K_{i}=1,\forall i$, the computational complexity per iteration in the proposed iterative algorithm is reduced to $GMN+G^{2}N$ flops, which is computationally cheap.
-For the multi-user unicast scenario, the MMF problem for RIS-assisted downlink beamforming under various system settings has been considered in the recent literature, and various approaches have been proposed [22, 23, 24].
-In [22], for the same setting considered in this work, an iterative algorithm that provides an alternating descent path is proposed to maximize the minimum SINR, and the computational complexity in each iteration as $O((GN)^{3}+M^{4})$.
-In comparison, the computational complexity of our proposed fast algorithm only grows linearly with $G$, $M$, and $N$.
-A different system model for RIS-assisted transmission is considered in [23], where only the RIS path is available and the channel correlation matrix between RIS and users are assumed to be the same.
-Using an approximation of the optimal BS transmit beamformer, the authors propose an algorithm based on projected gradient ascent to obtain the RIS beamformer based on the large-scale channel statistics.
-The computational complexity of the proposed algorithm is in the higher polynomial order of $N$ and $M$.
-RIS-assisted multi-cell coordinated transmit beamforming is considered in [24], where the RIS can adjust both reflecting phases and amplitudes.
-The AO approach is adopted, with two subproblems being solved by the second-order cone programming (SOCP) and SDR.
-An inexact AO approach with SCA is further considered to reduce the computational complexity.
-However, since these proposed approaches are second-order algorithms, their computational complexity is still high.
-VII Simulation Results
+V-B3 Special Case $K_{i}=1$
+When $K_{i}=1$, $\forall i$, problem ${\cal S}_{o}$ reduces to the RIS-assisted multiuser unicast scenario for $G$ users with the max-min fair objective.
+In this case, the computational complexity per iteration in Algorithm 2 reduces to $GMN+G^{2}N$ flops, which grows linearly with $M$ and $N$.
+V-B4 Improved Initialization Method for QoS Problem ${\cal P}_{o}$
+We can adopt the APSA algorithm as an improved initialization method to AMBF (Algorithm 1) for the QoS problem.
+Specifically, following the random phase initialization in Section IV-C1, we have the initial point $({\bf e}^{0},{\bf w}^{0})$.
+We feed this point to APSA with power budget set to $P=\sum_{i=1}^{G}\|{\bf w}_{i}^{0}\|^{2}$.
+The output ${\bf e}^{\text{final}}$ from APSA is then used as the updated initial point for ${\bf e}$ in Algorithm 1: ${\bf e}^{0}\leftarrow{\bf e}^{\text{final}}$.
+Also, as the purpose of APSA here is for providing an initial point, we can set a pre-set number of iterations for Algorithm 2 for the output before waiting for the algorithm converges,
+VI Simulation Results
 We consider a downlink RIS-assisted transmission scenario with a BS and an RIS.
-The locations of the BS and the RIS using the $(x,y,z)$-coordinates in meters are $(0,0,0)$ and $(70,70,0)$, respectively.
+The locations of the BS and the RIS using the $(x,y,z)$-coordinates in meters are $(x_{\text{BS}},y_{\text{BS}},z_{\text{BS}})=(0,0,20)$ and $(x_{\text{RIS}},y_{\text{RIS}},z_{\text{RIS}})=(70,70,20)$, respectively.
 The RIS is a rectangular surface with the array elements placed on the $(y,z)$-plane.
-The users are located randomly within a circle centered at the RIS with a radius of $20$ m on the $(x,y)$-plane.
-We set the target SINR $\gamma_{ik}=\gamma$, $\forall i,k$.
-Unless specified otherwise, our default system setup is given as $N=16$ antennas at the BS, and $M=10\times 10$ reflective elements at the RIS, $G=2$ groups, $K_{i}=K=2$ users, $\forall i$.
-We assume a line-of-sight (LOS) channel between the BS and the RIS and model ${\bf H}_{\text{r}}$ as a Rician fading channel matrix given by ${\bf H}_{\text{r}}=\beta^{\text{BR}}\big{(}\sqrt{\frac{K_{\text{r}}}{1+K_{\text{r}}}}{\bf H}_{\text{r}}^{\text{LOS}}+\sqrt{\frac{1}{1+K_{\text{r}}}}{\bf H}_{\text{r}}^{\text{NLOS}}\big{)}$, where Rician factor $K_{\text{r}}=10$, $\beta^{\text{BR}}$ is the path gain modeled as $\beta^{\text{BR}}\text{[dB]}=-30-22\log 10(d^{\text{B-R}})$ with $d^{\text{B-R}}$ being the BS-RIS distance in meters, ${\bf H}_{\text{r}}^{\text{NLOS}}$ is the non-line-of-sight (NLOS) component modeled as $[{\bf H}_{\text{r}}^{\text{NLOS}}]_{ij}\sim{\cal C}{\cal N}(0,{\bf I})$, and ${\bf H}_{\text{r}}^{\text{LOS}}$ is the LOS component.
+The users are located randomly within a rectangle area $[45,65]\times[75,125]$ on the $(x,y)$-plane, each at location $(x_{ik},y_{ik},0)$.
+We set the target SINR for each user the same: $\gamma_{ik}=\gamma$, $\forall i,k$.
+Unless specified otherwise, our default system setup is given as $N=32$ antennas at the BS, and $M=10\times 10$ reflective elements at the RIS, $G=2$ groups, $K_{i}=K=3$ users, $\forall i$, and SINR target $\gamma=10~{}$dB.
+We consider the system operates at carrier frequency $f_{c}=3.5$ GHz.
+We assume a line-of-sight (LOS) channel between the BS and the RIS and model ${\bf H}_{\text{r}}$ as a Rician fading channel matrix given by ${\bf H}_{\text{r}}=\beta^{\text{BR}}\big{(}\sqrt{\frac{K_{\text{B}}}{1+K_{\text{B}}}}{\bf H}_{\text{r}}^{\text{LOS}}+\sqrt{\frac{1}{1+K_{\text{B}}}}{\bf H}_{\text{r}}^{\text{NLOS}}\big{)}$, where Rician factor $K_{\text{B}}=10$, $\beta^{\text{BR}}$ is the path gain modeled as $\beta^{\text{BR}}\text{[dB]}=-30-22\log 10(d_{\text{B-R}})$ with $d_{\text{B-R}}$ being the BS-RIS distance in meters, ${\bf H}_{\text{r}}^{\text{NLOS}}$ is the non-line-of-sight (NLOS) component modeled as $[{\bf H}_{\text{r}}^{\text{NLOS}}]_{ij}\sim{\cal C}{\cal N}(0,{\bf I})$, and ${\bf H}_{\text{r}}^{\text{LOS}}$ is the LOS component.
 The LOS component ${\bf H}_{\text{r}}^{\text{LOS}}$ is a function of the BS and RIS locations and is modeled as
-${\bf H}_{\text{r}}^{\text{LOS}}={\bf b}_{\text{RIS}}(\psi_{2},\theta_{2}){\bf b}_{\text{BS}}(\psi_{1},\theta_{1})^{H}$ (43)
-where $\psi_{1}$ and $\theta_{1}$ are the azimuth and elevation angles of departure (AoD) from the BS to the IRS, and ${\bf b}_{\text{BS}}(\psi_{1},\theta_{1})$ is the BS antenna array steering vector, with its $n$th entry as
-$[{\bf b}_{\text{BS}}(\psi_{1},\theta_{1})]_{n}=e^{\frac{2\pi(n-1)\Delta_{\text{BS}}}{\lambda_{c}}}\cos(\psi_{1})\cos(\theta_{1}),\ n=1,\ldots,N$
-where $\Delta_{\text{BS}}$ is the distance between two adjacent BS antennas, $\lambda_{c}$ is the carrier wavelength, and $\cos(\psi_{1})\cos(\theta_{1})=\frac{x_{\text{RIS}}-x_{\text{BS}}}{d^{\text{B-R}}}$.
+${\bf H}_{\text{r}}^{\text{LOS}}={\bf b}_{\text{RIS}}(\phi_{2},\theta_{2}){\bf b}_{\text{BS}}(\phi_{1},\theta_{1})^{H}$ (45)
+where $\phi_{1}$ and $\theta_{1}$ are the azimuth and elevation angles of departure (AoD) from the BS to the IRS, and ${\bf b}_{\text{BS}}(\phi_{1},\theta_{1})$ is the BS antenna array 3D steering vector.
+Assume the BS antenna elements are along the $x$-coordinate, the $n$th entry of ${\bf b}_{\text{BS}}(\phi_{1},\theta_{1})$ is given by
+$[{\bf b}_{\text{BS}}(\phi_{1},\theta_{1})]_{n}=e^{-j\frac{2\pi(n-1)\Delta_{\text{BS}}}{\lambda_{c}}\cos(\phi_{1})\sin(\theta_{1})},\ n=1,\ldots,N$
+where $\Delta_{\text{BS}}$ is the distance between two adjacent BS antennas, $\lambda_{c}$ is the carrier wavelength, and $\cos(\phi_{1})\sin(\theta_{1})=\frac{x_{\text{RIS}}-x_{\text{BS}}}{d_{\text{B-R}}}$.
 We assume $\Delta_{\text{BS}}=\frac{\lambda_{c}}{2}$.
-Similarly, $\psi_{2}$ and $\theta_{2}$ are the azimuth and elevation angles of arrival (AoA) from the BS to the RIS, and ${\bf b}_{\text{RIS}}(\psi_{2},\theta_{2})$ is the RIS element streaming vector, with the $m$th entry, for $m\in{\cal M}$, given by
-$[{\bf b}_{\text{RIS}}(\psi_{2},\theta_{2})]_{m}=e^{\frac{2\pi\Delta_{\text{RIS}}}{\lambda_{c}}}[y_{m}\sin(\psi_{2})\cos(\theta_{2})+z_{m}\sin(\theta_{2})]$
-where $\Delta_{\text{RIS}}$ is the distance between two adjacent RIS elements, $y_{m}=\text{mod}(m-1,M_{y})$ with $M_{y}$ being the number of RIS elements along the $y$-axis, $z_{m}=\lfloor m/M_{y}\rfloor+1$, $\sin(\psi_{2})\cos(\theta_{2})=\frac{y_{\text{BS}}-y_{\text{RIS}}}{d_{\text{B-R}}}$, and $\sin(\theta_{2})=\frac{z_{\text{BS}}-z_{\text{RIS}}}{d_{\text{B-R}}}$.
-We assume $\Delta_{\text{RIS}}=\frac{\lambda_{c}}{2}$.
-We consider a NLOS channel between the RIS and each user modeled as Rayleigh fading.
-The channel ${\bf h}^{\text{r}}_{ik}$ between the RIS and user $k$ in group $i$ is generated i.i.d as ${\bf h}^{\text{r}}_{ik}\sim{\cal C}{\cal N}({\bf 0},\beta^{\text{r}}_{ik}{\bf I})$, where $\beta^{\text{r}}_{ik}$ is the path gain modeled as $\beta^{\text{r}}_{ik}[\text{dB}]=-30-22\log 10(d^{\text{R-U}}_{ik})$, with $d^{\text{R-U}}_{ik}$ being the distance between the RIS and the user.
-We also model the channel between BS and user $k$ in group $i$ as Rayleigh fading, which is generated i.i.d as ${\bf h}^{\text{d}}_{ik}\sim{\cal C}{\cal N}({\bf 0},\beta^{\text{d}}_{ik}{\bf I})$, where $\beta^{\text{d}}_{ik}$ is the path gain modeled as $\beta^{\text{d}}_{ik}[\text{dB}]=-32.6-36.7\log 10(d^{\text{B-U}}_{ik})$, with $d^{\text{B-U}}_{ik}$ being the respective BS-user distances.
+Similarly, ${\bf b}_{\text{RIS}}(\phi_{2},\theta_{2})$ is the $M\times 1$ RIS element streaming vector, with $\phi_{2}$ and $\theta_{2}$ being the azimuth and elevation angles of arrival from the BS to the RIS, and the $m$th element, for $m\in{\cal M}$, is given by
+$[{\bf b}_{\text{RIS}}(\phi_{2},\theta_{2})]_{m}=e^{-j\frac{2\pi\Delta_{\text{RIS}}}{\lambda_{c}}[y_{m}\sin(\phi_{2})\sin(\theta_{2})+z_{m}\cos(\theta_{2})]}$ (46)
+where $\Delta_{\text{RIS}}=\frac{\lambda_{c}}{2}$ is the distance between two adjacent RIS elements, $y_{m}=\text{mod}(m-1,M_{y})$ with $M_{y}$ being the number of RIS elements along the $y$-axis, $z_{m}=-\lfloor(m-1)/M_{y}\rfloor$, $\sin(\phi_{2})\sin(\theta_{2})=\frac{y_{\text{BS}}-y_{\text{RIS}}}{d_{\text{B-R}}}$, and $\cos(\theta_{2})=\frac{z_{\text{BS}}-z_{\text{RIS}}}{d_{\text{B-R}}}$.
+We assume a LOS channel between the RIS and each user.
+The channel ${\bf h}^{\text{r}}_{ik}$ between the RIS and user $k$ in group $i$ is generated i.i.d as ${\bf h}^{\text{r}}_{ik}=\beta^{\text{r}}_{ik}\big{(}\sqrt{\frac{K_{\text{r}}}{1+K_{\text{r}}}}{\bf h}_{ik}^{\text{r,LOS}}+\sqrt{\frac{1}{1+K_{\text{r}}}}{\bf h}_{ik}^{\text{r,NLOS}}\big{)}$, where Rician factor $K_{\text{r}}=10$, $\beta^{\text{r}}_{ik}$ is the path gain modeled as $\beta^{\text{r}}_{ik}[\text{dB}]=-30-22\log 10(d^{\text{R-U}}_{ik})$, with $d^{\text{R-U}}_{ik}$ being the distance between the RIS and the user.
+The NLOS component is ${\bf h}_{ik}^{\text{r,NLOS}}\sim{\cal C}{\cal N}({\bf 0},{\bf I})$.
+The LOS component ${\bf h}_{ik}^{\text{r,LOS}}={\bf b}_{\text{RIS}}(\phi_{ik}^{\text{r}},\theta^{\text{r}})$ is similar as in (46), where $\phi_{ik}^{\text{r}}$ and $\theta^{\text{r}}$ being the azimuth and elevation AoD from the IRS to each user, with $\sin(\phi_{ik}^{\text{r}})\sin(\theta^{\text{r}})=\frac{y_{ik}-y_{\text{RIS}}}{d^{\text{R-U}}_{ik}}$, and $\cos(\theta^{\text{r}})=\frac{z_{ik}-z_{\text{RIS}}}{d^{\text{R-U}}_{ik}}$.
+The channel between BS and user $k$ in group $i$ is modeled as a NLOS channel with Rayleigh fading, which is generated i.i.d as ${\bf h}^{\text{d}}_{ik}\sim{\cal C}{\cal N}({\bf 0},\beta^{\text{d}}_{ik}{\bf I})$, where $\beta^{\text{d}}_{ik}$ is the path gain modeled as $\beta^{\text{d}}_{ik}[\text{dB}]=-32.4-36.7\log 10(d^{\text{B-U}}_{ik})$, with $d^{\text{B-U}}_{ik}$ being the respective BS-user distances.
 The receiver noise power is set to $\sigma^{2}=-100~{}\text{dBm}$.
 […]
-VII-A Convergence Behavior of the Proposed Algorithms
-We first study the convergence behavior of our proposed AMBF algorithm in Algorithm 1 for the QoS problem ${\cal P}_{o}$, which alternatingly solves two multicast subproblems at the BS and the RIS using our proposed fast algorithms.
+VI-A Performance for the QoS Problem
+VI-A1 Convergence Behavior
+We first study the convergence behavior of the proposed AMBF (Algorithm 1) for ${\cal P}_{o}$.
+Denote the total transmit power at the BS as $P_{\text{tot}}=\sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$.
 Fig.
-0(a) shows the convergence behavior of solving subproblem $\widetilde{{\cal S}}_{{\bf w}}$ using PSA in a specific iteration of the AMBF algorithm.
+2(a) shows the trajectory of $P_{\text{tot}}$ over AO iterations by Algorithm 1 for three random channel realizations.
+We observe that Algorithm 1 converges fast, especially given that fast algorithms are used for each subproblem.
+Fig.
+2(b) shows the convergence behavior of solving subproblem $\widetilde{{\cal S}}_{{\bf w}}$ for RIS phase shift ${\bf e}$ using PSA in the first AO iteration of AMBF.
 The convergence behavior for three random channel realizations are shown.
-We see PSA converges after $1000\sim 2000$ iterations.
-Note that the closed-form updates for PSA in each iteration is computationally cheap, and thus, the overall computation time for $1000\sim 2000$ iterations is fast.
-Furthermore, denote the total transmit power at the BS as $P_{\text{tot}}=\sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}$.
-Fig.
-0(b) shows the trajectory of $P_{\text{tot}}$ over iterations by Algorithm 1 for three random channel realizations.
-We observe that our proposed AMBF approach converges within about $6\sim 12$ iterations, which is relatively fast, especially given that we use fast algorithms for computing the solution to each subproblem.
-(a) The subproblem $\widetilde{{\cal S}}_{{\bf w}}$.
-(b) The overall AMBF Algorithm.
-Figure 1: Convergence behavior of the AMBF algorithm in Algorithm 1 ($N=16,M=100$, $(G,K)=(2,2)$, $\gamma=10$ dB).
-Fig.
-2 shows the convergence behavior of the MMF problem ${\cal S}_{o}$ solved by Algorithm 2 under three random channel realizations.
-We see that for two channel realizations, Algorithm 2 converges very fast within just a few iterations.
-Even for a worse scenario under the third channel realization, Algorithm 2 converges after $800$ iterations.
-Note that the PSA updates in each iteration is closed-form and computationally cheap.
-Figure 2: Convergence behavior of the Algorithm 2 ($N=30$, $M=100$, $(G,K)=(2,2)$, $\gamma=10$ dB, $P=10$ dB).
-VII-B Performance Comparison for the QoS Problem
-We now evaluate the performance of our proposed AMBF algorithm in Algorithm 1. For performance comparison, we consider the following methods:
+We see PSA converges after $500\sim 1400$ iterations.
+Note that the convergence speed becomes faster in the subsequent AO iterations as the algorithm converges.
+Furthermore, the closed-form updates for PSA in each iteration is computationally inexpensive as shown in the complexity analysis.
+Thus, the overall computation time is fast despite that the number of iterations can be large especially in the initial AO iteration.
+(a) The overall AMBF for ${\cal P}_{o}$.
+(b) The subproblem $\widetilde{{\cal S}}_{{\bf w}}$.
+Figure 2: Convergence behavior of Algorithm 1 ($N=32,M=100$, $(G,K)=(2,3)$, $\gamma=10$ dB).
+VI-A2 Performance Comparison
+For performance comparison, we consider the following methods:
   1. i)
@@
   3. iii)
-    Direct AO [19]: Apply the AO approach directly to the QoS problem ${\cal P}_{o}$, where the subproblem w.r.t.
-${\bf e}$ is a feasibility problem with the SINR constraints in (3b) and the unit-modulus constraints in (3c).
-  4. iv)
-    AMBF-SDR: Apply Algorithm 1, except that instead of solving subproblem $\tilde{{\cal S}}_{\bf w}$ w.r.t.
-${\bf e}$, we relax ${\cal S}_{{\bf w}}^{\text{eq}}$ by dropping the constraint in (10c) and solve the relaxed problem:
-    ${\cal S}_{{\bf w}}^{\text{Relaxed}}:\max_{{\bf e}}\min_{i,k}$ $\ \frac{\textrm{SINR}_{ik}}{\gamma_{ik}}$ s.t. $|e_{m}|^{2}\leq 1,m\in{\cal M},$
-    We solve ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ by using its epigraph form and applying SDR with a bi-section search for the objective value (i.e., let ${\bf X}={\bf e}{\bf e}^{H}$ and drop the rank-1 condition).
-Then, we recover ${\bf e}$ from ${\bf X}$ using the Gaussian randomization method.
+    DirectSCA [34]: An SCA method to directly solve ${\cal P}_{o}$.
+Note that this SCA algorithm was developed originally for the RIS-assisted joint beamforming in the multiuser unicast scenario.
+We can straightforwardly extend it to solve ${\cal P}_{o}$ in the multicast scenario.
+The SINR constraints w.r.t.
+${\bf w}$ and ${\bf e}$ in (2b) are convexified and iteratively solved via a convex solver such as CVX [35].
+This method serves as a competitive benchmark.
+  4. vi)
+    IP Method: Apply the interior-point (IP) algorithm for the joint optimization problem. In particular, we adopt the MATLAB built-in non-linear optimization solver Fmincon to solve ${\cal P}_{o}$.
   5. v)
-    AMBF-SCA: Similar to AMBF-SDR that apply Algorithm 1, except that we solve ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ by SCA with bi-section search over $t$.
-In particular, we convexify ${\cal S}_{{\bf w}}^{\text{Relaxed}}$ into a sequence of convex approximation problems and solved them using a standard convex solver.
-  6. vi)
-    IP Method: Apply MATLAB non-linear optimization solver Fmincon to ${\cal P}_{o}$. It uses the interior-point (IP) algorithm to solve ${\cal P}_{o}$.
-Figure 3: Total transmit power $P_{\text{tot}}$ vs. $M$ ($N=16$, $(G,K)=(2,2)$, $\gamma=10$ dB).
-TABLE I: Average Computation Time (sec).
-$M$
-25
-50
-100
-200
-300
-RIS-SDR
-24.73
-46.95
-84.68
-672
-1743
-RIS-SCA
-23.12
-38.52
-62.17
-436
-1349
-IP Method
-3.25
-10.73
-58.33
-796
-2062
-Algorithm 1
-2.19
-3.12
-5.22
-8.83
-13.62
-VII-B1 Performance over the Number of RIS elements $M$
+    DirectAO [29]: An algorithm that directly applies AO to ${\cal P}_{o}$, where the subproblem for ${\bf e}$ is a feasibility problem to satisfy the SINR constraints in (2b) and the unit-modulus constraints in (2c).
+Each subproblem is solved by an SDR-based algorithm.
 Fig.
@@
 the number of RIS elements $M$ by different methods, for the target SINR $\gamma=10$ dB.
-We see that the BS transmit power for random RIS, direct AO, and IP Method remain roughly flat as $M$ increases.
-This indicates that they do not provide an effective RIS beamforming solution.
-In contrast, under Algorithm 1, the resulting BS transmit power is reduced substantially by $4$ dB as $M$ increases from 25 to 300.
-RIS-SCA has a nearly identical performance to Algorithm 1, while RIS-SDR performs worse than Algorithm 1 with a noticeable gap as large as $2$ dB.
-This is expected, as SDR is an approximation method with a degraded performance when there are many constraints, and the problem size is relatively large.
-We also compare the average computation time of different algorithms in Table I.
-The computational advantage of Algorithm 1 over RIS-SDR, RIS-SCA, and IP Method as the number of RIS elements $M$ increases is clearly observed, demonstrating the scalability of Algorithm 1 over $M$.
-In particular, the computational complexity of Algorithm 1 only grows mildly with $M$.
-Thus, Algorithm 1 provides both the best performance and the fastest computational time than the other three methods, especially for the large value of $M$.
-VII-B2 Performance over Target SINR $\gamma$
+The performance of random RIS remains roughly flat as $M$ increases, as it does not provide an effective RIS beamforming solution.
+DirectAO can only provides a small improvement over random RIS, which shows solving the feasibility problem for RIS subproblem does not give an effective solution.
+Also, the algorithm entails very high computation complexity, and thus its performance is only shown for $M\leq 100$.
+In contrast, a substantial gain can be achieved when RIS beamforming is effectively optimized.
+As we see, DirectSCA performs the best with the lowest power.
+This is expected as it is a joint optimization approach to solve ${\cal P}_{o}$ directly, and SCA performs well in general.
+However, the main drawback of this approach is the cost of impractically high computational complexity as we will show later.
+In addition, as $M$ becomes large, the numerical stability of this method is reduced noticeably.
+In particular, the frequency that the CVX solver fails to find a solution increases to $40\%\sim 50\%$ when $M>100$.
+In comparison, the performance of the proposed Algorithm 1 closely follows that of DirectSCA, with a gap about $1$ dB.
+Also, the algorithm always provides a converged solution.
+IP method for this non-convex problem often has trouble to converge properly, and the resulting performance deteriorates as $M$ increases, with a large gap observed for $M>200$.
+Figure 3: Total transmit power $P_{\text{tot}}$ vs. $M$ ($N=32$, $(G,K)=(2,3)$, $\gamma=10$ dB).
+TABLE I: Average Computation Time (sec).
+$M$ 50 100 200 300 400
+Algorithm 1 6.49 17.03 22.2 24.8 37.5
+DirectSCA [34] 646 935 2364 2425 4403
+IP Method 15.9 19.9 35.1 39.1 47.7
+DirectAO 269 1015   –   –    –
+We now compare the average computation time of different algorithms in Table I.
+The computational advantage of Algorithm 1 over DirectSCA, IP Method, and DirectAO is clearly observed as $M$ increases, demonstrating the excellent scalability of Algorithm 1 over $M$.
+DirectSCA has prohibitively high computational complexity as $M$ increases, making it impractical for real implementation.
+In contrast, the computational time of Algorithm 1 is several magnitudes lower and only grows mildly with $M$.
+Thus, Algorithm 1 is a much more efficient algorithm offering competitive performance with significantly lower computational complexity as compared to DirectSCA, especially for the large value of $M$.
 Fig.
 4 shows $P_{\text{tot}}$ vs.
-target SINR $\gamma$ by different methods.
-We see that direct AO, Random RIS, and IP method provide either no or only very small improvement over the case of No RIS.
-A substantial gain can be achieved when the RIS passive beamforming is effectively optimized to improve the channel environment between the BS and users.
-In particular, our proposed Algorithm 1 provides about $4$ dB power reduction over No RIS for $\gamma\geq 4$ dB.
-Similar to Fig.
-3, RIS-SCA has nearly identical performance to Algorithm 1, while RIS-SDR performs worse than Algorithm 1 with a noticeable $2\sim 3$ dB gap.
-Figure 4: Total transmit power $P_{\text{tot}}$ vs. target SINR $\gamma$ ($N=16,M=100$, $(G,K)=(2,2)$).
-VII-C Performance Comparison for the MMF Problem
-We now evaluate the proposed RIS-assisted multicast beamforming design for the MMF problem. For performance comparison, besides our proposed Algorithm 2, we consider the following methods:
+target SINR $\gamma$ by different methods, for $M=256~{}(16\times 16)$.
+We see that the proposed Algorithm 1 closely follows DirectSCA and provides substantial power reduction over Random RIS and IP method for all values of $\gamma$, with a roughly constant gap.
+Figure 4: Total transmit power $P_{\text{tot}}$ vs. target SINR $\gamma$ ($N=32,M=256$, $(G,K)=(2,3)$).
+VI-B Performance for the MMF Problem
+VI-B1 Convergence Behavior
+Fig.
+5 shows the convergence behavior of Algorithm 2 for the MMF problem ${\cal S}_{o}$ under three random channel realizations.
+We see that APSA converges in $1000\sim 2000$ iterations.
+Since the closed-form updates for APSA in each iteration is computationally inexpensive, the overall computation time is fast despite the number of iterations is large.
+As we will see later, the overall computational time is low.
+Figure 5: Convergence behavior of the Algorithm 2 ($N=32$, $M=100$, $(G,K)=(2,3)$).
+VI-B2 Performance Comparison
+For performance comparison, besides our proposed Algorithm 2, we consider the following methods:
   1. 1.
@@
   3. 3.
-    Relaxed RIS-MMF: Only solve the relaxed problem $\widetilde{{\cal S}}_{o}$ in (VI-A), i.e., Algorithm 2 without the final processing step to obtain the feasible solution ${\bf e}$.
-We use this method to measure the performance loss in the final processing step in Algorithm 2 and quantify how good the relaxed problem $\widetilde{{\cal S}}_{o}$ is as an approximation of the original problem ${\cal S}_{o}$.
+    QoS2MMF: Solving ${\cal S}_{o}$ by iteratively solving ${\cal P}_{o}$ as discussed below Proposition 1 in Section V.
   4. 4.
-    IP Method: Use MATLAB non-linear optimization solver Fmincon to solve the MMF problem ${\cal S}_{o}$.
-  5. 5.
-    The alternating projected gradient (APG) method [18]: APG is a recently proposed iterative algorithm to maximize the achievable sum group-rate of all groups.
-The group rate is defined as the minimum rate among users in the group, which corresponds to the minimum SINR in the group.
-Note that the design objective for the APG method [18] is different from the max-min SINR (or rate) objective in our MMF problem.
-However, for the single-group multicast scenario when $G=1$, the two objectives becomes identical.
-Since there are no existing RIS-assisted methods designed for the MMF problem, we compare Algorithm 2 with APG in this special single-group case.
-Besides [18], we note that [17] has also proposed the RIS-assisted multicast beamforming algorithms for the sum-group-rate maximization.
-Since it is shown in [18] that APG outperforms the method in [17] in both performance and computational complexity, we choose APG as the state-of-the-art method for comparison.
-In the following simulation results, we set the BS transmit power $P=10~{}{\text{dBm}}$.
-VII-C1 Multiple Groups
+    IP Method: Adopt the IP algorithm as a joint optimization method to directly solve the MMF problem ${\cal S}_{o}$ by using the MATLAB non-linear optimization solver Fmincon.
+Figure 6: Average minimum SINR vs. $M$ ($N=32$, $(G,K)=(2,3)$).
+TABLE II: Average Computation Time (sec) ($N=32$, $(G,K)=(2,3)$).
+$M$ 50 100 200 300 400
+Algorithm 1 1.19 1.22 1.64 1.75 1.96
+QoS2MMF 27.4 32.5 37.8 34.1 58.4
+IP Method 7.47 9.84 15.3 22.2 27.5
+Note that we select the above methods for comparison as there are no existing methods available for the RIS-assisted MMF problem,555 The sum-group-rate maximization problem is considered as the objective in [27, 28], where the group rate is defined as the minimum rate (i.e., SINR) within the group.
+Since this objective cannot ensure fairness among groups, the solutions can result in some group having nearly zero rate, which is undesirable.
+nor a tight upper bound for meaningful comparison.
+In the following simulation results, we set the BS transmit power $P=10~{}{\text{dBm and $\gamma_{ik}=\gamma=10$~{}dB.}}$..
 Fig.
 6 shows the average minimum SINR vs.
-the number of RIS elements $M$ by different methods for $N=30$.
-We see that the minimum SINR achieved by Algorithm 2 increases noticeably as $M$ increases, significantly better than random RIS and IP Method, especially when $M>N$.
-This demonstrates the effectiveness of Algorithm 2 to design RIS beamforming to improve the user channel conditions.
-Furthermore, we see that Algorithm 2 and Relaxed RIS-MMF provide a nearly identical performance for all values of $M$.
-This indicates that the solution to $\widetilde{{\cal S}}_{o}$ in Algorithm 2 is mostly feasible to ${\cal S}_{o}$, and the final processing step causes a negligible performance loss.
-This also demonstrates that the relaxed problem $\widetilde{{\cal S}}_{o}$ is a good approximation of the original problem ${\cal S}_{o}$.
-Figure 5: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(2,2)$).
-Figure 5: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(2,2)$).
-Figure 6: Average minimum SINR vs. $N$ ($M=100$, $(G,K)=(2,2)$).
+the number of RIS elements $M$ by different methods.
+We see that the minimum SINR achieved by Algorithm 2 is substantially higher than random RIS and IP method.
+Also, the SINR gain under Algorithm 2 by increasing $M$ from $50$ to $500$ is nearly 3 dB, significantly more than random RIS and IP method.
+This shows that Algorithm 2 can provide an effective beamforming solution for RIS elements to improve the performance.
+Also, the performance of Algorithm 2 and QoS2MMF are close, demonstrating that Algorithm 2 is an effective first-order algorithm that provides a good performance.
+In contrast, IP method cannot produce an effective solution for the MMF problem and its performance is similar to random RIS.
+We compare the average computation times of these methods in Table II.
+The computational advantage of Algorithm 2 over QoS2MMF and IP method can be clearly seen.
 Fig.
 7 shows the average minimum SINR vs.
-the number of BS antennas $N$ by different methods for $M=100$.
-Again, we see that the gain of random RIS and IP Method over no RIS is very small.
-Our Algorithm 2 provides about $1.5\sim 2$ dB additional SINR gain over these methods for different values of $N$.
+the number of BS antennas $N$.
+Since IP method is not effective, we exclude this method in comparison.
+Again, we see that Algorithm 2 provides about $3$ dB SINR gain over random RIS under $M=100$ for different values of $N$, and its performance is close to that of QoS2MMF.
 In Fig.
-8, we show the average minimum SINR vs.
-the number of users $K$ per group for $N=30$ and $M=100$.
-Algorithm 2 outperforms no RIS, random RIS, and IP Method for all values of $K$, especially for smaller values of $K$.
-This is expected, since the gain of the optimized multicast beamforming is higher for a smaller user group.
-The gain reduces as the group size increases, and the difference among algorithms reduces.
-Figure 7: Average minimum SINR vs. $K$ ($N=30,M=100$, $(G,K)=(2,2)$).
-Figure 7: Average minimum SINR vs. $K$ ($N=30,M=100$, $(G,K)=(2,2)$).
-Figure 8: Average minimum SINR vs. $M$ ($N=30$, $(G,K)=(1,4)$).
-VII-C2 Single Group
-We now compare our proposed algorithm with other algorithms, including the APG method [18], in the single group scenario.
-Fig.
-8 shows the average minimum SINR vs.
-the number of RIS elements $M$ for $G=1$, $K=4$, and $N=30$.
-We see that Algorithm 2 outperforms all other algorithms significantly.
-In particular, our proposed algorithm provides $1\sim 2.5$ dB gain over APG for $M$ ranging from $25$ $\sim 300$.
-We also compare the average computation time of different methods in Table II for the setting used in Fig.
-8.
-IP method has the highest computational complexity.
-Algorithm 2 and APG have a similar computation time as $M$ increases.
-Although APG is slightly faster than Algorithm 2, the difference is not significant.
-TABLE II: MMF – Average Computation Time (sec).
-$M$
-25
-50
-100
-200
-300
-APG [18]
-0.29
-0.71
-1.61
-2.05
-2.7
-IP Method
-8.73
-12.17
-36.63
-197
-476
-Algorithm 2
-0.45
-0.98
-2.1
-2.6
-3.2
-VII-C3 Unicast Scenario
-We also consider the RIS-assisted downlink beamforming in a unicast scenario with $4$ users by setting $G=4$ and $K=1$.
-Note that in this case, Algorithm 2 and APG have different design objectives: APG aims to maximize the sum rate of all users, while Algorithm 2 aims to maximize the minimum SINR among users.
-For comparison, we plot the average minimum SINR and the sum rate performance over $M$ in Figs.
-10 and 10, respectively, for $N=30$.
-We see that Algorithm 2 outperforms the APG method and other methods for both the minimum SINR and the sum rate objectives.
-Figure 9: Unicast: Average minimum SINR vs. $M$ ($N=30$, $G=4$ users).
-Figure 9: Unicast: Average minimum SINR vs. $M$ ($N=30$, $G=4$ users).
-Figure 10: Unicast: Average sum rate vs. $M$ ($N=30$, $G=4$ users).
-VIII Conclusion
-In this paper, we consider joint BS and RIS beamforming design for RIS-assisted downlink multi-user beamforming, aiming to provide efficient and scalable algorithms.
-We have shown that the joint BS and RIS beamforming optimization for the QoS problem can be broken down into a BS multicast beamforming QoS problem and an RIS passive multicast beamforming MMF problem, and we have proposed an AMBF approach to solve the two subproblems alternatingly.
-We have utilized the optimal multicast beamforming structure to solve the BS multicast beamforming QoS problem efficiently.
-We have also developed a PSA-based fast iterative algorithm to solve the challenging RIS passive multicast MMF problem, which only contains closed-form updates with cheap computation.
-We have shown that our AMBF approach is also applicable to the unicast scenario, making it an effective method for general RIS-assisted downlink beamforming.
-For the RIS-assisted MMF problem, based on the relation of the QoS and MMF problems, we have used the optimal BS beamforming structure again to transform the MMF problem and propose a PSA-based iterative algorithm.
-Similar to that for the QoS problem, The algorithm uses the closed-form updates per iteration, leading to a highly computationally efficient solution.
-Simulation results show that our proposed algorithms outperform the alternative methods and that the computational complexity is substantially lower than most existing methods and grows linearly with the number of RIS reflective elements and BS antennas.
+8, we show the average minimum SINR as the number of users $K$ per group increases.
+Again, Algorithm 2 and QoS2MMF perform similarly.
+The gain of Algorithm 2 over no RIS or random RIS is consistent for all values of $K$ with $2.5\sim 3$ dB gain.
+Figure 7: Average minimum SINR vs. $N$ ($M=100$, $(G,K)=(2,3)$).
+Figure 8: Average minimum SINR vs. $K$ ($N=32,M=100$, $G=2$).
+VII Conclusion
+In this paper, we consider joint BS and RIS beamforming design for RIS-assisted downlink multicasting, aiming to provide efficient and scalable algorithms.
+We show that the QoS problem can be broken down into a BS multicast beamforming QoS problem and an RIS passive multicast beamforming MMF problem, and we propose an AMBF approach to solve the two subproblems alternatingly.
+Our proposed fast algorithm for AMBF utilizes the optimal BS mutlicast beamforming structure for computational efficiency and employs a PSA-based fast first-order algorithm to solve the RIS passive multicast MMF problem via iterative closed-form updates.
+The overall computational complexity grows linearly with the number of BS antennas and RIS elements.
+For the RIS-assisted MMF problem, we utilize the optimal BS beamforming structure to transform the MMF problem and propose an APSA algorithm using all closed-form updates, leading to a highly computationally efficient solution with complexity linear the number of BS antennas and RIS elements.
+Simulation results shows that our proposed algorithms provide favorable performance with low cost in computational time as compared with existing methods.
 Appendix A Proof of Proposition 1
-Proof:
-We can equivalently rewrite ${\cal S}_{o}$ as
+Proof: We can equivalently rewrite ${\cal S}_{o}$ as
 $\max_{{\bf w}}$ $\ \bigg{(}\max_{{\bf e}:|e_{m}|^{2}=1,m\in{\cal M}}\min_{i,k}\frac{\textrm{SINR}_{ik}}{\gamma_{ik}}\bigg{)}$ s.t. $\ \sum_{i=1}^{G}\|{\bf w}_{i}\|^{2}\leq P.$
-For the QoS problem ${\cal P}_{o}$, consider its equivalent formulation ${\cal P}_{2}$ in (III-A) derived in Section III-A.
+For the QoS problem ${\cal P}_{o}$, consider its equivalent formulation ${\cal P}_{2}$ in (5) derived in Section III-A.
 Comparing the above problem and ${\cal P}_{2}$, we see that the objective functions and the constraints in these two problems have the same forms but switched, indicating that they are the inverse problems (and equivalently ${\cal S}_{o}$ and ${\cal P}_{o}$ are the inverse problems).
@@
 As a result, we have the relations in (27).
-∎